### v1

# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

Abstract
Learning disentangled representations in an unsupervised manner is a fundamental challenge in machine learning. Solving it may unlock other problems, such as generalization, interpretability, or fairness. While remarkably difficult to solve in general, recent works have shown that disentanglement is provably achievable under additional assumptions that can leverage geometrical constraints, such as local isometry. To use these insights, we propose a novel perspective on disentangled representation learning built on quadratic optimal transport. Specifically, we formulate the problem in the Gromov-Monge setting, which seeks isometric mappings between distributions supported on different spaces. We propose the Gromov-Monge-Gap (GMG), a regularizer that quantifies the geometry-preservation of an arbitrary push-forward map between two distributions supported on different spaces. We demonstrate the effectiveness of GMG regularization for disentanglement on four standard benchmarks. Moreover, we show that geometry preservation can even encourage unsupervised disentanglement without the standard reconstruction objective - making the underlying model decoder-free, and promising a more practically viable and scalable perspective on unsupervised disentanglement.
1 Introduction
Learning low-dimensional representations of high-dimensional data is a fundamental challenge in unsupervised deep learning [5, 43], where the emphasis is put on learning representations that allow for efficient adaptation across a wide range of tasks [5, 28, 43]. Disentanglement [5, 27, 28, 43, 60] has shown significant promise in facilitating such generalization [5, 27, 28, 43, 44, 29, 60, 30, 4], alongside interpretability and fairness [42, 67]. Most works [5, 27, 33, 13, 43, 60] regard disentanglement as a one-to-one map between learned representations and ground-truth latent factors; seeking to recover these factors from data alone in an unsupervised fashion.
While unsupervised disentanglement is theoretically impossible [43], the inductive biases of autoencoder architectures ensure effective disentanglement in practice [59, 74, 29]. Most approaches operate on variational autoencoder (VAE) frameworks [35], using objectives that match latent VAE posteriors to factorized priors [27, 33, 38, 10, 13]. Recent works [29, 51, 31] provide a new perspective, showing how geometric constraints on representation spaces may enable disentanglement. In particular, Horan et al. [29] demonstrate that unsupervised disentanglement is always possible under the assumption of local isometry to the data and sufficient non-Gaussianity of the generative factors, thus supporting the isometry desideratum.
In this work, we show how these geometric desiderata can be effectively quantified through the lens of optimal transport (OT) theory [61, 55], by treating mapping to or from the latent space as transport maps $T$ from or to the data manifold, respectively. While the more classical tools described in the OT toolbox [55] provide tools to align distributions supported on the same space $\mathcal{X}$, using an inter-domain cost $c(\mathbf{x},\mathbf{y})$ for any two points $\mathbf{x},\mathbf{y}\in\mathcal{X}$, such approaches are not well defined when mapping between latent and data space with very different dimensionalities, for which no “natural” cost function exists.
Instead, we rely in this work on the Gromov-Wasserstein (GW) formalism [64, 48, 69, 63] formulation of OT, which considers instead intra-domain costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ in each space. This formulation seeks instead the most geometry-preserving mapping between two distributions by favoring maps that minimize the distortion between intra-domain costs by comparing $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime}),c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$. While this distortion itself can be used as a regularization for a given map $T$ as done in Nakagawa et al. [51], we borrow a page from the Monge gap formalism proposed by Uscidda and Cuturi [68] to propose a modified and better-posed regularizer that substracts to the distortion of $T$, the best distortion that can be achieved by any map (or a proxy thereof).
In a nod to [68], we propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving). In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results, showing notably that the GMG and its finite sample version are weakly convex. We show how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41]. Motivated by this formalism, we build upon the Monge gap - a regularizer introduced in Uscidda and Cuturi [68] that measures whether a map $T$ transports a reference distribution at minimal displacement cost - to propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving). In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results highlighting that the GMG and its finite sample version are weakly convex. We lay out how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41].
Our experiments on four standard disentangled representation learning benchmarks show that the integration of these geometry-preserving desiderata through the Gromov-Monge Gap (GMG) significantly improves disentanglement performance across various methods, from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with support factorization [60], outperforming a distortion-based regularization. Moreover, we demonstrate that these geometric regularizations can replace the standard reconstruction loss, enabling measurable unsupervised disentanglement even without a decoder, which is not feasible in standard frameworks that rely on the decoder-based reconstruction term to prevent collapse. This finding suggests the potential for more scalable unsupervised disentangled representation learning approaches and bridges to popular, weakly- or self-supervised encoder-only representation learning methods [14, 72, 3, 24].
2 Background and Related Works
2.1 Disentangled Representation Learning
The Disentanglement Formalism.
Disentanglement has varying operational definitions [28, 43, 60]. In this work, we follow the common understanding [43, 44, 67, 60] where data $\mathbf{x}$ is generated by a process $p(\mathbf{x}|\mathbf{z})$ operating on ground-truth latent factors $\mathbf{z}\sim p(\mathbf{z})$, modeling underlying source of variations (s.a. object shape, color, background…). Given a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, unsupervised disentangled representation learning aims to find a mapping $e_{\phi}$ s.t. $e_{\phi}(\mathbf{x}_{i})\approx\mathbb{E}[\mathbf{z}|\mathbf{x}_{i}]$, up to element-wise transformations. Notably, this is to be achieved without prior information on $p(\mathbf{z})$ and $p(\mathbf{x}|\mathbf{z})$.
Unsupervised Disentanglement through Prior Matching.
Most unsupervised disentanglement methods operate on variational autoencoders (VAEs)[35], which define a generative model of the form $p_{\theta}(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z})$. Here, $p_{\theta}(\mathbf{x}|\mathbf{z})$ is a product of exponential family distributions with parameters computed by a decoder $d_{\theta}(\mathbf{z})$. The latent prior $p(\mathbf{z})$ is usually chosen to be a normal Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$, and the probabilistic encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ is realized through a neural network $e_{\phi}(\mathbf{x})$ that predicts the parameters of the latent such that $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|e_{\phi}(\mathbf{x}))$. The $\beta$-VAE [27]
$\mathcal{L}_{\beta}(\theta,\phi)\vcentcolon=\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})-\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))]$ (1)
achieves disentanglement by enforcing stronger, $\beta$-weighted prior matching on top of the reconstruction objective, assuming statistical factor independence [60]. Several follow-ups refine latent prior matching through different objectives or prior choices [13, 38, 10, 59].
Disentanglement through a Geometric Lens.
Recent studies [26, 12, 41, 51, 31] indicate that disentanglement can arise by encouraging learned representations to preserve meaningful geometric features of the data, such as scaled distances between samples. Notably, Horan et al. [29] demonstrated that disentanglement is provably feasible when the generative factors are sufficiently non-Gaussian and locally isometric to the data. In this work, we explore how to promote geometry preservation using quadratic OT between the latent and data spaces, which we introduce in the next section.
2.2 Quadratic Optimal Transport
Gromov-{Monge, Wasserstein} Formulations.
OT [55] involves transferring one probability distribution to another while incorporating inductive biases. When these distributions lie on incomparable domains, the task is addressed using the Gromov-Monge and GW problems, also known as OT quadratic formulations. Formally, consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the Gromov-Monge problem [50] seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
$\inf_{T:T\sharp p=q}\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))|^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})\,.$ (GMP)
When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. However, solving this problem is difficult, and existence is not guaranteed in general [19]. Moreover, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty. Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e. distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the Gromov-Wasserstein (GW) problem [48, 64]
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
A solution $\pi^{\star}$ of (GWP) always exists, making $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q)$ a well-defined quantity. It quantifies the minimal distortion of the geometries induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ achievable when coupling $p$ and $q$.
Discrete Solvers.
When both $p$ and $q$ are instantiated as samples, the GW Prob. ((GWP)) translates to a quadratic assignment problem, whose objective can be regularized using entropy [15]. For empirical measures ${p}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$, ${q}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{\mathbf{y}_{j}}$ and $\varepsilon\geq 0$, we set:
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({p}_{n},{q}_{n})\vcentcolon=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,$ (EGWP)
where $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})$. As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. In addition to yielding better statistical [73] and regularity [58] properties, entropic regularization also enhances computational performance. In practice, we can solve (EGWP) using a mirror descent scheme that iterates the Sinkhorn algorithm [54, 62].
Neural Solvers.
While for classical OT, numerous neural methods have been proposed [45, 37, 22, 68, 66], the GW setting has received less attention. To our knowledge, the only neural Gromov-Monge formulation proposed thus far is [52], which involves a min-max-min optimization procedure. On the other hand, Klein et al. [36] recently proposed an approach to compute neural GW couplings.
3 Disentanglement with the Gromov-Monge gap
This section details our novel optimal transport perspective to achieve disentanglement from geometric considerations (see § 2.1), using the VAE framework. To achieve this, we first investigate how one can promote an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains $\mathcal{X}$ and $\mathcal{Y}$ to preserve predefined geometric features. In a VAE, $T$ can represent either the encoder $e_{\phi}$, which produces latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from the latent codes. As a result, in the former, the source domain $\mathcal{X}$ is the data, and the target domain $\mathcal{Y}$ is the latent space, with roles swapped in the latter. If we assume that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, it is equivalent whether $e_{\phi}$ preserves the geometric features from data to latents or $d_{\theta}$ preserves them from latents to data. Consequently, in what follows, $T$ can refer to either the encoder or the decoder without distinction.
Outline.
Leveraging this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1. We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves points while preserving geometric features as much as possible, i.e., minimizing distortion. §3.3 then shows how the GMG can be estimated and computed from samples to be practically applicable in the VAE framework, which transitions into §3.4 studying convexity properties of the GMG. Put together, §3.13.4 define the practical GMG which allows us to learn a latent space that matches, as much as possible, geometrical constraints in the data space. Finally in §3.5, we use the GMG with different choices of costs to propose effective disentangled representation learning objectives.
3.1 From the distortion…
We encode the geometric features of interest through two cost functions defined on each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. We then want $T$ to preserve these costs, i.e., $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$. In practice, two types of cost functions are particularly meaningful.
  • [i]
    (Scaled) squared Euclidean distance: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$ and $c_{\mathcal{Y}}=\alpha^{2}\|\mathbf{y}-\mathbf{y}\|_{2}^{2}$, with $\alpha>0$. A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ preserves the scaled distances between the points, i.e. it is a scaled isometry. When $\alpha=1$, we recover the standard definition of an isometry.
  • [ii]
[…]
In the following, we say that $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] if they belong to these families of costs. Introducing a reference distribution $r\in\mathcal{P}(\mathcal{X})$, weighting the areas of $\mathcal{X}$ where we penalize deviations of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ from $c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$, we can quantify this property using the following criterion.
Definition 3.1 (Distortion).
The distortion (DST) of a map $T$ is defined as:
$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T):=\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})$ (DST)
$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies how much $T$ distorts the geometric features induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ on the support of $r$, i.e., when $\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$. In disentangled representation learning, it can be desirable to regularize the decoder to be isometric [29, 51, 31]. However, a fully geometry-preserving mapping might not necessarily exist between the latent space and the data distribution. This means there usually exists an inherent trade-off between the accurate reconstruction of the data distribution and this reconstruction being, e.g., a "fully isometric" map. If such a map does not exist, the reconstruction loss and the distortion term cannot be 00 simultaneously. In practice, this means that the distortion loss will move away from accurately reconstructing the data, which negatively impacts the quality of the learned latent representations. This naturally raises the question of how to formulate a geometric regularization that takes the most geometry-preserving mapping into account.
[…]
3.2 …to The Gromov-Monge Gap
Recently, Uscidda and Cuturi [68] introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost. In practice, this regularizer is combined with fitting losses to compute Monge maps, which are defined by two main features: (i) they fit a marginal constraint with (ii) minimal displacement cost. Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost. In § 3.5, we use it, alongside fitting losses, to compute Gromov-Monge maps, as defined in Eq. (GMP), which are similarly defined by (i) fitting a marginal constraint with (ii) minimal distortion cost.
Definition 3.2 (Gromov-Monge gap).
The Gromov-Monge gap (GMG) of a map $T$ is defined as:
$\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ (GMG)
From Eq. (GWP), we recall that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ is the minimal distortion achievable when transporting $r$ to $T\sharp r$. Thus, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies the difference between the distortion incurred when transporting $r$ to $T\sharp r$ via $T$, to this minimal distortion. More formally, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is the optimality gap of $T$ in the Gromov-Monge Prob. (GMP) between $r$ and $T\sharp r$, which is always feasible, even when $r$ is discrete, as $T$ belongs to the constraint set. In light of this, it is a well-defined and
  • The GMG measures how close $T$ is to be a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. Indeed, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0$ with equality i.f.f. $T$ is a Gromov-Monge map solution of Prob. (GMP) between $r$ and $T\sharp r$, i.e., $T$ moves $r$ with minimal (but eventually non zero) distortion.
  • When transport without distortion is possible, the GMG coincides with the distortion. When there exists another map $U:\mathcal{X}\to\mathcal{Y}$ transporting $r$ to $T\sharp r$ with zero distortion, i.e., $U\sharp r=T\sharp r$ and $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(U)=0$, then $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$. Indeed, $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)=0$ in that case, as the coupling $\pi=(\mathrm{Id},U)\sharp r$ sets the GW objective to zero, thereby minimizing it.
The last point (ii) is fundamental and illustrates how the GMG functions as a debiased distortion. Indeed, it compares the distortion induced by $T$ to a baseline distortion, defined as the minimal achievable distortion when transforming the reference distribution into its image under $T$. Thus, when transformation without any distortion is achievable, the reference distortion becomes zero, and the GMG precisely aligns with the distortion itself, i.e., $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$. In this context, the GMG offers the optimal compromise: it avoids the over-penalization induced by the distortion when fully preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ is not feasible, yet it coincides with it when such preservation is feasible.
The Influence of the Reference Distribution.
A crucial property of $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$. Formally, if $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\operatorname{supp}(s)\subseteq\operatorname{supp}(r)$, then $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$. This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion? We answer this question positively with Prop. (3.3) when the costs are the (scaled) Euclidean distances or the cosine similarity. Intuitively, this means that if $T$ moves $r$ while preserving (scaled) distances or angles as much as possible, it will also preserve these properties as much as possible when moving any "smaller" distribution within $r$.
Proposition 3.3.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] (see § 3.1), if $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$, then for any $s\in\mathcal{P}(\mathcal{X})$ s.t. $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, one has $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$.
3.3 Estimation and Computation from Samples
Plug-In Estimation.
In practice, we estimate Eq. (DST) and Eq. (GMG) using i.i.d. samples $\mathbf{x}_{1},...,\mathbf{x}_{n}$ from the reference distribution $r$. We then consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use a plug-in estimator for both cases, i.e., we estimate the distortion via
$\begin{split}\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (2)
and the GMG via $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$. To better understand what the discrete GMG quantifies, we can reformulate it using the minimal distortion achieved by a permutation $\sigma\in\mathcal{S}_{n}$ between the $\mathbf{x}_{i}$ and the $T(\mathbf{x}_{i})$.
Proposition 3.4.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] , the empirical GMG reads:
$\begin{split}\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\sigma\in\mathcal{S}_{n}}\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\left(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{\sigma(i)}),T(\mathbf{x}_{\sigma(j)}))\right)^{2}\end{split}$ (3)
As a Monte Carlo estimator, $\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is naturally consistent. We can ask the same question for $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)$, which requires studying the convergence of the empirical GW distance $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$. For the costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ of interest, we show that consistency holds.
Proposition 3.5.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii], $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\rightarrow\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ almost surely.
Efficient Computation.
Computing $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ requires solving a discrete GW problem between $r_{n}$ and $T\sharp r_{n}$ to obtain $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$. To alleviate computational challenges, we estimate this term using an entropic regularization $\varepsilon\geq 0$, as introduced in Eq. (EGWP):
$\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})\,.$ (4)
Choosing $\varepsilon=0$, we recover the unregularized one $\mathcal{GM}_{{r}_{n},0}^{c_{\mathcal{X}},c_{\mathcal{Y}}}=\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. Moreover, the entropic estimator preserves the positivity, as for $\varepsilon\geq 0$, we have $\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}\geq 0$ (see A.1). As described in § 2, we compute $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al. [54]’s solver. While it always has $\mathcal{O}(n^{2})$ memory complexity, when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, this solver runs in $\mathcal{O}(n^{2}d)$ time [62, Alg. 2]. Since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, the computation of the GMG for the costs of interest [i] or [ii] scales as $\mathcal{O}(n^{2}d)$ in time. We use ott-jax’s [16] implementation of this scheme.
3.4 (Weak) Convexity of the Gromov-Monge gap
As laid out, the GMG can be used as a regularization loss to push any model $T$ to be more geometry-preserving. A natural question that arises when defining such a regularizer is: what are its regularity properties, and in particular, is it convex? In the following, we study the convexity of $T\mapsto\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$, and its finite-sample counterpart $T\mapsto\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$. We focus on the costs of interest [i] or [ii]. For simplicity, we replace cosine similarity with inner product—i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$—as they are equivalent, up to normalization of $r$ and $T$. We then study the convexity of the GMG for (i) the (scaled) squared Euclidean distances and (ii) the inner product, denoted respectively by (i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$. To that end, we introduce a weaker notion of convexity, previously defined for functions on $\mathbb{R}^{d}$ [17], which we extend here to $L_{2}(r)=\{T\,|\,\|T\|_{L_{2}(r)}^{2}\vcentcolon=\int_{\mathcal{X}}\|T(\mathbf{x})\|_{2}^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})<+\infty\}$.
Definition 3.6 (Weak convexity.).
With $\gamma>0$, a functional $\mathcal{F}:L_{2}(r)\to\mathbb{R}$ is $\gamma$-weakly convex if $\mathcal{F}_{\gamma}:T\mapsto\mathcal{F}(T)+\tfrac{\gamma}{2}\|T\|_{L_{2}(r)}^{2}$ is convex.
A weakly convex functional is convex up to an additive quadratic perturbation. The weak convexity constant $\gamma$ quantifies the magnitude of this perturbation and indicates a degree of non-convexity of $\mathcal{F}$. A lower $\gamma$ suggests that $\mathcal{F}$ is closer to being convex, while a higher $\gamma$ indicates greater non-convexity.
Theorem 3.7.
Both $\mathcal{GM}_{r}^{2}$ and $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, as well as their finite sample versions, are weakly convex.
  • Finite sample. We note $\mathbf{X}\in\mathbb{R}^{n\times d}$ the matrix that stores the $\mathbf{x}_{i}$, i.e. the support of $r_{n}$, as rows. Then, (i) $\mathcal{GM}_{r_{n}}^{2}$ and (ii) $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2,n}$ and (ii) $\gamma_{\textrm{inner},n}$-weakly convex, where: $\gamma_{\textrm{inner},n}=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})$ and $\gamma_{2,n}=\gamma_{\textrm{inner},n}+\max_{i=1\dots n}\|\mathbf{x}_{i}\|_{2}^{2}$.
  • Asymptotic. (i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2}$ and (ii) $\gamma_{\textrm{inner}}$-weakly convex, where: $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ and $\gamma_{2,n}=\gamma_{\textrm{inner}}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
From a practitioner’s perspective, we analyze the insights provided by Thm. (3.7) in three parts.
  • First, we have $\gamma_{2}\geq\gamma_{\textrm{inner}}$. Therefore, $\mathcal{GM}_{r}^{2}$ is less convex than $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, making it harder to optimize, and the same argument holds for their estimator. In other words, we provably recover that, in practice, preserving the (scaled) distances is harder than simply preserving the angles.
  • Second, as $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])\geq\lambda_{\max}(\mathrm{Cov}_{\mathbf{x}\sim r}[\mathbf{x}])$, this exhibits a tradeoff w.r.t. Prop. (3.3): by choosing a bigger reference distribution $r$, we trade the convexity of the GMG. For $\gamma_{2}$, the dependency in $r$ is even worse. In practice, we then choose $r$ with support as small as possible, precisely where we want $T$ to move points with minimal distortion.
  • Third, and probably the most surprising, the finite sample GMG is more convex in high dimension. Indeed, $\gamma_{\textrm{inner},n}$ is the spectral width of $\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top}$, containing the (rescaled) inner-products between the $\mathbf{x}_{i}\sim r$. When $n>d$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})=0$ as $\mathrm{rank}(\mathbf{X}\mathbf{X}^{\top})=d$. Then, $\gamma_{\textrm{inner},n}$ increases, which in turn decreases the GMG’s convexity. On the other hand, when $d>n$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})>0$ if $\mathbf{X}$ is full rank. Intuitively, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is nearly convex when $\mathbf{X}\mathbf{X}^{\top}$ is well conditioned. Assuming that the $\mathbf{x}_{i}$ are normalized, this might happen in high dimension, as those points will be orthogonal with high probability. This property suggests that, in practice, and contrary to the insights provided by the statistical OT literature [71, 25, 56, 73], the GMG might not benefit a large sample size.
3.5 Learning with the Gromov-Monge gap
General Learning Procedure.
Provided a source distribution $p$ and a target $q$ defining a marginal constraint, learning with the GMG remains to optimize a loss of the form
$\mathcal{L}(\theta)\vcentcolon=\Delta(T_{\theta},p,q)+\lambda_{\textrm{GMG}}\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T_{\theta})$ (5)
where $\Delta$ is a fitting loss, which can access paired, or unpaired, samples of $p$ and $q$. In theory, from Prop. (3.3), we can choose any reference $r$ s.t. $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$. In practice, given the insights of Thm. (3.6), we usually consider $r=p$. Note that replacing $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$ by $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ in Eq. (5), we similarly define the learning procedure with the distortion. We compare their effect in Figure 1.
VAE Learning Procedure.
In the VAE setting, (i) when we apply the GMG (or the distortion) to the encoder $e_{\phi}$, the fitting loss is defined through the prior matching constraint, as described in § 2.1. Conversely, (ii) when we apply the GMG to the decoder $d_{\phi}$, the fitting loss is defined through the reconstruction loss. Additionally, in both cases, our goal is to promote the latent space to preserve certain geometric features of the data. Therefore, in (i) we use $r=p_{\textrm{data}}$ the data distribution as reference $r$, while in (ii) we use the latent distribution $r=q_{\phi}$. Introducing weightings $\lambda_{\textrm{enc}},\lambda_{\textrm{dec}}\geq 0$, determining which mapping we regularize, this remains to optimize the loss
$\mathcal{L}_{\beta\textrm{-GMG}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{GM}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{GM}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (6)
where $\mathcal{L}_{\beta}$ is introduce in § 2.1. Note that this loss can easily be extended to $\beta$-TCVAE and the combination of other regularization terms. While previous work [51, 41] chooses to apply the geometric regularizations to the decoder, we investigate regularizing both, separately or simultaneously. For completeness, we also derive the VAE-loss when learning with the distortion:
$\mathcal{L}_{\beta\textrm{-DST}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{D}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{D}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (7)
The choice of $c_{\mathcal{X}},c_{\mathcal{Y}}$.
[…]
4 Experiments
Table 1: Effect of different geometric regularization on disentanglement (DCI-D, Shapes3D [33]). We highlight the best method per regularization, and the best/second best per column.
$\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
Base $65.8$$\pm 15.6$ $75.0$$\pm 3.4$ $88.1$$\pm 7.4$ $90.2$$\pm 7.5$
Isometric (IR)
+ (DST) 71.5 ±$\pm 3.6$ 75.8 ±$\pm 6.6$ 92.1 ±$\pm 9.7$ 90.9 ±$\pm 7.6$
+ (GMG) 72.0 ±$\pm 12.5$ 78.9 ±$\pm 5.0$ 92.5 ±$\pm 4.4$ 91.7 ±$\pm 6.0$
Scaled Isometric (SIR)
+ Jacobian 61.4 ±$\pm 12.8$ 76.7 ±$\pm 4.5$ 90.5 ±$\pm 3.8$ 91.5 ±$\pm 5.6$
+ (DST) 67.4 ±$\pm 7.1$ 77.9 ±$\pm 4.5$ 93.2 ±$\pm 9.7$ 94.5 ±$\pm 6.9$
+ (GMG) 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
Conformal (CR)
+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
Experimental setup.
We evaluate the effectiveness of the (GMG) as regularizer in disentangled representation learning. We use the standard $\beta$-VAE and $\beta$-TCVAE as our base models and incorporate the GMG on top of them. Moreover, we consider the recently proposed HFS [60] regularization on top of both $\beta$-VAE and $\beta$-TCVAE, totaling four base models. Our primary goal is to investigate the differences between using the GMG and the (DST) as regularizers, specifically examining whether the (GMG) leads to more disentangled representations compared to the raw distortion. Additionally, we aim to determine which geometric regularization (IR, SIR, CR) is most beneficial for disentanglement and what part of the pipeline should be regularized. Lastly, we investigate whether a geometric regularization can help prevent the collapse of learned representation in the Decoder-free setting. We evaluate the learned latents with DCI-D [20] as it was found to be the metric most suitable to measure disentanglement [44, 18]. We benchmark over multiple datasets commonly used to assess the performances of disentangled representation learning methods: Shapes3D [33], DSprites [27], SmallNORB [40], and Cars3D [57].
4.1 Evaluating Different Geometric Regularizations
Regularizing the Decoder.
First, we focus on the difference between optimizing for different geometry-preserving regularizations. We compare between IR, SIR, and CR [41] realized through either the (DST), or (GMG). Additionally, we include the Jacobian-based SIR as introduced in Lee et al. [41]. We report full results on Shapes3D [9] over 5 seeds in Table 1. We observe that the GMG outperforms the sole distortion loss for all levels of regularization and baselines. Moreover, we find that a CR performs best with respect to disentanglement compared to both IR and SIR. Note, that employing a CR has not been benchmarked for disentangled representation learning before. These results elucidate the clear benefit of using the GMG in its CR implementation in terms of learning more disentangled representations significantly improving upon previously proposed regularization.
Table 2: Effect of (GMG) and (DST) leveraged as a conformal regularization (CR) on the disentanglement of learned representations as measured by DCI-D over four datasets. We highlight the best, and second best result for each dataset and method.
CR $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
Shapes3D [33]
Base $65.8$$\pm 15.6$ $75.0$$\pm 3.4$ $88.1$$\pm 7.4$ $90.2$$\pm 7.5$
+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
DSprites [27]
Base $26.2$$\pm 18.5$ ${32.3}$$\pm 19.3$ $33.6$$\pm 17.9$ $48.7$$\pm 10.2$
+ (DST) $\underline{28.6}$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
+ (GMG) 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
SmallNORB [40]
Base $26.8$$\pm 0.2$ $29.8$$\pm 0.4$ $26.8$$\pm 0.2$ $29.8$$\pm 0.4$
+ (DST) $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$ $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$
+ (GMG) 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$ 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$
Cars3D [57]
Base $29.6$$\pm 5.7$ $32.3$$\pm 4.6$ $29.6$$\pm 5.7$ $32.3$$\pm 4.6$
+ (DST) $26.8$$\pm 3.6$ $33.7$$\pm 4.2$ $26.8$$\pm 3.6$ $33.7$$\pm 4.2$
+ (GMG) 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
Thus, next we benchmark the GMG in its CR form against its distortion counterpart across three more datasets again over four different baselines. We report full results in Table 2. Again we observe that the GMG outperforms or performs equally well to its distortion equivalent, confirming the benefits of accounting for the optimal possible mapping in the regularization. Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it. We emphasize that using the GMG as CR significantly improves results for all datasets versus not using any isometric regularization. This establishes the GMG as an effective regularization method beneficial for disentangled representation learning.
Regularizing the Encoder.
Lastly, we also analyze a CR on $e_{\phi}$, as well as regularizing both $d_{\theta}$ and $e_{\phi}$ together. We report full results over two datasets in Table 4. Again, the GMG on $d_{\theta}$ achieves best DCI-D over all baselines. This result is expected in the light of Theorem 3.7. Interestingly, regularizing solely $d_{\theta}$ outperforms regularizing both $e_{\phi}$ and $d_{\theta}$. We hypothesize this is due to the regularization of the decoder also offering a stronger signal as its gradients impact both the decoder and the encoder, as in this case, the reference $r$ is the distribution of encoded images.
4.2 Towards Decoder-free Disentanglement
Recently, works such as [11, 70, 21, 47, 1] have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation
Table 3: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D [9].
Decoder-free $\beta$-VAE $\beta$-TCVAE Base $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ Isometric (IR) + (DST) $38.2$$\pm 0.8$ $42.7$$\pm 1.6$ + (GMG) $13.9$$\pm 0.4$ $20.5$$\pm 0.5$ Scaled Isometric (SIR) + (DST) 45.6 ±$\pm 1.2$ 53.5 ±$\pm 1.0$ + (GMG) $15.2$$\pm 0.3$ $25.2$$\pm 0.6$ Conformal (CR) + (DST) $37.0$$\pm 0.4$ $46.1$$\pm 1.5$ + (GMG) $37.0$$\pm 0.9$ $38.8$$\pm 1.1$
learning [14, 72, 3, 24]. However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
As the goal of geometry preservation connects the data manifold and the latent domain through a minimal distortion objective and is applicable to both the encoder and decoder of a VAE (§3, Table 4), we posit that its application may provide sufficient training signal to learn meaningful representations and encourage disentanglement, eliminating the need for a reconstruction loss and decoder. Table 3 shows preliminary results on unsupervised decoder-free disentangled representation learning on the Shapes3D benchmark, where the decoder and associated reconstruction objective have been removed.
Standard approaches such as $\beta$-VAE or $\beta$-TCVAE collapse and do not achieve measurable disentanglement (DCI-D of $0.0$). However, the inclusion of either DST or GMG significantly raises achievable disentanglement and, combined with the $\beta$-TCVAE matching objective, can achieve DCI-D scores of up to $53.5$ without needing any decoder or reconstruction loss. While these are preliminary insights, we believe they offer promise for more scalable approaches to unsupervised disentangled representation learning and potential bridges to popular and scalable self-supervised representation learning approaches. Note, that here the distortion loss significantly outperforms the GMG. This is expected due to the nature of the GMG, as the distortion loss offers a more restrictive and, thus, stronger signal for learning representations, which is necessary in the absence of a reconstruction objective. This highlights that while in most scenarios (§ 4.1, Figure 1), the GMG is preferable over the distortion loss, there also exist settings where a more restrictive optimization signal is desirable.
5 Conclusion
In this work, we introduce an optimal transport (OT) perspective on unsupervised disentangled representation learning to incorporate general latent geometrical constraints. We derive the Gromov-Monge gap (GMG), a provably weakly convex OT regularizer that measures the preservation of geometrical properties by a transport map $T$. By formulating disentangled representation learning as a transport problem, we integrate the GMG into standard training objectives, allowing for incorporating and studying various geometric constraints on the disentanglement of learned representation spaces. Including these geometry preserving regularization offers significant performance benefits across four standard disentanglement benchmarks when applied to existing disentanglement methods. Moreover, we show promising results on decoder-free unsupervised disentanglement. We demonstrate that optimizing for geometric constraints through the OT lens can provide sufficient training signal and regularization on the model encoder to achieve measurable disentanglement without explicit reconstruction objectives. This opens a possible door towards more scalable unsupervised disentanglement and bridges to weakly- & self-supervised encoder-only representation learning efforts.
Acknowledgements
Co-funded by the European Union (ERC, DeepCell - 101054957). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Luca Eyring and Karsten Roth thank the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. Karsten Roth also thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. Zeynep Akata was supported by BMBF FKZ: 01IS18039A, by the ERC (853489 - DEXIM), by EXC number 2064/1 – project number 390727645. Fabian J. Theis consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity.
Appendix
Appendix A Proofs
A.1 Positivity of the Entropic GMG
Recall that
$\begin{split}\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}(T)&\vcentcolon=\tfrac{1}{n}\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
For any coupling $\mathbf{P}\in U_{n}$, since $-\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
$\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})<\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
As a result, applying minimization on both sides yields that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})<\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}({r}_{n},T\sharp{r}_{n})$, and therefore:
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}(T)>\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}(T)=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0.$
A.2 Reminders on Monge and Kantorovich OT
In this section, we recall the Monge and Kantorovich formulations of OT, which we will use to prove various results. These are the classical formulations of OT. Although we introduce them here after discussing the Gromov-Monge and Gromov-Wasserstein formulations, it should be noted that they are generally introduced beforehand. Indeed, the Gromov-Monge and Gromov-Wasserstein formulations were historically developed to derive OT formulations for comparing measures supported on incomparable spaces.
Monge Formulation.
Instead of intra-domain cost functions, we consider here an inter-domain continuous cost function $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. This assumes that we have a meaningful way to compare elements $\mathbf{x},\mathbf{y}$ from the source and target domains. The Monge [49] problem (MP) between $p\in\mathcal{P}(\mathcal{X})$ and $p\in\mathcal{P}(\mathcal{Y})$ consists of finding a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $p$, while minimizing the average displacement cost quantified by $c$
$\inf_{T:T\sharp p=p}\int_{\mathcal{X}}c(\mathbf{x},T(\mathbf{x}))\mathop{}\!\mathrm{d}p(\mathbf{x})\,.$ (MP)
We call any solution $T^{\star}$ to this problem a Monge map between $p$ and $q$ for cost $c$. Similarly to the Gromov-Monge Problem (GMP), solving the Monge Problem (MP) is difficult, as the constraint set is not convex and might be empty, especially when $p,q$ are discrete.
Kantorovich Formulation.
Instead of transport maps, the Kantorovich problem (KP) seeks a couplings $\pi\in\Pi(p,q)$:
$\mathrm{W}_{c}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\,.$ (KP)
An optimal coupling $\pi^{\star}$ solution of (KP), always exists. Studying the equivalence between (MP) and (KP) is easier than in the Gromov-Monge and Gromov-Wasserstein cases. Indeed, when (MP) is feasible, the Monge and Kantorovich formulations coincide and $\pi^{\star}=(\mathrm{Id},T^{\star})\sharp p$.
A.3 Conditionally Positive Kernels
In this section, we recall the definition of a conditionally positive kernel, which is involved in multiple proofs relying on the linearization of the Gromov-Wasserstein problem as a Kantorovich problem.
Definition A.1.
A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is conditionally positive if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t. $\mathbf{a}^{\top}\mathbf{1}_{n}=0$, one has
$\sum_{i,j=1}^{n}\mathbf{a}_{i}\mathbf{a}_{j}\,k(\mathbf{x}_{i},\mathbf{x}_{j})\geq 0$
Conditionally positive kernels include all positive kernels, such as the inner-product $k(\mathbf{x},\mathbf{y})=\langle\mathbf{x},\mathbf{y}\rangle$, the cosine similarity $k(\mathbf{x},\mathbf{y})=\textrm{cos-sim}(\mathbf{x},\mathbf{y})=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{y}}{\|\mathbf{y}\|_{2}}\rangle$, but also the negative squared Euclidean distance $k(\mathbf{x},\mathbf{y})=-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. Therefore, each of the costs of interest is either a conditionally positive kernel - for the inner product and the cosine distance - or its opposite is - the squared Euclidean distance.
Appendix B Proofs of § 3.2
See 3.3
Proof.
Let $T,r,s$ as described and suppose that $\mathcal{GM}_{r}^{c}(T)=0$. Then, $\pi^{r}:=(\mathrm{Id},T)\sharp r$ is an optimal Gromov-Wasserstein coupling, solution of Problem (GWP) between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. Therefore, from [65, Theorem. 3], $\pi^{r}$ is an optimal Kantorvich coupling, solution of Problem (KP) between $r$ and $T\sharp r$ for the linearized cost:
$\tilde{c}:(\mathbf{x},\mathbf{y})\in\mathcal{X}\times\mathcal{Y}\mapsto\int_{\mathcal{X}\times\mathcal{Y}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi^{r}(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ (8)
Additionally, $\mathcal{X}\times\mathcal{Y}$ is a compact set as a product of compact sets, so since $(\mathbf{x},\mathbf{y})\mapsto|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$ is continuous as $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ are continuous, it is bounded on $\mathcal{X}\times\mathcal{Y}$. Afterward, since $\pi^{r}$ has finite mass, by Lebesgue’s dominated convergence Theorem, it follows that $\tilde{c}$ is continuous, and hence uniformly continuous, again since $\mathcal{X}\times\mathcal{Y}$ is compact.
Afterwards, by virtue of [61, Theorem 1.38], $\mathrm{Spt}\left(\pi^{r}\right)$ is a $\tilde{c}$-cyclically monotone (CM) set (see [61, Definition. 1.36]). From the definition of cyclical monotonicity, this property translates to subsets. Then, by defining $\pi^{s}=(\mathrm{Id},T)\sharp s$, as $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, one has $\mathrm{Spt}(\pi^{s})=\mathrm{Spt}((\mathrm{Id},T)\sharp s)\subset\mathrm{Spt}((\mathrm{Id},T)\sharp r)=\mathrm{Spt}(\pi^{r})$, so $\mathrm{Spt}(\pi^{s})$ is $\tilde{c}$-CM. Finally, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $\tilde{c}$ is uniformly continuous, the $\tilde{c}$-cyclical monotonicity of its support implies that the coupling $\pi^{p}$ is a Kantorovich optimal coupling between its marginals for cost $\tilde{c}$, thanks to [61, Theorem 1.49]. By re-applying [65, Theorem. 3], we get that $\pi^{s}$ solves the Gromov-Wasserstein problem between its marginals for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$. In other words, $\pi^{s}=(\mathrm{Id},T)\sharp s$ is Gromov-Wasserstein optimal coupling between $s$ and $T\sharp s$ so $T$ is a Gromov-Monge map between $s$ and $T\sharp s$ and $\mathcal{GM}_{s}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$. ∎
Appendix C Proofs of § 3.3
See 3.4
Proof.
We start by showing a more general results, stating that when $c_{\mathcal{X}},c_{\mathcal{Y}}$ are conditionally positive kernels (see A.1), the discrete GW couplings between uniform, empirical distributions supported on the same number of points, ae permutation matrices.
Proposition C.1 (Equivalence between Gromov-Monge and Gromov-Wasserstein problems in the discrete case.).
Let ${p}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and ${q}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{y}_{i}}$ two uniform, empirical measures, supported on the same number of points. We denote by $P_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n},\exists\sigma\in\mathcal{S}_{n},\mathbf{P}_{ij}:=\delta_{j,\sigma(i)}\}$ the set set of permutation matrices. Assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels (see A.1). Then, the GM and GW formulations coincide, in the sense that we can restrict the GW problem to permutations, namely
$\begin{split}\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({p}_{n},{p}_{n})&=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\sigma\in\mathcal{S}_{n}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{\sigma(i)},\mathbf{y}_{\sigma(j)}))^{2}\end{split}$ (9)
Proof.
Let $\mathbf{P}^{\star}\in U_{n}$ solution of the Gromov-Wasserstein between ${p}_{n}$ and ${p}_{n}$, i.e.
$\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
that always exists by continuity of the GW objective function on the compact $U_{n}$. We show that $\mathbf{P}^{\star}$ can be chosen as a (rescaled) permutation matrix without loss of generality.
As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from [65, Theorem. 3], $\mathbf{P}^{\star}$ also solves:
$\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{Q}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}^{\star}_{ij}\mathbf{Q}_{i^{\prime}j^{\prime}}$ (10)
We then define the linearized cost matrix $\tilde{C}\in\mathbb{R}^{n\times n}$, s.t.
$\tilde{\mathbf{C}}_{ij}=\sum_{i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}^{\star}_{ij}$
which allows us to reformulate Eq. (10) as
$\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{Q}\in U_{n}}\langle\tilde{\mathbf{C}},\mathbf{Q}\rangle$ (11)
Birkhoff’s theorem states that the extremal points of $U_{n}$ are equal to the permutation matrices $P_{n}$. Moreover, a seminal theorem of linear programming [6, Theorem 2.7] states that the minimum of a linear objective on a bounded polytope, if finite, is reached at an extremal point of the polyhedron. Therefore, as $\mathbf{P}^{\star}$ solves Eq. (11), it is an extremal point of $U_{n}$, so it can always be chosen as a permutation matrix. Therefore, the equivalence between GW and GM follows. ∎
To conclude the proof of Prop. 3.4, we simply remark that:
  • $r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $T\sharp r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$ are uniform, empirical distribution, and supported on the same number of points;
  • The costs of interests [i] or [ii] are either conditionally positive, or their opposite is, as detailed below  Def (A.1).
See 3.5
Proof.
We first note that the empirical estimator of the distortion is consistent, as both costs [i] or [ii] are continuous, and $\mathcal{X}$ is compact. We then need to study, in both cases, the convergence of $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r_{n})$ to $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r)$.
To that end, we first remark that as, almost surely, $r_{n}\to r$ in distribution, one also has that, almost surely, $T\sharp r_{n}\to T\sharp r$ in distribution. Indeed, since $\mathcal{Y}$ is compact, $T$ is bounded so for any bounded and continuous $f:\mathcal{Y}\rightarrow\mathbb{R}$ and $X\sim r$, $f\circ T(X)$ is well defined and bounded so integrable. Afterwards, one can simply adapt the proof of the almost sure weak convergence of empirical measure based on the strong law of large numbers to show that, almost surely, $T\sharp{r}_{n}\rightarrow T\sharp r$ in distribution. See for instance [39, Theorem 10.4.1].
[i] We start by the (scaled) squared euclidean distances. Up to replacing $r$ by $\alpha^{2}\sharp r$ and $T$ by $T\circ(\tfrac{1}{\alpha^{2}})$, and similarly for $r_{n}$, we can assume without loss of generality that $\alpha=1$. As, almost surely, both $r_{n}\to r$ and $T\sharp{r}_{n}\to T\sharp r$ in distribution, the results follows from [48, Thm 5.1, (e)].
[ii] We continue with the cosine similarity. To that end, we first consider the inner product, i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, and show that if $p_{n}\to p$ and $q_{n}\to q$ in distribution, then $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$. As noticed by Rioux et al. [58, Lemma 2]–in the first version of the paper– the GW for inner product costs can be reformulated as:
$\begin{split}\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)=&\int_{\mathcal{X}\times\mathcal{X}}\langle\mathbf{x},\mathbf{x}^{\prime}\rangle\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})+\int_{\mathcal{Y}\times\mathcal{Y}}\langle\mathbf{y},\mathbf{y}^{\prime}\rangle\mathop{}\!\mathrm{d}q(\mathbf{y})\mathop{}\!\mathrm{d}q(\mathbf{y}^{\prime})\\ +&\min_{\mathbf{M}\in\mathcal{M}}\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+4\|\mathbf{M}\|_{2}^{2}\,,\end{split}$ (12)
where we define $\mathcal{M}=[-M/2,M/2]^{d_{\mathcal{X}}\times d_{\mathcal{Y}}}$ with $M=\sqrt{\int_{\mathcal{X}}\|\mathbf{x}\|_{2}^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})\int_{\mathcal{Y}}\|\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}q(\mathbf{y})}$. In particular, they show this result for the entropic GW problem with $\varepsilon>0$, but their proof is also valid for $\varepsilon=0$. The above terms only involving the marginal, i.e., not involved in the minimization, are naturally stable under convergence in distribution, as $\mathcal{X}$ and $\mathcal{Y}$ are compact, so as $\mathcal{X}\times\mathcal{X}$ and $\mathcal{Y}\times\mathcal{Y}$. As a result, we only need to study the stability of this quantity under the convergence in distribution of the following functional:
$\mathcal{F}(p,q)=\min_{\mathbf{M}\in\mathcal{M}}\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+4\|\mathbf{M}\|_{2}^{2}\,,$ (13)
We first remark that:
$\begin{split}&|\mathcal{F}(p,q)-\mathcal{F}(p_{n},q_{n})|\\ \leq&\sup_{M\in\mathcal{M}}|\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})-\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})|\\ \leq&\sup_{M\in\mathcal{M}}|\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}2\|\mathbf{M}\mathbf{x}-\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})-\min_{\pi\in\Pi(p_{n},q_{n})}\int_{\mathcal{X}\times\mathcal{Y}}2\|\mathbf{M}\mathbf{x}-\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})2|\\ +&2\cdot\,\sup_{M\in\mathcal{M}}|\int_{\mathcal{X}}\|\mathbf{M}\mathbf{x}\|_{2}^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})-\int_{\mathcal{X}}\|\mathbf{M}\mathbf{x}\|_{2}^{2}\mathop{}\!\mathrm{d}p_{n}(\mathbf{x})|\\ +&2\cdot\,|\int_{\mathcal{Y}}\|\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}q(\mathbf{y})-\int_{\mathcal{Y}}\|\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}q_{n}(\mathbf{y})|\end{split}$ (14)
Then, we show the convergence of each term separately.
  • For the first term, we remark that (up to a constant factor) it can be reformulated:
    $\sup_{M\in\mathcal{M}}|\mathrm{W}_{2}^{2}(\mathbf{M}\sharp p,q)-\mathrm{W}_{2}^{2}(\mathbf{M}\sharp p_{n},q_{n})|$
    where we remind that that $\mathrm{W}_{2}^{2}$ is the (squared) Wasserstein distance, solution of Eq. (KP) induced by $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. By virtue of [46, Theorem 2], there exists a constant $C>0$, s.t. we can uniformly bound
    $\sup_{M\in\mathcal{M}}|\mathrm{W}_{2}^{2}(\mathbf{M}\sharp p,q)-\mathrm{W}_{2}^{2}(\mathbf{M}\sharp p_{n},q_{n})|\leq Cn^{-1/d}$
    and the convergence follows.
  • For the second one, this follows from from the convergence in distribution of $p_{n}$ to $p$ along with the Ascoli-Arzela theorem, since both $\mathcal{M}$ and $\mathcal{X}$ are compact sets, so the $\{f_{\mathbf{M}}\,|\,f_{\mathbf{M}}:\mathbf{x}\mapsto\|\mathbf{M}\mathbf{x}\|_{2}^{2}\}$ are uniformly bounded and equi-continuous.
  • For the third one, this follows from the convergence in distribution of $q_{n}$ to $q$.
As a result, we finally get $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$. Finally, we remark that for any $p,q$, $\mathrm{GW}^{\textrm{cos-sim}}(p,q)=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)$, where $\mathrm{proj}_{S^{d-1}}(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$. Using similar arguments invoked previously, as $p_{n}\to p$ in distribution $p$, $\mathrm{proj}_{S^{d-1}}\sharp p_{n}\to\mathrm{proj}_{S^{d-1}}\sharp p$ in distribution, and similarly $\mathrm{proj}_{S^{d-1}}\sharp q_{n}\to\mathrm{proj}_{S^{d-1}}\sharp q$ in distribution. As a result:
$\begin{split}\mathrm{GW}^{\textrm{cos-sim}}(p_{n},q_{n})&=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p_{n},\mathrm{proj}_{S^{d-1}}\sharp q_{n})\\ &\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)\\ &=\mathrm{GW}^{\textrm{cos-sim}}(p,q)\end{split}$ (15)
which yields the desired convergence by using $p_{n}=r_{n}$ and $q_{n}=T\sharp r_{n}$.
Appendix D Proofs of § 3.4
See 3.7
We start by recalling the standard definition of weakly convex function on $\mathbb{R}^{d}$, along with technical lemmas that we will in the proof of Thm. (3.7).
Definition D.1.
A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\gamma$-weakly convex if $f+\gamma\|\cdot\|_{2}^{2}$ is convex.
Lemma D.2.
Let $\mathbf{A}\in S_{d}(\mathbb{R})$ a symmetric matrix and define the quadratic form $f_{\mathbf{A}}:\mathbf{x}\in\mathbb{R}^{d}\mapsto\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$. Then, $f_{\mathbf{A}}$ is $\max(0,-\lambda_{\min}(\mathbf{A}))$-weakly convex.
Proof.
We use the fact that a twice continuously differentiable function is convex i.f.f. its hessian is positive semi-definite [8, §(3.1.4)]. Therefore, $f_{\mathbf{A}}$ is convex i.f.f. $\nabla^{2}f_{\mathbf{A}}=\mathbf{A}\geq 0$. If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e. 00-weakly convex. Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $-\lambda_{\min}(\mathbf{A})$-weakly convex. ∎
Lemma D.3.
Let $(f_{i})_{i\in I}$ a family of $\gamma$-weakly convex functions, with potentially infinite $I$. Then, $f:\mathbf{x}\in\mathbb{R}^{d}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})$ is $\gamma$-weakly convex.
Proof.
As the $f_{i}$ are $\gamma$-weakly convex, $f_{i}+\tfrac{1}{2}\gamma$ is convex, so $\mathbf{x}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})+\tfrac{1}{2}\gamma\|x\|_{2}^{2}=(\sup_{i\in I}f_{i}(\mathbf{x}))+\tfrac{1}{2}\gamma\|x\|_{2}^{2}$ is convex [8, Eq. (3.7)]. Therefore, the $\gamma$-weak convexity of $f$ follows ∎
Proof of Thm. (3.7)..
Finite sample. We first study the weak convexity of $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$, i.e. the Gromov-Monge gap for the inner product. For a map $T\in L_{2}(r)$, it reads
$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
As ${r}_{n}$ and $T\sharp{r}_{n}$ are uniform empirical supported on the same number of points, using Prop. C.1, we can reformulate the RHS with permutation matrices, which yields
$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
From this expression, $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ can be reformulated as a matrix input function. Indeed, it only depends on the map $T$ via its values on the support of ${r}_{n}$, namely $\mathbf{x}_{1},...,\mathbf{x}_{n}$. Therefore, we write $\mathbf{t}_{i}\vcentcolon=T(\mathbf{x}_{i})$, and define $\mathbf{X},\mathbf{T}\in\mathbb{R}^{n\times d}$ which contain observations $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ respectively, stored as rows. Then, studying $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ remains to study
$f(\mathbf{T})\vcentcolon=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle\mathbf{t}_{i},\mathbf{t}_{j}\rangle|^{2}-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle\mathbf{t}_{j},\mathbf{t}_{j^{\prime}}\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
By developing each term and exploiting that for any $\mathbf{P}\in P_{n}$, $\mathbf{P}\mathrm{1}_{n}=\mathbf{P}^{\top}\mathrm{1}_{n}=\tfrac{1}{n}\mathrm{1}_{n}$, we derive
$\begin{split}f(\mathbf{T})&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}-\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle\cdot\langle\mathbf{t}_{i},\mathbf{t}_{j}\rangle-\min_{\mathbf{P}\in P_{n}}\tfrac{1}{n^{2}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}-\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle\cdot\langle\mathbf{t}_{j},\mathbf{t}_{j^{\prime}}\rangle\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\max_{\mathbf{P}\in P_{n}}\tfrac{1}{n^{2}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle\cdot\langle\mathbf{t}_{j},\mathbf{t}_{j^{\prime}}\rangle\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle\cdot\langle\mathbf{t}_{i},\mathbf{t}_{j}\rangle\\ &=\max_{\mathbf{P}\in P_{n}}\langle\tfrac{1}{n^{2}}\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P},\mathbf{T}\mathbf{T}^{\top}\rangle-\langle\tfrac{1}{n^{2}}\mathbf{X}\mathbf{X}^{\top},\mathbf{T}\mathbf{T}^{\top}\rangle\\ &=\max_{\mathbf{P}\in P_{n}}\langle\tfrac{1}{n^{2}}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top}),\mathbf{T}\mathbf{T}^{\top}\rangle\\ &=\max_{\mathbf{P}\in P_{n}}\langle\tfrac{1}{n^{2}}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top})\mathbf{T},\mathbf{T}\rangle\\ &=\max_{\mathbf{P}\in P_{n}}\langle\mathbf{A}_{\mathbf{X},\mathbf{P}}\mathbf{T},\mathbf{T}\rangle\end{split}$
where we define $\mathbf{A}_{\mathbf{X},\mathbf{P}}\vcentcolon=\tfrac{1}{n^{2}}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top})\in\mathbb{R}^{n\times n}$. To study the convexity of this matrix input function, we vectorize it. From [53, Eq. (520)], we note that, for any $\mathbf{M}\in\mathbb{R}^{n\times n}$
$\begin{split}\langle\mathbf{M}\mathbf{T},\mathbf{T}\rangle=\mathbf{vec}(\mathbf{T})^{\top}\mathbf{vec}(\mathbf{M}\mathbf{T})=\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{M}\otimes I_{n})\mathbf{vec}(\mathbf{T})\end{split}$
where $\mathbf{vec}$ is the vectorization operator, raveling a matrix along its rows, and $\otimes$ is the Kronecker product. Applying this identity, we reformulate:
$f(\mathbf{T})=\max_{\mathbf{P}\in U_{n}}\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})\mathbf{vec}(\mathbf{T})$ (16)
To study the convexity of $r$, we study the convexity of each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}(\mathbf{T})\vcentcolon=\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})\mathbf{vec}(\mathbf{T})$, which are quadratic forms induced by the $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. This remains to study the (semi-) positive definiteness of the matrices $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. As each $\mathbf{A}_{\mathbf{X},\mathbf{P}}\in\mathbb{R}^{n\times n}$ is symmetric and square, $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is also symmetric and from [53, Eq. (519)] its eigenvalues are the outer products of the eigenvalues of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$ and $I_{n}$, namely
$\begin{split}\mathbf{eig}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})&=\{\lambda_{i}(\mathbf{A}_{\mathbf{X},\mathbf{P}})\cdot\lambda_{j}(I_{n})\}_{1\leq i,j\leq n}\\ &=\{\underbrace{\lambda_{1}(\mathbf{A}_{\mathbf{X},\mathbf{P}}),\dots,\lambda_{1}(\mathbf{A}_{\mathbf{X},\mathbf{P}})}_{\textrm{n times}},\dots,\underbrace{\lambda_{n}(\mathbf{A}_{\mathbf{X},\mathbf{P}}),\dots,\lambda_{n}(\mathbf{A}_{\mathbf{X},\mathbf{P}})}_{\textrm{n times}}\}\end{split}$ (17)
It follows that the minimal eigenvalue of $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is $\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})=\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}})$. Utilizing the expression of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$
$\begin{split}\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}})&=\tfrac{1}{n^{2}}\lambda_{\min}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top})\\ &\geq\tfrac{1}{n^{2}}(\lambda_{\min}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P})+\lambda_{\min}(-\mathbf{X}\mathbf{X}^{\top}))\\ &=\tfrac{1}{n^{2}}(\lambda_{\min}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P})-\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top}))\end{split}$ (18)
Reminding that $\mathbf{P}\in U_{n}$, one has $\mathbf{P}^{\top}=\mathbf{P}^{-1}$, so $\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}$ and $\mathbf{X}\mathbf{X}^{\top}$ are similar, and they have the same eigenvalues. In particular $\lambda_{\min}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P})=\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})$. Combining these results, it follows that
$\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})=\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}})\geq\tfrac{1}{n^{2}}(\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top}))$ (19)
We then remind that each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}$ is the quadratic form defined by $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$, so by applying Prop. D.2, it is $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$-weakly convex, and hence $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex. Therefore, applying Prop. (D.3), $r$ is $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex, in $\mathbb{R}^{d}$. Reminding that $\gamma_{\mathrm{inner}}=\tfrac{1}{n}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$, $r$ is $\tfrac{1}{n}\gamma_{\mathrm{inner}}$ weakly convex. This implies that $\mathbf{T}\mapsto f(\mathbf{T})+\tfrac{1}{n}\gamma_{\mathrm{inner}}\|\mathbf{T}\|_{2}^{2}$ is convex. By reminding that $\mathbf{T}$ stores the $T(\mathbf{x}_{i})$ as rows, $\tfrac{1}{n}\|\mathbf{T}\|_{2}^{2}=\|T\|_{L_{2}(r_{n})}$. Consequently, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is $\gamma_{\mathrm{inner}}$ in $L_{2}(r_{n})$.
We then study the convexity of $\mathcal{GM}_{r_{n}}^{2}$. We follow exactly the same approach. One has:
$\begin{split}\mathcal{GM}^{2}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
Similarly, studying the convexity of $\mathcal{GM}^{2}_{{r}_{n}}(T)$ remains to study the convexity of the matrix input function:
$\begin{split}g(\mathbf{T})&\vcentcolon=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|\mathbf{t}_{i}-\mathbf{t}_{j}\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|\mathbf{t}_{i}-\mathbf{t}_{j}\|_{2}^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
As before, by developing each term, one has:
$\begin{split}g(\mathbf{T})=\max_{\mathbf{P}\in P_{n}}&\tfrac{1}{n^{2}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle\cdot\langle\mathbf{t}_{j},\mathbf{t}_{j^{\prime}}\rangle\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}+\tfrac{1}{2n}\sum_{i,j=1}^{n}\mathbf{P}_{ij}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}\\ -&\left(\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle\cdot\langle\mathbf{t}_{i},\mathbf{t}_{j}\rangle+\tfrac{1}{2n}\sum_{i,j=1}^{n}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}\right)\end{split}$
The quadratic terms in $\mathbf{P}$ can be factorized as before using $\mathbf{A}_{\mathbf{X},\mathbf{P}}$. For the new terms w.r.t. the inner product case, we introduce $\mathbf{D}_{\mathbf{X}}\vcentcolon=\mathrm{diag}(\|\mathbf{x}_{1}\|_{2}^{2},\dots,\|\mathbf{x}_{n}\|_{2}^{2})$, and remark that we can rewrite:
$\begin{split}\tfrac{1}{2n}\sum_{i,j=1}^{n}\mathbf{P}_{ij}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}-\tfrac{1}{2n}\sum_{i,j=1}^{n}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}=\mathbf{vec}(T)^{\top}\left(\tfrac{1}{2n}(\mathbf{P}^{\top}-I_{n})\otimes\mathbf{D}_{\mathbf{X}}\right)\mathbf{vec}(T)\end{split}$
As we can always symetrize the matrix when considering its associated quadratic form, we have:
$\begin{split}\tfrac{1}{2n}\sum_{i,j=1}^{n}\mathbf{P}_{ij}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}-\tfrac{1}{2n}\sum_{i,j=1}^{n}\|\mathbf{x}_{i}\|_{2}^{2}\|\mathbf{t}_{i}\|_{2}^{2}=\mathbf{vec}(T)^{\top}\left(\tfrac{1}{2}(\tfrac{1}{2n}(\mathbf{P}^{\top}+\mathbf{P})-I_{n})\otimes\mathbf{D}_{\mathbf{X}}\right)\mathbf{vec}(T)\end{split}$
As a result, we denote $\mathbf{B}_{\mathbf{X},\mathbf{P}}=\tfrac{1}{n}(\tfrac{1}{2}(\mathbf{P}^{\top}+\mathbf{P})-I_{n})\otimes\mathbf{D}_{\mathbf{X}}$ and finally get:
$\begin{split}g(\mathbf{T})&=\max_{\mathbf{P}\in P_{n}}\mathbf{vec}(T)^{\top}\left(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}+\mathbf{B}_{\mathbf{X},\mathbf{P}}\right)\mathbf{vec}(T)\end{split}$
As we did for $f$, studying the weak convexity of $f$ remains to lower bound the minimal eigenvalue of $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}+\mathbf{B}_{\mathbf{X},\mathbf{P}}$. First, one remark that:
$\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}+\mathbf{B}_{\mathbf{X},\mathbf{P}})\geq\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})+\lambda_{\min}(\mathbf{B}_{\mathbf{X},\mathbf{P}})$
As we we have already lower bounded $\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})\geq\tfrac{1}{n^{2}}(\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top}))$, we focus on the RHS. Similarly, one has:
$\begin{split}\lambda_{\min}(\mathbf{B}_{\mathbf{X},\mathbf{P}})&=\lambda_{\min}\left(\tfrac{1}{2n}(\tfrac{1}{2}(\mathbf{P}^{\top}+\mathbf{P})-I_{n})\otimes\mathbf{D}_{\mathbf{X}}\right)\\ &\geq\lambda_{\min}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)+\lambda_{\min}\left(-\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\\ &\geq\lambda_{\min}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)-\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\end{split}$ (20)
For both terms, we apply again [53, Eq. (519)]. For the LHS, one has:
$\mathbf{eig}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)=\{\lambda_{i}(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P}))\lambda_{j}(\mathbf{D}_{\mathbf{X}})\}_{1\leq i,j\leq n}$ (21)
We remark that $\tfrac{1}{2}(\mathbf{P}^{\top}+\mathbf{P})$ is a symetric bi-stochastic matrix, so $\lambda_{\min}(\tfrac{1}{2}(\mathbf{P}^{\top}+\mathbf{P}))\geq-1$. Therefore, $\lambda_{\min}(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P}))\geq-\tfrac{1}{2n}$. As a result, since the eigenvalues of $\mathbf{D}_{\mathbf{X}}$ are the $\|\mathbf{x}_{i}\|_{2}^{2}$, this yields:
$\lambda_{\min}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
Similarly, we have:
$-\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
from which we deduce that:
$\lambda_{\min}(\mathbf{B}_{\mathbf{X},\mathbf{P}})\geq-\tfrac{1}{n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
We can then lower bound:
$\begin{split}\lambda_{\min}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}+\mathbf{B}_{\mathbf{X},\mathbf{P}})&\geq\tfrac{1}{n^{2}}(\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top}))-\tfrac{1}{n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}\\ &=-\tfrac{1}{n}\gamma_{2,n}\end{split}$ (22)
which yields the $\tfrac{1}{n}\gamma_{2,n}$-weak convexity of $g$, and finally the $\gamma_{2,n}$-weak convexity of $\mathcal{GM}^{2}_{{r}_{n}}$.
Asymptotic. For any $T$, we note that, almost surely, $\|T\|^{2}_{L_{2}(r_{n})}\to\|T\|^{2}_{L_{2}(r)}$. As a result, since convexity is preserved under pointwise convergence and by virtue of Prop. (3.5), we study the (almost sure) convergence of $\gamma_{\textrm{inner},n}$ and $\gamma_{2,n}$.
We start by $\gamma_{\textrm{inner},n}$. We first remark that $\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}^{\top}\mathbf{X})$. Moreover, as $\mathbf{A}\in S_{d}^{+}(\mathbb{R})\mapsto\lambda_{\max}(\mathbf{A})$ is continuous and $\tfrac{1}{n}\mathbf{X}^{\top}\mathbf{X}\to\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}]$ almost surely, one has $\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ almost surely. Moreover, for any $n>d$, $\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})=0$. As a result, $\gamma_{\textrm{inner},n}\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ almost surely, which provides the desired asymptotic result.
We continue with $\gamma_{2,n}$. We first remark that $\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}\leq\sup_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$. As a result, by defining $\tilde{\gamma}_{2,n}=\gamma_{\textrm{inner},n}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$, $\mathcal{GM}^{2}_{{r}_{n}}$ is also $\tilde{\gamma}_{2,n}$-weakly convex. Moreover, $\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ does not depends on $n$, $\tilde{\gamma}_{2,n}\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ almost surely, which also provides the desired asymptotic result.
Appendix E Additional Empirical Results
Table 4: Disentanglement of regularizing the Encoder and the Encoder and Decoder as measured by DCI-D on two different datasets. We highlight best, second best, and third best results for each method and dataset.
DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
Shapes3D [33]
Base $67.7$$\pm 7.8$ $75.6$$\pm 8.7$ $88.1$$\pm 7.4$ $89.5$$\pm 7.9$
+ Enc-(DST) $69.2$$\pm 9.1$ $77.2$$\pm 7.5$ $87.7$$\pm 7.7$ $90.5$$\pm 5.9$
+ Enc-(GMG) $70.9$$\pm 9.5$ 79.6 ±$\pm 6.6$ 92.5 ±$\pm 5.9$ 93.5 ±$\pm 6.9$
+ Dec-(DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ 91.9 ±$\pm 9.4$
+ Dec-(GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
+ Enc-Dec-(GMG) 72.8 ±$\pm 7.7$ $79.3$$\pm 13.9$ 93.3 ±$\pm 5.0$ $91.8$$\pm 7.3$
DSprites [27]
Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
+ Enc-(DST) 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
+ Enc-(GMG) $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
+ Dec-(DST) $28.6$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
+ Dec-(GMG) 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
+ Enc-Dec-(GMG) 33.1 ±$\pm 14.9$ 40.2 ±$\pm 7.0$ $28.7$$\pm 14.6$ $46.0$$\pm 11.3$
Appendix F Experimental Details
All our experiments build on python 3 and the jax-framework [2], alongside ott-jax for optimal transport utilities.
To effectively conduct comprehensive and representative research on disentangled representation learning, we convert the public PyTorch framework proposed in  [60] to an equivalent jax variant. We verify our implementation through replications of baseline and HFS results in Roth et al. [60], mainting relative performance orderings and close absolute disentanglement scores (as measured using DCI-D, whose implementation directly follows from [43] and leverages gradient boosted tree implementations from scikit-learn).
For exact and fair comparison, we utilize standard hyperparamater choices from Roth et al. [60] (which leverages hyerparameters directly from [43], [44] and https://github.com/google-research/disentanglement_lib). Consequently, the base VAE architecture utilized across all experiment is the same as the one utilized in [60] and [44]: With image input sizes of $64\times 64\times N_{c}$ (with $N_{c}$ the number of input image channels, usually 3). The latent dimensionality, if not otherwise specified, is set to 10. The exact VAE model architecture is as follows:
  • Encoder: [conv($32,4\times 4$, stride 2) + ReLU] $\times$ 2, [conv($64,4\times 4$, stride 2) + ReLU] $\times$ 2, MLP(256), MLP(2 $\times$ 10)
  • Decoder: MLP(256), [upconv($64,4\times 4$, stride 2) + ReLU] $\times$ 2, [upconv($32,4\times 4$, stride 2) + ReLU], [upconv($n_{c},4\times 4$, stride 2) + ReLU]
Similar, we retain all training hyperparameters from [60] and [44]: Using an Adam optimizer ([34], $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) and a learning rate of $10^{-4}$. Similarly, we utilize a batch-size of $64$, for which we also ablate all baseline methods. The total number of training steps is set to $300000$.
As commonly done for this setting [43, 44, 60], we also perform a small grid search over all the hyperparameters. We report the full details in Tab. 5.
For $\lambda_{enc}$ and $\lambda_{dec}$, we set $\lambda_{enc}=0$ and $\lambda_{dec}=1$ for the Decoder setting and $\lambda_{enc}=1$ and $\lambda_{dec}=0$ for the Encoder setting. For the joint one, we set $\lambda_{enc}=0.5$ and $\lambda_{dec}=0.5$. All runs run on an RTX 2080TI GPU.
Table 5: Hyperparameter grid searches for different baseline and proposed methods.
Method Parameter Values
$\beta$-VAE $\beta$ [2, 4, 6, 8, 10, 16]
$\beta$-TCVAE $\beta$ [2, 4, 6, 8, 10, 16]
+ HFS $\gamma$ [1, 10]
+ DST $\lambda$ [0.1, 1, 5, 10, 20]
+ GMG $\lambda$ [0.1, 1, 5, 10, 20]

### v2

@@
-Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap
+Disentangled Representation Learning with the Gromov-Monge Gap
 Abstract
-Learning disentangled representations in an unsupervised manner is a fundamental challenge in machine learning.
+Learning disentangled representations from unlabelled data is a fundamental challenge in machine learning.
 Solving it may unlock other problems, such as generalization, interpretability, or fairness.
-While remarkably difficult to solve in general, recent works have shown that disentanglement is provably achievable under additional assumptions that can leverage geometrical constraints, such as local isometry.
-To use these insights, we propose a novel perspective on disentangled representation learning built on quadratic optimal transport.
-Specifically, we formulate the problem in the Gromov-Monge setting, which seeks isometric mappings between distributions supported on different spaces.
-We propose the Gromov-Monge-Gap (GMG), a regularizer that quantifies the geometry-preservation of an arbitrary push-forward map between two distributions supported on different spaces.
-We demonstrate the effectiveness of GMG regularization for disentanglement on four standard benchmarks.
-Moreover, we show that geometry preservation can even encourage unsupervised disentanglement without the standard reconstruction objective - making the underlying model decoder-free, and promising a more practically viable and scalable perspective on unsupervised disentanglement.
+Although remarkably challenging to solve in theory, disentanglement is often achieved in practice through prior matching.
+Furthermore, recent works have shown that prior matching approaches can be enhanced by leveraging geometrical considerations, e.g., by learning representations that preserve geometric features of the data, such as distances or angles between points.
+However, matching the prior while preserving geometric features is challenging, as a mapping that fully preserves these features while aligning the data distribution with the prior does not exist in general.
+To address these challenges, we introduce a novel approach to disentangled representation learning based on quadratic optimal transport.
+We formulate the problem using Gromov-Monge maps that transport one distribution onto another with minimal distortion of predefined geometric features, preserving them as much as can be achieved.
+To compute such maps, we propose the Gromov-Monge-Gap (GMG), a regularizer quantifying whether a map moves a reference distribution with minimal geometry distortion.
+We demonstrate the effectiveness of our approach for disentanglement across four standard benchmarks, outperforming other methods leveraging geometric considerations.
 1 Introduction
-Learning low-dimensional representations of high-dimensional data is a fundamental challenge in unsupervised deep learning [5, 43], where the emphasis is put on learning representations that allow for efficient adaptation across a wide range of tasks [5, 28, 43].
-Disentanglement [5, 27, 28, 43, 60] has shown significant promise in facilitating such generalization [5, 27, 28, 43, 44, 29, 60, 30, 4], alongside interpretability and fairness [42, 67].
-Most works [5, 27, 33, 13, 43, 60] regard disentanglement as a one-to-one map between learned representations and ground-truth latent factors; seeking to recover these factors from data alone in an unsupervised fashion.
-While unsupervised disentanglement is theoretically impossible [43], the inductive biases of autoencoder architectures ensure effective disentanglement in practice [59, 74, 29].
-Most approaches operate on variational autoencoder (VAE) frameworks [35], using objectives that match latent VAE posteriors to factorized priors [27, 33, 38, 10, 13].
-Recent works [29, 51, 31] provide a new perspective, showing how geometric constraints on representation spaces may enable disentanglement.
-In particular, Horan et al.
-[29] demonstrate that unsupervised disentanglement is always possible under the assumption of local isometry to the data and sufficient non-Gaussianity of the generative factors, thus supporting the isometry desideratum.
-In this work, we show how these geometric desiderata can be effectively quantified through the lens of optimal transport (OT) theory [61, 55], by treating mapping to or from the latent space as transport maps $T$ from or to the data manifold, respectively.
-While the more classical tools described in the OT toolbox [55] provide tools to align distributions supported on the same space $\mathcal{X}$, using an inter-domain cost $c(\mathbf{x},\mathbf{y})$ for any two points $\mathbf{x},\mathbf{y}\in\mathcal{X}$, such approaches are not well defined when mapping between latent and data space with very different dimensionalities, for which no “natural” cost function exists.
-Instead, we rely in this work on the Gromov-Wasserstein (GW) formalism [64, 48, 69, 63] formulation of OT, which considers instead intra-domain costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ in each space.
-This formulation seeks instead the most geometry-preserving mapping between two distributions by favoring maps that minimize the distortion between intra-domain costs by comparing $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime}),c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$.
-While this distortion itself can be used as a regularization for a given map $T$ as done in Nakagawa et al.
-[51], we borrow a page from the Monge gap formalism proposed by Uscidda and Cuturi [68] to propose a modified and better-posed regularizer that substracts to the distortion of $T$, the best distortion that can be achieved by any map (or a proxy thereof).
-In a nod to [68], we propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving).
-In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results, showing notably that the GMG and its finite sample version are weakly convex.
-We show how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41].
-Motivated by this formalism, we build upon the Monge gap - a regularizer introduced in Uscidda and Cuturi [68] that measures whether a map $T$ transports a reference distribution at minimal displacement cost - to propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving).
-In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results highlighting that the GMG and its finite sample version are weakly convex.
-We lay out how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41].
-Our experiments on four standard disentangled representation learning benchmarks show that the integration of these geometry-preserving desiderata through the Gromov-Monge Gap (GMG) significantly improves disentanglement performance across various methods, from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with support factorization [60], outperforming a distortion-based regularization.
-Moreover, we demonstrate that these geometric regularizations can replace the standard reconstruction loss, enabling measurable unsupervised disentanglement even without a decoder, which is not feasible in standard frameworks that rely on the decoder-based reconstruction term to prevent collapse.
-This finding suggests the potential for more scalable unsupervised disentangled representation learning approaches and bridges to popular, weakly- or self-supervised encoder-only representation learning methods [14, 72, 3, 24].
+Learning low-dimensional representations of high-dimensional data is a fundamental challenge in unsupervised deep learning (Bengio et al., 2014).
+Emphasis is put on learning representations that allow for efficient and robust adaptation across a wide range of tasks (Higgins et al., 2018, Locatello et al., 2019a).
+The fundamental property of disentanglement has shown significant promise to improve generalization (Locatello et al., 2020, Roth et al., 2023, Hsu et al., 2023, Barin-Pacela et al., 2024), interpretability and fairness (Locatello et al., 2019b, Träuble et al., 2021).
+Most works regard disentanglement as a one-to-one map between learned representations and ground-truth latent factors, effectively seeking to recover these factors from data alone in an unsupervised fashion.
+While unsupervised disentanglement is theoretically impossible (Locatello et al., 2019a), the inductive biases of autoencoder architectures ensure effective disentanglement in practice (Rolinek et al., 2019, Zietlow et al., 2021).
+Most approaches operate using variational autoencoder (VAE) frameworks (Kingma and Welling, 2014), using objectives that match latent VAE posteriors to factorized priors (Higgins et al., 2017, Kim and Mnih, 2018, Kumar et al., 2018, Burgess et al., 2018, Chen et al., 2018).
+More recently, studies such as Gropp et al.
+(2020), Chen et al.
+(2020a), Lee et al.
+(2022), Horan et al.
+(2021), Nakagawa et al.
+(2023), Huh et al.
+(2023), Hahm et al.
+(2024) have provided a new perspective, showing that geometric constraints on representation spaces may also enable disentanglement.
+Typically, latent representations are encouraged to preserve key geometric features of the data distribution, such as (scaled) distances or angles between samples.
+Horan et al.
+(2021) even demonstrate that unsupervised disentanglement is always possible provided that the latent space is locally isometric to the data, further supporting the geometric desiderata.
+However, combining prior matching with these geometric aspects is challenging.
+In general, a mapping that perfectly aligns the data distribution with the prior while fully preserving the geometric features of interest may not exist.
+This leads to an inherent trade-off: Practitioners must carefully fine-tune regularization terms, either by altering prior matching to prioritize geometry preservation, or vice-versa.
+In this work, we demonstrate how to effectively combine geometric desiderata with prior matching within the VAE framework, using optimal transport (OT) theory (Santambrogio, 2015, Peyré and Cuturi, 2019).
+By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2020, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports.
+Specifically, we consider cost functions $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ on $\mathcal{X}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$ on $\mathcal{Y}$ that encode geometric features such as scaled distances or angles.
+Consequently, the resulting mapping transforms one distribution onto the other while preserving these geometric features as much as possible.
+Our Contribution: A novel OT-based approach to disentanglement through geometric considerations.
+  • (i)
+    We address the challenge of learning disentangled representations using geometric constraints by leveraging Gromov-Monge mappings between the data and prior distributions.
+Since fully preserving geometric features—such as (scaled) distances or angles between points—during the alignment of these two distributions is generally impossible, we aim to find an alignment that, instead, minimizes the distortion of these features, thereby preserving them as much as possible.
+  • (ii)
+    Inspired by (Uscidda and Cuturi, 2023), we introduce the Gromov-Monge Gap (GMG), a regularizer that measures how closely a map $T$ approximates a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+GMG measures whether $T$ transports distributions with minimal distortion w.r.t.
+$c_{\mathcal{X}},c_{\mathcal{Y}}$.
+We propose an efficient procedure to compute GMG and describe how to integrate it within the VAE framework.
+  • (iii)
+    We show that when $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ encode scaled distances or angles, the GMG and its finite-sample counterpart are weakly convex functions.
+In both cases, we precisely characterize the weak convexity constants and analyze their practical implications for practitioners.
+  • (iv)
+    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods—from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023)—notably outperforming other methods that leverage geometric constraints.
 2 Background and Related Works
@@
 The Disentanglement Formalism.
-Disentanglement has varying operational definitions [28, 43, 60].
-In this work, we follow the common understanding [43, 44, 67, 60] where data $\mathbf{x}$ is generated by a process $p(\mathbf{x}|\mathbf{z})$ operating on ground-truth latent factors $\mathbf{z}\sim p(\mathbf{z})$, modeling underlying source of variations (s.a.
+Disentanglement has varying operational definitions.
+In this work, we follow the common understanding (Higgins et al., 2017, Locatello et al., 2019a; 2020, Träuble et al., 2021, Roth et al., 2023) where data $\mathbf{x}$ is generated by a process $p(\mathbf{x}|\mathbf{z})$ operating on ground-truth latent factors $\mathbf{z}\sim p(\mathbf{z})$, modeling underlying source of variations (s.a.
 object shape, color, background…).
-Given a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, unsupervised disentangled representation learning aims to find a mapping $e_{\phi}$ s.t.
+Given a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, $\mathbf{x}_{i}\sim p_{\textrm{data}}$, unsupervised disentangled representation learning aims to find a mapping $e_{\phi}$ s.t.
 $e_{\phi}(\mathbf{x}_{i})\approx\mathbb{E}[\mathbf{z}|\mathbf{x}_{i}]$, up to element-wise transformations.
@@
 Unsupervised Disentanglement through Prior Matching.
-Most unsupervised disentanglement methods operate on variational autoencoders (VAEs)[35], which define a generative model of the form $p_{\theta}(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z})$.
+Most unsupervised disentanglement methods operate on variational autoencoders (VAEs, Kingma and Welling (2014)), which define a generative model of the form $p_{\theta}(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z})$.
 Here, $p_{\theta}(\mathbf{x}|\mathbf{z})$ is a product of exponential family distributions with parameters computed by a decoder $d_{\theta}(\mathbf{z})$.
-The latent prior $p(\mathbf{z})$ is usually chosen to be a normal Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$, and the probabilistic encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ is realized through a neural network $e_{\phi}(\mathbf{x})$ that predicts the parameters of the latent such that $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|e_{\phi}(\mathbf{x}))$.
-The $\beta$-VAE [27]
-$\mathcal{L}_{\beta}(\theta,\phi)\vcentcolon=\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})-\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))]$ (1)
-achieves disentanglement by enforcing stronger, $\beta$-weighted prior matching on top of the reconstruction objective, assuming statistical factor independence [60].
-Several follow-ups refine latent prior matching through different objectives or prior choices [13, 38, 10, 59].
+The latent prior $p(\mathbf{z})$ is usually chosen as a standard Gaussian $\mathcal{N}(\mathbf{z}|\mathbf{0}_{d},\mathbf{I}_{d})$, and the probabilistic encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ is implemented through neural networks $e_{\phi}(\mathbf{x}),\sigma_{\phi}(\mathbf{x})$ that predicts the latent parameters so that $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|e_{\phi}(\mathbf{x}),\sigma^{2}_{\phi}(\mathbf{x}))$.
+The $\beta$-VAE (Higgins et al., 2017) achieves disentanglement by minimizing
+$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}],$ (1)
+which enforces $\beta$-weighted prior matching on top of the reconstruction loss, assuming statistical factor independence (Roth et al., 2023).
+Several follow-ups refine latent prior matching through different losses or prior choices (Kim and Mnih, 2018, Chen et al., 2018, Kumar et al., 2018, Burgess et al., 2018, Rolinek et al., 2019, Moor et al., 2021, Balabin et al., 2024).
 Disentanglement through a Geometric Lens.
-Recent studies [26, 12, 41, 51, 31] indicate that disentanglement can arise by encouraging learned representations to preserve meaningful geometric features of the data, such as scaled distances between samples.
-Notably, Horan et al.
-[29] demonstrated that disentanglement is provably feasible when the generative factors are sufficiently non-Gaussian and locally isometric to the data.
-In this work, we explore how to promote geometry preservation using quadratic OT between the latent and data spaces, which we introduce in the next section.
+Recent research suggests that disentanglement can arise by imposing geometric constraints on representation spaces (Horan et al., 2021).
+Most works (Karras et al., 2020, Huh et al., 2023) promote latent representations to preserve scaled distances between data points.
+Closest to our work, Lee et al.
+(2022) use a Jacobian-based regularization within a VAE, while Nakagawa et al.
+(2023) employ a distortion-based regularization, directly comparing pairwise squared Euclidean distances, within a Wasserstein AE (WAE, Tolstikhin et al.
+(2019)).
 2.2 Quadratic Optimal Transport
 Gromov-{Monge, Wasserstein} Formulations.
-OT [55] involves transferring one probability distribution to another while incorporating inductive biases.
-When these distributions lie on incomparable domains, the task is addressed using the Gromov-Monge and GW problems, also known as OT quadratic formulations.
-Formally, consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$.
-For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the Gromov-Monge problem [50] seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
+OT (Peyré and Cuturi, 2019) involves transferring a probability distribution to another with inductive biases.
+When these distributions lie on incomparable domains, one relies on the Gromov-Monge (GM) and GW problems, also known as quadratic formulations of OT.
+Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$.
+For technical reasons, we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are CPD kernels (see Def.
+(A.1)).
+For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
 $\inf_{T:T\sharp p=q}\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))|^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})\,.$ (GMP)
 When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
-However, solving this problem is difficult, and existence is not guaranteed in general [19].
-Moreover, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty.
+However, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty in that case.
 Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e.
-distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the Gromov-Wasserstein (GW) problem [48, 64]
-$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
-A solution $\pi^{\star}$ of (GWP) always exists, making $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q)$ a well-defined quantity.
+distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2020)
+$\mathrm{GW}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
+A solution $\pi^{\star}$ to (GWP) always exists, making $\mathrm{GW}(p,q)$ a well-defined quantity.
 It quantifies the minimal distortion of the geometries induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ achievable when coupling $p$ and $q$.
 Discrete Solvers.
-When both $p$ and $q$ are instantiated as samples, the GW Prob.
-((GWP)) translates to a quadratic assignment problem, whose objective can be regularized using entropy [15].
+When both $p$ and $q$ are instantiated as samples, GW Prob.
+(GWP) translates to a quadratic assignment problem, whose objective can be regularized using entropy (Cuturi, 2013, Peyré et al., 2016).
 For empirical measures ${p}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$, ${q}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{\mathbf{y}_{j}}$ and $\varepsilon\geq 0$, we set:
-$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({p}_{n},{q}_{n})\vcentcolon=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,$ (EGWP)
-where $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})$.
+$\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})\vcentcolon=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(\mathbf{C}_{\mathcal{X}_{i,i^{\prime}}}-\mathbf{C}_{\mathcal{Y}_{j,j^{\prime}}})^{2}\,\mathbf{P}_{i,j}\mathbf{P}_{i^{\prime},j^{\prime}}-\varepsilon H(\mathbf{P})\,,$ (EGWP)
+with $\mathbf{C}_{\mathcal{X}}=[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{i,i^{\prime}}$, $\mathbf{C}_{\mathcal{Y}}=[c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}})]_{j,j^{\prime}}\in\mathbb{R}^{n\times n}$, $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{i,j}\log(\mathbf{P}_{i,j})$.
 As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$.
-In addition to yielding better statistical [73] and regularity [58] properties, entropic regularization also enhances computational performance.
-In practice, we can solve (EGWP) using a mirror descent scheme that iterates the Sinkhorn algorithm [54, 62].
-Neural Solvers.
-While for classical OT, numerous neural methods have been proposed [45, 37, 22, 68, 66], the GW setting has received less attention.
-To our knowledge, the only neural Gromov-Monge formulation proposed thus far is [52], which involves a min-max-min optimization procedure.
-On the other hand, Klein et al.
-[36] recently proposed an approach to compute neural GW couplings.
+In addition to yielding greater convexity (Rioux et al., 2023) and better statistical (Zhang et al., 2023) and properties, entropic regularization improves computational performance.
+We can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (Peyré et al., 2016).
+This solver has $\mathcal{O}(n^{2})$ memory complexity.
+Its time complexity depends on $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+For general $c_{\mathcal{X}},c_{\mathcal{Y}}$, it runs in $\mathcal{O}(n^{3})$.
+However, for the most common practical choices of $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, it can also be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$, as detailed by Scetbon et al.
+(2022, §3 & Alg.
+2).
+Neural GW.
+While for classical OT, numerous neural methods have been proposed (Makkuva et al., 2020, Korotin et al., 2022, Tong et al., 2023, Eyring et al., 2024), the GW setting has received less attention.
+Carrasco et al.
+(2024) propose an approach for solving GM with inner product costs, that involves a min-max-min optimization procedure.
+Klein et al.
+(2024) propose a framework to computeentropic GW couplings.
+Yet, these formulations are end-to-end, and not based on a regularizer that can be flexibly added to a loss function.
+Therefore, they cannot be combined with other models, as we do in this work with the $\beta$-VAE.
+In a concurrent work, Sotiropoulou and Alvarez-Melis (2024) introduced the Gromov-Monge gap.
+However, (i) they do not apply it for disentangled representational learning, and (ii) they do not investigate its properties, as we do in this work.
 3 Disentanglement with the Gromov-Monge gap
-This section details our novel optimal transport perspective to achieve disentanglement from geometric considerations (see § 2.1), using the VAE framework.
-To achieve this, we first investigate how one can promote an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains $\mathcal{X}$ and $\mathcal{Y}$ to preserve predefined geometric features.
-In a VAE, $T$ can represent either the encoder $e_{\phi}$, which produces latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from the latent codes.
-As a result, in the former, the source domain $\mathcal{X}$ is the data, and the target domain $\mathcal{Y}$ is the latent space, with roles swapped in the latter.
-If we assume that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, it is equivalent whether $e_{\phi}$ preserves the geometric features from data to latents or $d_{\theta}$ preserves them from latents to data.
-Consequently, in what follows, $T$ can refer to either the encoder or the decoder without distinction.
+This section outlines our novel OT-based approach to disentanglement through geometric considerations within the VAE framework.
+First, we explore how an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains can be encouraged to preserve predefined geometric features.
+In a VAE, $T$ may represent either the encoder $e_{\phi}$, which generates latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from these codes.
+In the case of the encoder, the source domain $\mathcal{X}$ corresponds to the data, and the target domain $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped.
+Assuming that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, i.e., $e_{\phi}\circ d_{\theta}=\mathrm{I}d$, the preservation of geometric features by either $e_{\phi}$ or $d_{\theta}$ becomes equivalent.
+Therefore, in the following sections, we refer to $T$ as either the encoder or the decoder without loss of generality.
 Outline.
-Leveraging this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1.
-We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves points while preserving geometric features as much as possible, i.e., minimizing distortion.
-§3.3 then shows how the GMG can be estimated and computed from samples to be practically applicable in the VAE framework, which transitions into §3.4 studying convexity properties of the GMG.
-Put together, §3.13.4 define the practical GMG which allows us to learn a latent space that matches, as much as possible, geometrical constraints in the data space.
-Finally in §3.5, we use the GMG with different choices of costs to propose effective disentangled representation learning objectives.
+Building on this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1.
+We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion.
+§3.3 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework.
+This transitions into §3.4 studying (weak) convexity properties of the GMG, as an operator.
+Finally, in §3.5, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
 3.1 From the distortion…
-We encode the geometric features of interest through two cost functions defined on each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$.
-We then want $T$ to preserve these costs, i.e., $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$.
-In practice, two types of cost functions are particularly meaningful.
+We encode the geometric features of interest using two cost functions, one for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$.
+We aim for $T$ to preserve the geometric feature, which means that $T$ preserves these costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$.
+In practice, two types of cost function from this family are particularly relevant:
   • [i]
-    (Scaled) squared Euclidean distance: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$ and $c_{\mathcal{Y}}=\alpha^{2}\|\mathbf{y}-\mathbf{y}\|_{2}^{2}$, with $\alpha>0$.
+    (Scaled) squared L2 distance: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})=\alpha^{2}\|\mathbf{y}-\mathbf{y}\|_{2}^{2}$, with $\alpha>0$.
 A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ preserves the scaled distances between the points, i.e.
@@
 […]
-In the following, we say that $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] if they belong to these families of costs.
-Introducing a reference distribution $r\in\mathcal{P}(\mathcal{X})$, weighting the areas of $\mathcal{X}$ where we penalize deviations of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ from $c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$, we can quantify this property using the following criterion.
+We refer to these costs via $\mathbf{L2^{2}}$ for [i] with $\alpha=1$, $\mathbf{ScL2^{2}}$ for [i] with $\alpha\neq 1$ and $\mathbf{Cos}$ for [ii].
+Introducing a reference distribution $r\in\mathcal{P}(\mathcal{X})$, weighting the areas of $\mathcal{X}$ where we penalize deviations of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ from $c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$, we can quantify this property using the following criterion:
 Definition 3.1 (Distortion).
-The distortion (DST) of a map $T$ is defined as:
-$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T):=\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})$ (DST)
-$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies how much $T$ distorts the geometric features induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ on the support of $r$, i.e., when $\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$.
-In disentangled representation learning, it can be desirable to regularize the decoder to be isometric [29, 51, 31].
-However, a fully geometry-preserving mapping might not necessarily exist between the latent space and the data distribution.
-This means there usually exists an inherent trade-off between the accurate reconstruction of the data distribution and this reconstruction being, e.g., a "fully isometric" map.
-If such a map does not exist, the reconstruction loss and the distortion term cannot be 00 simultaneously.
-In practice, this means that the distortion loss will move away from accurately reconstructing the data, which negatively impacts the quality of the learned latent representations.
-This naturally raises the question of how to formulate a geometric regularization that takes the most geometry-preserving mapping into account.
-[…]
+The distortion (DST) of a map $T$, for cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$ and reference distribution $r$, is defined as:
+$\mathrm{DST}_{r}(T):=\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\,.$ (DST)
+$\mathrm{DST}_{r}(T)$ quantifies how much $T$ distorts geometric features encoded by $c_{\mathcal{X}},c_{\mathcal{Y}}$ on the support of $r$, that is, when $\mathrm{DST}_{r}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$.
+Nakagawa et al.
+(2023) suggest promoting geometry preservation by regularizing the encoder $d_{\theta}$ using the DST, with $\mathbf{ScL2^{2}}$ as costs, and the latent representation as reference distribution, namely $r=e_{\phi}\sharp p_{\textrm{data}}$.
+While they use it within a WAE, translated in our VAE setting, this corresponds to minimizing
+$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda\,\mathrm{DST}_{r}(d_{\theta})}_{\textrm{(iii) geom.
+preservation}},$ (2)
+where $\lambda>0$.
+Given the choice of the costs, $d_{\theta}$ is distortion-free (i.e., $\mathrm{DST}_{r}(d_{\theta})=0$), if it is a scaled isometry.
+Since a scaled isometry that maps the prior onto the data distribution may not exist, there is an inherent trade-off between minimizing terms (ii) and (iii), which are responsible for achieving practical disentanglement.
+As these terms cannot be simultaneously minimized to 00, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations.
+This naturally raises the question of how to avoid this over-penalization.
+Instead of seeking a distortion-free decoder, we should seek a decoder that transports the prior to the data distribution with minimal distortion of the costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+In other words, the decoder should be a Gromov-Monge map between the prior and the data distributions for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ (see 2.2).
+Conversely, if we choose to regularize the encoder, the same reasoning applies by swapping the roles of the prior and the data distribution.
+In the next section, we introduce a regularizer to fit Gromov-Monge maps, which we will use as a replacement to the DST.
 3.2 …to The Gromov-Monge Gap
-Recently, Uscidda and Cuturi [68] introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost.
-In practice, this regularizer is combined with fitting losses to compute Monge maps, which are defined by two main features: (i) they fit a marginal constraint with (ii) minimal displacement cost.
+Figure 1: Sketch of $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})$ for two different maps $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$.
+We use a discrete reference distribution $\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}$ on 3 points, and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}$, hence measuring if $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ minimally distorts the distances.
+On the left, $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ is the optimal map, $\boldsymbol{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T^{\star}}}$, and maps the three points with minimal (yet non-zero) distortion, which is measured as the sum of the squared lengths of the orange segments.
+This results in $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})=0$.
+On the right, $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ swaps two points compared to $\boldsymbol{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T^{\star}}}$—specifically, $\boldsymbol{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}_{2}}}$ and $\boldsymbol{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}_{3}}}$—causing a higher distortion than the minimal one, and measured as the sum of the squared lengths of the red segments.
+This results in $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})>0$, equal to the gap between these distortions.
+Recently, Uscidda and Cuturi (2023) introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost.
 Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
-In § 3.5, we use it, alongside fitting losses, to compute Gromov-Monge maps, as defined in Eq.
-(GMP), which are similarly defined by (i) fitting a marginal constraint with (ii) minimal distortion cost.
 Definition 3.2 (Gromov-Monge gap).
-The Gromov-Monge gap (GMG) of a map $T$ is defined as:
-$\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ (GMG)
-From Eq.
-(GWP), we recall that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ is the minimal distortion achievable when transporting $r$ to $T\sharp r$.
-Thus, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies the difference between the distortion incurred when transporting $r$ to $T\sharp r$ via $T$, to this minimal distortion.
-More formally, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is the optimality gap of $T$ in the Gromov-Monge Prob.
-(GMP) between $r$ and $T\sharp r$, which is always feasible, even when $r$ is discrete, as $T$ belongs to the constraint set.
-In light of this, it is a well-defined and
+The Gromov-Monge gap (GMG) of a map $T$, for cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$ and reference distribution $r$, is defined as:
+$\mathrm{GMG}_{r}(T)\vcentcolon=\mathrm{DST}_{r}(T)-\mathrm{GW}(r,T\sharp r)$ (GMG)
+We recall from Eq.
+(GWP) that $\mathrm{GW}(r,T\sharp r)$ represents the minimal distortion of $c_{\mathcal{X}},c_{\mathcal{Y}}$ achievable when transporting $r$ to $T\sharp r$.
+Thus, the GMG quantifies the difference between the distortion incurred when transporting $r$ to $T\sharp r$ via $T$, and this minimal distortion.
+Formally, when Prob.
+(GMP) and Prob.
+(GWP) between $r$ and $T\sharp r$ are equivalent, the the GMG is the suboptimality gap of $T$ in Prob.
+(GMP).
+This is the case, for example, when $r$ is a density and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ (Dumont et al., 2022).
+Otherwise, the GMG is the suboptimality gap of $\pi=(\mathrm{I}_{d},T)\sharp r$ in Prob.
+(GWP) between $r$ and $T\sharp r$.
+In light of this, it is a well-defined quantity and:
   • The GMG measures how close $T$ is to be a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
-Indeed, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0$ with equality i.f.f.
-$T$ is a Gromov-Monge map solution of Prob.
-(GMP) between $r$ and $T\sharp r$, i.e., $T$ moves $r$ with minimal (but eventually non zero) distortion.
+Indeed, $\mathrm{GMG}_{r}(T)\geq 0$ with equality if $T$ is a Gromov-Monge map solution of Prob.
+(GMP) between $r$ and $T\sharp r$, i.e., $T$ moves $r$ with minimal (but eventually non zero) distortion; see App.
+B.1.
   • When transport without distortion is possible, the GMG coincides with the distortion.
-When there exists another map $U:\mathcal{X}\to\mathcal{Y}$ transporting $r$ to $T\sharp r$ with zero distortion, i.e., $U\sharp r=T\sharp r$ and $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(U)=0$, then $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$.
-Indeed, $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)=0$ in that case, as the coupling $\pi=(\mathrm{Id},U)\sharp r$ sets the GW objective to zero, thereby minimizing it.
-The last point (ii) is fundamental and illustrates how the GMG functions as a debiased distortion.
-Indeed, it compares the distortion induced by $T$ to a baseline distortion, defined as the minimal achievable distortion when transforming the reference distribution into its image under $T$.
-Thus, when transformation without any distortion is achievable, the reference distortion becomes zero, and the GMG precisely aligns with the distortion itself, i.e., $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$.
-In this context, the GMG offers the optimal compromise: it avoids the over-penalization induced by the distortion when fully preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ is not feasible, yet it coincides with it when such preservation is feasible.
+When there exists another map $U:\mathcal{X}\to\mathcal{Y}$ transporting $r$ to $T\sharp r$ with zero distortion, i.e., $U\sharp r=T\sharp r$ and $\mathrm{DST}_{r}(U)=0$, then $\mathrm{GMG}_{r}(T)=\mathrm{DST}_{r}(T)$.
+Indeed, $\mathrm{GW}(r,T\sharp r)=0$ in that case, as the coupling $\pi=(\mathrm{Id},U)\sharp r$ sets the GW objective to zero, thereby minimizing it.
+The last point illustrates how the GMG functions as a debiased distortion.
+It compares the distortion induced by $T$ to a baseline distortion, defined as the minimal achievable distortion when transforming $r$ into $T\sharp r$.
+Thus, when transformation without distortion is achievable, this baseline becomes zero, and the GMG equals the distortion.
+Consequently, the GMG offers the optimal compromise: it avoids the over-penalization induced by the distortion when fully preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ is not feasible, yet it coincides with it when such full preservation is feasible.
+See Fig.
+1 for a simple illustration.
 The Influence of the Reference Distribution.
-A crucial property of $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$.
-Formally, if $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\operatorname{supp}(s)\subseteq\operatorname{supp}(r)$, then $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$.
+A crucial property of $\mathrm{DST}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$.
+Formally, if $\mathrm{DST}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, then $\mathrm{DST}_{s}(T)=0$.
 This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion?
 We answer this question positively with Prop.
-(3.3) when the costs are the (scaled) Euclidean distances or the cosine similarity.
-Intuitively, this means that if $T$ moves $r$ while preserving (scaled) distances or angles as much as possible, it will also preserve these properties as much as possible when moving any "smaller" distribution within $r$.
+(3.3).
 Proposition 3.3.
-When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] (see § 3.1), if $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$, then for any $s\in\mathcal{P}(\mathcal{X})$ s.t.
-$\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, one has $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$.
+If $\mathrm{GMG}_{r}(T)=0$, $\forall s\in\mathcal{P}(\mathcal{X})$ s.t. $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, one has $\mathrm{GMG}_{s}(T)=0$.
 3.3 Estimation and Computation from Samples
@@
 (DST) and Eq.
-(GMG) using i.i.d.
-samples $\mathbf{x}_{1},...,\mathbf{x}_{n}$ from the reference distribution $r$.
-We then consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use a plug-in estimator for both cases, i.e., we estimate the distortion via
-$\begin{split}\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (2)
-and the GMG via $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$.
-To better understand what the discrete GMG quantifies, we can reformulate it using the minimal distortion achieved by a permutation $\sigma\in\mathcal{S}_{n}$ between the $\mathbf{x}_{i}$ and the $T(\mathbf{x}_{i})$.
-Proposition 3.4.
-When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] , the empirical GMG reads:
-$\begin{split}\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\sigma\in\mathcal{S}_{n}}\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\left(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{\sigma(i)}),T(\mathbf{x}_{\sigma(j)}))\right)^{2}\end{split}$ (3)
-As a Monte Carlo estimator, $\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is naturally consistent.
-We can ask the same question for $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)$, which requires studying the convergence of the empirical GW distance $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$.
-For the costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ of interest, we show that consistency holds.
-Proposition 3.5.
-When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii], $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\rightarrow\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ almost surely.
-Efficient Computation.
-Computing $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ requires solving a discrete GW problem between $r_{n}$ and $T\sharp r_{n}$ to obtain $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$.
-To alleviate computational challenges, we estimate this term using an entropic regularization $\varepsilon\geq 0$, as introduced in Eq.
+(GMG) from samples $\mathbf{x}_{1},...,\mathbf{x}_{n}\sim r$.
+We consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use plug-in estimators, i.e.
+$\begin{split}\mathrm{DST}_{{r}_{n}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (3)
+and $\mathrm{GMG}_{{r}_{n}}(T)=\mathrm{DST}_{{r}_{n}}(T)-\mathrm{GW}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$.
+Algorithm 1 GMG$(\mathbf{x}_{1},\dots\mathbf{x}_{n},T,\varepsilon)$.
+[…]
+2:$\mathbf{t}_{1},\dots\mathbf{t}_{n}\leftarrow T(\mathbf{x}_{1}),\dots,T(\mathbf{x}_{n})$.
+3:$\mathbf{C}_{\mathcal{X}}\leftarrow[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{1\leq i,i^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{X}})$
+4:$\mathbf{C}_{\mathcal{Y}}\leftarrow[c_{\mathcal{Y}}(\mathbf{t}_{j},\mathbf{t}_{j^{\prime}})]_{1\leq j,j^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{Y}}$)
+5:$\textsc{DST}\leftarrow\texttt{mean}((\mathbf{C}_{\mathcal{X}}-\mathbf{C}_{\mathcal{Y}})^{2})$triangleright $n^{2}$
+6:$\varepsilon\leftarrow\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$triangleright usually $\mathcal{O}(n^{2})$
+7:$\textsc{GW}\leftarrow\texttt{GW}(\mathbf{C}_{\mathcal{X}},\mathbf{C}_{\mathcal{Y}},\varepsilon)$triangleright $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$
+8:return DST - GW
+Efficient and Stable Computation.
+Computing the GMG requires solving a discrete GW problem between $r_{n}$ and $T\sharp r_{n}$ to get $\mathrm{GW}({r}_{n},T\sharp{r}_{n})$.
+We compute this term using an entropic regularization $\varepsilon\geq 0$, as in Eq.
 (EGWP):
-$\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})\,.$ (4)
-Choosing $\varepsilon=0$, we recover the unregularized one $\mathcal{GM}_{{r}_{n},0}^{c_{\mathcal{X}},c_{\mathcal{Y}}}=\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$.
-Moreover, the entropic estimator preserves the positivity, as for $\varepsilon\geq 0$, we have $\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}\geq 0$ (see A.1).
-As described in § 2, we compute $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al.
-[54]’s solver.
-While it always has $\mathcal{O}(n^{2})$ memory complexity, when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, this solver runs in $\mathcal{O}(n^{2}d)$ time [62, Alg.
-2].
-Since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, the computation of the GMG for the costs of interest [i] or [ii] scales as $\mathcal{O}(n^{2}d)$ in time.
-We use ott-jax’s [16] implementation of this scheme.
+$\mathrm{GMG}_{{r}_{n},\varepsilon}(T)\vcentcolon=\mathrm{DST}_{{r}_{n}}(T)$ (4)
+$\,\,-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,.$
+Choosing $\varepsilon=0$, we recover $\mathrm{GMG}_{{r}_{n},0}(T)=\mathrm{GMG}_{{r}_{n}}(T)$.
+Moreover, the entropic estimator preserves positivity, as for $\varepsilon\geq 0$, $\mathrm{GMG}_{r_{n},\varepsilon}(T)\geq 0$ (see B.3).
+We compute $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al.
+(2016)’s solver introduced in 2.2.
+We use the implementation provided by ott-jax (Cuturi et al., 2022).
+In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$,$\mathbf{C}_{\mathcal{Y}}$.
+We define a scale $\varepsilon_{0}$ and set $\varepsilon=\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$.
+Standard options for the statistic include $\texttt{stat}\in\{\texttt{mean},\texttt{max},\texttt{std}\}$.
+This procedure is equivalent to running the entropic GW solver on the re-scaled cost matrices $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$ with $\varepsilon=\varepsilon_{0}$; see App.
+B.2.
+It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013), then extended to the entropic GW solver (Cuturi et al., 2022).
+Additionally, as suggested by Scetbon et al.
+(2022), to promote the solver’s stability, we always initialize it with $\mathbf{P}_{\textrm{init}}=\tfrac{1}{n^{2}}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$.
+Computational Complexity.
+For usual costs, including inner products, $\ell_{p}^{q}$ distances, and standard CPD kernels, the computation of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ (resp.
+$c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$) can be done in $\mathcal{O}(d_{\mathcal{X}})$ time (resp.
+$\mathcal{O}(d_{\mathcal{Y}})$).
+Consequently, the DST can be computed in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ time.
+Furthermore, as discussed in § 2.2, the time complexity of the entropic GW solver is $\mathcal{O}(n^{3})$ in general, but can be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, or $c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$.
+Therefore, since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, this solver runs in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ for the costs of interest $\mathbf{(Sc)L2^{2}}$ and $\mathbf{Cos}$.
+The complete algorithm, along with a time complexity analysis of each step, is described in Alg.
+1.
 3.4 (Weak) Convexity of the Gromov-Monge gap
-As laid out, the GMG can be used as a regularization loss to push any model $T$ to be more geometry-preserving.
-A natural question that arises when defining such a regularizer is: what are its regularity properties, and in particular, is it convex?
-In the following, we study the convexity of $T\mapsto\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$, and its finite-sample counterpart $T\mapsto\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$.
-We focus on the costs of interest [i] or [ii].
-For simplicity, we replace cosine similarity with inner product—i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$—as they are equivalent, up to normalization of $r$ and $T$.
-We then study the convexity of the GMG for (i) the (scaled) squared Euclidean distances and (ii) the inner product, denoted respectively by (i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$.
-To that end, we introduce a weaker notion of convexity, previously defined for functions on $\mathbb{R}^{d}$ [17], which we extend here to $L_{2}(r)=\{T\,|\,\|T\|_{L_{2}(r)}^{2}\vcentcolon=\int_{\mathcal{X}}\|T(\mathbf{x})\|_{2}^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})<+\infty\}$.
+[…]
 Definition 3.4 (Weak convexity.).
@@
 Theorem 3.5.
-Both $\mathcal{GM}_{r}^{2}$ and $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, as well as their finite sample versions, are weakly convex.
+Both $\mathrm{GMG}_{r}^{2}$ and $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$, as well as their finite sample versions, are weakly convex.
   • Finite sample.
@@
 the support of $r_{n}$, as rows.
-Then, (i) $\mathcal{GM}_{r_{n}}^{2}$ and (ii) $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2,n}$ and (ii) $\gamma_{\textrm{inner},n}$-weakly convex, where: $\gamma_{\textrm{inner},n}=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})$ and $\gamma_{2,n}=\gamma_{\textrm{inner},n}+\max_{i=1\dots n}\|\mathbf{x}_{i}\|_{2}^{2}$.
+Then, (i) $\mathrm{GMG}_{r_{n}}^{2}$ and (ii) $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2,n}$ and (ii) $\gamma_{\textrm{inner},n}$-weakly convex, where: $\gamma_{\textrm{inner},n}=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})$ and $\gamma_{2,n}=\gamma_{\textrm{inner},n}+\max_{i=1\dots n}\|\mathbf{x}_{i}\|_{2}^{2}$.
   • Asymptotic.
-(i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2}$ and (ii) $\gamma_{\textrm{inner}}$-weakly convex, where: $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ and $\gamma_{2,n}=\gamma_{\textrm{inner}}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
+(i) $\mathrm{GMG}_{r}^{2}$ and (ii) $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2}$ and (ii) $\gamma_{\textrm{inner}}$-weakly convex, where: $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ and $\gamma_{2,n}=\gamma_{\textrm{inner}}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
 From a practitioner’s perspective, we analyze the insights provided by Thm. (3.5) in three parts.
   • First, we have $\gamma_{2}\geq\gamma_{\textrm{inner}}$.
-Therefore, $\mathcal{GM}_{r}^{2}$ is less convex than $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, making it harder to optimize, and the same argument holds for their estimator.
+Therefore, $\mathrm{GMG}_{r}^{2}$ is less convex than $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$, making it harder to optimize, and the same argument holds for their estimator.
 In other words, we provably recover that, in practice, preserving the (scaled) distances is harder than simply preserving the angles.
@@
 In practice, we then choose $r$ with support as small as possible, precisely where we want $T$ to move points with minimal distortion.
-  • Third, and probably the most surprising, the finite sample GMG is more convex in high dimension.
-Indeed, $\gamma_{\textrm{inner},n}$ is the spectral width of $\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top}$, containing the (rescaled) inner-products between the $\mathbf{x}_{i}\sim r$.
+  • Third, and most surprising, the finite sample GMG is more convex in high dimension.
+Indeed, $\gamma_{\textrm{inner},n}$ is the spectral width of $\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top}$, which contains the (rescaled) inner products between $\mathbf{x}_{i}\sim r$.
 When $n>d$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})=0$ as $\mathrm{rank}(\mathbf{X}\mathbf{X}^{\top})=d$.
 Then, $\gamma_{\textrm{inner},n}$ increases, which in turn decreases the GMG’s convexity.
-On the other hand, when $d>n$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})>0$ if $\mathbf{X}$ is full rank.
-Intuitively, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is nearly convex when $\mathbf{X}\mathbf{X}^{\top}$ is well conditioned.
-Assuming that the $\mathbf{x}_{i}$ are normalized, this might happen in high dimension, as those points will be orthogonal with high probability.
-This property suggests that, in practice, and contrary to the insights provided by the statistical OT literature [71, 25, 56, 73], the GMG might not benefit a large sample size.
+However, when $d>n$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})>0$ if $\mathbf{X}$ is full rank.
+Intuitively, $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is nearly convex when $\mathbf{X}\mathbf{X}^{\top}$ is well conditioned.
+Assuming that the $\mathbf{x}_{i}$ are normalized, this might happen in high dimension, as they will be orthogonal with high probability.
+This suggests that, contrary to the insights provided by the statistical OT literature (Weed and Bach, 2017, Zhang et al., 2023), the GMG might not benefit a large sample size.
 3.5 Learning with the Gromov-Monge gap
 General Learning Procedure.
-Provided a source distribution $p$ and a target $q$ defining a marginal constraint, learning with the GMG remains to optimize a loss of the form
-$\mathcal{L}(\theta)\vcentcolon=\Delta(T_{\theta},p,q)+\lambda_{\textrm{GMG}}\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T_{\theta})$ (5)
-where $\Delta$ is a fitting loss, which can access paired, or unpaired, samples of $p$ and $q$.
-In theory, from Prop.
-(3.3), we can choose any reference $r$ s.t.
-$\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$.
-In practice, given the insights of Thm.
-(3.6), we usually consider $r=p$.
-Note that replacing $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$ by $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ in Eq.
-(5), we similarly define the learning procedure with the distortion.
-We compare their effect in Figure 1.
+Given a source $p$ and a target distribution $q$, we can use the GMG to guide a parameterized map $T_{\theta}$ towards approximating a Gromov-Monge map between $p$ and $q$.
+We handle the marginal constraint $T_{\theta}\sharp p=q$ separately through a fitting loss $\Delta(T_{\theta},p,q)$.
+Provided any reference $r$ s.t.
+$\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, we minimize
+$\min_{\theta}\Delta(T_{\theta},p,q)+\lambda\mathrm{GMG}_{r}(T_{\theta})$ (5)
+$\Delta$ can operate on paired (e.g., in VAE, the reconstruction loss), or unpaired (e.g., in VAE, the KL loss in VAE), samples of $p$ and $q$.
+Note that, in theory and as stated in Prop.
+(3.3), we can select any reference $r$ such that $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$.
+However, based on the insights from Thm.
+(3.4), we typically choose $r$ with minimal support size and, in practice, set $r=p$.
+This learning procedure is illustrated in Fig.
+2, where we also explore the effect of replacing $\mathrm{GMG}_{r}(T_{\theta})$ by $\mathrm{DST}_{r}(T_{\theta})$ in Eq.
+(5).
 VAE Learning Procedure.
-In the VAE setting, (i) when we apply the GMG (or the distortion) to the encoder $e_{\phi}$, the fitting loss is defined through the prior matching constraint, as described in § 2.1.
-Conversely, (ii) when we apply the GMG to the decoder $d_{\phi}$, the fitting loss is defined through the reconstruction loss.
-Additionally, in both cases, our goal is to promote the latent space to preserve certain geometric features of the data.
-Therefore, in (i) we use $r=p_{\textrm{data}}$ the data distribution as reference $r$, while in (ii) we use the latent distribution $r=q_{\phi}$.
-Introducing weightings $\lambda_{\textrm{enc}},\lambda_{\textrm{dec}}\geq 0$, determining which mapping we regularize, this remains to optimize the loss
-$\mathcal{L}_{\beta\textrm{-GMG}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{GM}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{GM}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (6)
-where $\mathcal{L}_{\beta}$ is introduce in § 2.1.
-Note that this loss can easily be extended to $\beta$-TCVAE and the combination of other regularization terms.
-While previous work [51, 41] chooses to apply the geometric regularizations to the decoder, we investigate regularizing both, separately or simultaneously.
-For completeness, we also derive the VAE-loss when learning with the distortion:
-$\mathcal{L}_{\beta\textrm{-DST}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{D}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{D}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (7)
-The choice of $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+In the VAE setting, we can use the GMG promote the (i) encoder $e_{\phi}$ or the (ii) decoder $d_{\theta}$ to mimic a Gromov-Monge map.
+In (i) we use $r_{e}=p_{\textrm{data}}$ the data distribution as reference $r$, while in (ii) we use the latent distribution $r_{d}=e_{\phi}\sharp p_{\textrm{data}}$.
+Introducing weightings $\lambda_{e},\lambda_{d}\geq 0$, determining which mapping we regularize, this remains to minimize
+$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda_{e}\mathrm{GMG}_{r_{e}}(e_{\phi})+\lambda_{d}\mathrm{GMG}_{r_{d}}(d_{\theta})}_{\textrm{(iii) geom.
+preservation}},$ (6)
+With this loss, prior matching and geometric desiderata can be efficiently combined, as terms (ii) and (iii) can simultaneously be 0.
+Note that this loss can be easily extended to more advanced prior matching objectives, such as the $\beta$-TCVAE loss (Chen et al., 2018), and can be combined with other regularizers, such as the HFS (Roth et al., 2023).
+We explore this strategy in experiments § 4.2.
 […]
 4 Experiments
-Table 1: Effect of different geometric regularization on disentanglement (DCI-D, Shapes3D [33]). We highlight the best method per regularization, and the best/second best per column.
+4.1 GMG vs. DST: Illustrative Example
+In Fig.
+2, we start by illustrating the difference between the DST and GMG on synthetic data.
+We train $T_{\theta}$ to map from a $3$D source $p$ to a $2$D target distribution $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019) and $\mathrm{R}$ a geometry-preserving regularizer.
+We compare three settings, [i] no regularization $\mathrm{R}=0$, [ii] $\mathrm{R}=\mathrm{DST}_{p}$ as regularizer , and [iii] $\mathrm{R}=\mathrm{GMG}_{p}$.
+For all cases, we plot the transported distribution $T_{\theta}\sharp p$ after training.
+Without regularization (third column), we fit the marginal constraint $T_{\theta}\sharp p=q$ but do not preserve the geometric features.
+With the DST (fourth column), we preserve geometric features but do not fit the marginal constraint, i.e., $T_{\theta}\sharp p\neq q$.
+On the other hand, with the GMG (fifth column), we get the best compromise by approximating a Gromov-Monge map: we fit $T_{\theta}\sharp p=q$, while preserving the geometric features as fully as possible.
+4.2 Leveraging the Gromov-Monge Gap for Disentanglement
+Experimental Setup.
+Having demonstrated how the DST enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning.
+Our primary goal is to investigate whether the GMG results in enhanced disentanglement compared to the DST by efficiently combining (i) prior matching with (ii) geometric constraints on the representation space.
+Moreover, we aim to determine which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose in the geometric regularization and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$.
+The experimental setup is described below.
+  • Baselines.
+We use the standard $\beta$-VAE and $\beta$-TCVAE as our starting models, with the option to apply the recent HFS regularization  (Roth et al., 2023) to each, resulting in a total of four base configurations.
+Note that the latter does not leverage geometric constraints; it is only used to enhance prior matching.
+We then investigate the effect of various geometry-preserving regularizations on disentanglement on top of these four base configurations.
+We consider GMG, DST and the Jacobian-based (Jac) regularization (Lee et al., 2022) discussed in 2.1.
+Given the inclusion of the DST, we naturally consider Nakagawa et al.
+(2023) as a baseline.
+  • Metrics.
+We evaluate the learned representations using the DCI-D (Eastwood and Williams, 2018) as it was found that it is the metric most suitable for measuring the disentanglement (Locatello et al., 2020, Dittadi et al., 2021).
+We report mean and standard deviation over 5 seeds.
+  • Datasets. We benchmark over four standard datasets: Shapes3D (Kim and Mnih, 2018), DSprites Higgins et al. (2017), SmallNORB LeCun et al. (2004), and Cars3D Reed et al. (2015).
+[…]
+Table 1: Effect of different regularization on disentanglement (DCI-D on Shapes3D).
+We highlight the best method per regularization type ($\mathbf{L2^{2}}$, $\mathbf{ScL2^{2}}$, or $\mathbf{Cos}$), and the best/second best per column.
 $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
 Base $65.8$$\pm 15.6$ $75.0$$\pm 3.4$ $88.1$$\pm 7.4$ $90.2$$\pm 7.5$
-Isometric (IR)
-+ (DST) 71.5 ±$\pm 3.6$ 75.8 ±$\pm 6.6$ 92.1 ±$\pm 9.7$ 90.9 ±$\pm 7.6$
-+ (GMG) 72.0 ±$\pm 12.5$ 78.9 ±$\pm 5.0$ 92.5 ±$\pm 4.4$ 91.7 ±$\pm 6.0$
-Scaled Isometric (SIR)
-+ Jacobian 61.4 ±$\pm 12.8$ 76.7 ±$\pm 4.5$ 90.5 ±$\pm 3.8$ 91.5 ±$\pm 5.6$
-+ (DST) 67.4 ±$\pm 7.1$ 77.9 ±$\pm 4.5$ 93.2 ±$\pm 9.7$ 94.5 ±$\pm 6.9$
-+ (GMG) 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
-Conformal (CR)
-+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
-+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
-Experimental setup.
-We evaluate the effectiveness of the (GMG) as regularizer in disentangled representation learning.
-We use the standard $\beta$-VAE and $\beta$-TCVAE as our base models and incorporate the GMG on top of them.
-Moreover, we consider the recently proposed HFS [60] regularization on top of both $\beta$-VAE and $\beta$-TCVAE, totaling four base models.
-Our primary goal is to investigate the differences between using the GMG and the (DST) as regularizers, specifically examining whether the (GMG) leads to more disentangled representations compared to the raw distortion.
-Additionally, we aim to determine which geometric regularization (IR, SIR, CR) is most beneficial for disentanglement and what part of the pipeline should be regularized.
-Lastly, we investigate whether a geometric regularization can help prevent the collapse of learned representation in the Decoder-free setting.
-We evaluate the learned latents with DCI-D [20] as it was found to be the metric most suitable to measure disentanglement [44, 18].
-We benchmark over multiple datasets commonly used to assess the performances of disentangled representation learning methods: Shapes3D [33], DSprites [27], SmallNORB [40], and Cars3D [57].
-4.1 Evaluating Different Geometric Regularizations
-Regularizing the Decoder.
-First, we focus on the difference between optimizing for different geometry-preserving regularizations.
-We compare between IR, SIR, and CR [41] realized through either the (DST), or (GMG).
-Additionally, we include the Jacobian-based SIR as introduced in Lee et al.
-[41].
-We report full results on Shapes3D [9] over 5 seeds in Table 1.
-We observe that the GMG outperforms the sole distortion loss for all levels of regularization and baselines.
-Moreover, we find that a CR performs best with respect to disentanglement compared to both IR and SIR.
-Note, that employing a CR has not been benchmarked for disentangled representation learning before.
-These results elucidate the clear benefit of using the GMG in its CR implementation in terms of learning more disentangled representations significantly improving upon previously proposed regularization.
-Table 2: Effect of (GMG) and (DST) leveraged as a conformal regularization (CR) on the disentanglement of learned representations as measured by DCI-D over four datasets.
-We highlight the best, and second best result for each dataset and method.
-CR $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
-Shapes3D [33]
-Base $65.8$$\pm 15.6$ $75.0$$\pm 3.4$ $88.1$$\pm 7.4$ $90.2$$\pm 7.5$
-+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
-+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
-DSprites [27]
+$\mathbf{L2^{2}}:c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$
++ Enc-DST $59.6$$\pm 6.9$ $75.7$$\pm 3.0$ $88.7$$\pm 7.1$ $90.3$$\pm 7.9$
++ Enc-GMG $62.3$$\pm 8.4$ $75.4$$\pm 5.3$ $88.4$$\pm 7.7$ $90.1$$\pm 4.3$
++ Dec-DST $71.5$$\pm 3.6$ $75.8$$\pm 6.6$ $92.1$$\pm 9.7$ $90.9$$\pm 7.6$
++ Dec-GMG $72.0$$\pm 8.5$ $78.9$$\pm 5.0$ $92.5$$\pm 4.4$ $91.7$$\pm 6.0$
+$\mathbf{ScL2^{2}}:c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$, $\alpha>0$ learnable
++ Jac $61.4$$\pm 12.8$ $76.7$$\pm 4.5$ $90.5$$\pm 3.8$ $91.5$$\pm 5.6$
++ Enc-DST $65.8$$\pm 11.9$ $73.0$$\pm 7.9$ $92.4$$\pm 3.7$ $89.2$$\pm 3.8$
++ Enc-GMG $65.1$$\pm 5.5$ $76.1$$\pm 7.7$ $90.8$$\pm 9.2$ $92.0$$\pm 5.3$
++ Dec-DST $67.4$$\pm 7.1$ $77.9$$\pm 4.5$ $93.2$$\pm 9.7$ $94.5$$\pm 6.9$
++ Dec-GMG 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
+[…]
++ Enc-DST $69.2$$\pm 9.1$ $77.2$$\pm 7.5$ $87.7$$\pm 7.7$ $90.5$$\pm 5.9$
++ Enc-GMG $70.9$$\pm 9.5$ $79.6$$\pm 6.6$ $92.5$$\pm 5.9$ $93.5$$\pm 6.9$
++ Dec-DST 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
++ Dec-GMG 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
+Which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ should we choose?
+The first question that naturally arises when using a geometry-preserving regularizer is: Which geometric features should be preserved?
+Previous works (Nakagawa et al., 2023, Lee et al., 2022, Huh et al., 2023) focused on preserving scaled distances between points, with the scale being learnable.
+We follow and extend this approach by also investigating plain distances and angles.
+This leads to three choices for $c_{\mathcal{X}},c_{\mathcal{Y}}$ ($\mathbf{L2^{2}},\mathbf{ScL2^{2}},\mathbf{Cos}$), as introduced in §3.1, following the hierarchy of geometry-preserving mappings proposed in Lee et al.
+(2022).
+We benchmark these on Shapes3D across various settings including DST, GMG, and Jac.
+Table 1 shows that angle preservation ($\mathbf{Cos}$), previously unconsidered for disentangled representation learning, consistently outperforms (scaled) distance preservation.
+This result is intuitive, as preserving angles imposes a weaker constraint, allowing for greater latent space expressiveness.
+In practice, preserving scaled distances seems to overly restrict the expressiveness of the latent space.
+Should we regularize the encoder $e_{\phi}$, or the decoder $d_{\theta}$?
+The next question we aim to answer is whether the decoder or encoder should be regularized.
+Therefore, we follow the previous setup on Shapes3D and benchmark all geometry-preserving regularizers on $e_{\phi}$ and $d_{\theta}$ as reported in Table 1.
+We find that regularizing the decoder is beneficial over regularizing the encoder.
+We hypothesize this is due to the regularization of $d_{\theta}$ offering a stronger signal as its gradients impact both $\phi$ and $\theta$, as in this case, the reference $r=e_{\phi}\sharp p_{\textrm{data}}$ is the distribution of encoded images.
+Our findings align with prior works (Lee et al., 2022, Nakagawa et al., 2023), which focus on regularizing $d_{\theta}$ yet do not offer this type of analysis.
+Additionally, we find that the GMG consistently outperforms the DST over all costs.
+Overall, the GMG on $d_{\theta}$ with $\mathbf{Cos}$ achieves best DCI-D results over all baselines.
+Consequently, moving forward we regularize the decoder for angle preservation.
+Table 2: Effect of GMG and DST with $\mathbf{Cos}$ as costs on disentanglement, as measured by DCI-D, over three datasets. We highlight the best, and second best result for each dataset and method.
+With $\mathbf{Cos}$ costs $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
+DSprites Higgins et al. (2017)
 Base $26.2$$\pm 18.5$ ${32.3}$$\pm 19.3$ $33.6$$\pm 17.9$ $48.7$$\pm 10.2$
-+ (DST) $\underline{28.6}$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
-+ (GMG) 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
-SmallNORB [40]
++ Dec-DST $\underline{28.6}$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
++ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
+SmallNORB LeCun et al. (2004)
 Base $26.8$$\pm 0.2$ $29.8$$\pm 0.4$ $26.8$$\pm 0.2$ $29.8$$\pm 0.4$
-+ (DST) $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$ $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$
-+ (GMG) 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$ 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$
-Cars3D [57]
++ Dec-DST $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$ $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$
++ Dec-GMG 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$ 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$
+Cars3D Reed et al. (2015)
 Base $29.6$$\pm 5.7$ $32.3$$\pm 4.6$ $29.6$$\pm 5.7$ $32.3$$\pm 4.6$
-+ (DST) $26.8$$\pm 3.6$ $33.7$$\pm 4.2$ $26.8$$\pm 3.6$ $33.7$$\pm 4.2$
-+ (GMG) 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
-Thus, next we benchmark the GMG in its CR form against its distortion counterpart across three more datasets again over four different baselines.
++ Dec-DST $26.8$$\pm 3.6$ $33.7$$\pm 4.2$ $26.8$$\pm 3.6$ $33.7$$\pm 4.2$
++ Dec-GMG 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
+GMG Consistently Enhances Disentanglement.
+To further validate our findings, we benchmark the GMG for decoder regularization with angle preservation ($\mathbf{Cos}$) against its distortion counterpart across three more datasets.
 We report full results in Table 2.
@@
 Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it.
-We emphasize that using the GMG as CR significantly improves results for all datasets versus not using any isometric regularization.
-This establishes the GMG as an effective regularization method beneficial for disentangled representation learning.
-Regularizing the Encoder.
-Lastly, we also analyze a CR on $e_{\phi}$, as well as regularizing both $d_{\theta}$ and $e_{\phi}$ together.
-We report full results over two datasets in Table 4.
-Again, the GMG on $d_{\theta}$ achieves best DCI-D over all baselines.
-This result is expected in the light of Theorem 3.7.
-Interestingly, regularizing solely $d_{\theta}$ outperforms regularizing both $e_{\phi}$ and $d_{\theta}$.
-We hypothesize this is due to the regularization of the decoder also offering a stronger signal as its gradients impact both the decoder and the encoder, as in this case, the reference $r$ is the distribution of encoded images.
-4.2 Towards Decoder-free Disentanglement
-Recently, works such as [11, 70, 21, 47, 1] have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation
-Table 3: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D [9].
-Decoder-free $\beta$-VAE $\beta$-TCVAE Base $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ Isometric (IR) + (DST) $38.2$$\pm 0.8$ $42.7$$\pm 1.6$ + (GMG) $13.9$$\pm 0.4$ $20.5$$\pm 0.5$ Scaled Isometric (SIR) + (DST) 45.6 ±$\pm 1.2$ 53.5 ±$\pm 1.0$ + (GMG) $15.2$$\pm 0.3$ $25.2$$\pm 0.6$ Conformal (CR) + (DST) $37.0$$\pm 0.4$ $46.1$$\pm 1.5$ + (GMG) $37.0$$\pm 0.9$ $38.8$$\pm 1.1$
-learning [14, 72, 3, 24]. However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
-As the goal of geometry preservation connects the data manifold and the latent domain through a minimal distortion objective and is applicable to both the encoder and decoder of a VAE (§3, Table 4), we posit that its application may provide sufficient training signal to learn meaningful representations and encourage disentanglement, eliminating the need for a reconstruction loss and decoder.
-Table 3 shows preliminary results on unsupervised decoder-free disentangled representation learning on the Shapes3D benchmark, where the decoder and associated reconstruction objective have been removed.
-Standard approaches such as $\beta$-VAE or $\beta$-TCVAE collapse and do not achieve measurable disentanglement (DCI-D of $0.0$).
-However, the inclusion of either DST or GMG significantly raises achievable disentanglement and, combined with the $\beta$-TCVAE matching objective, can achieve DCI-D scores of up to $53.5$ without needing any decoder or reconstruction loss.
-While these are preliminary insights, we believe they offer promise for more scalable approaches to unsupervised disentangled representation learning and potential bridges to popular and scalable self-supervised representation learning approaches.
-Note, that here the distortion loss significantly outperforms the GMG.
-This is expected due to the nature of the GMG, as the distortion loss offers a more restrictive and, thus, stronger signal for learning representations, which is necessary in the absence of a reconstruction objective.
-This highlights that while in most scenarios (§ 4.1, Figure 1), the GMG is preferable over the distortion loss, there also exist settings where a more restrictive optimization signal is desirable.
+We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets versus not using any isometric regularization.
+This establishes the GMG as an effective tool for enhanced disentanglement.
+Figure 3: Stability analysis for the DST and the GMG applied to the decoder $d_{\theta}$ with cost $\mathbf{Cos}$.
+For each regularizer, we assess stability by measuring the alignment of its gradients across 5 batches.
+We compute the cosine similarity between all pairs of gradients.
+Values closer to 1 indicates better alignment, hence better stability.
+Stability of the GMG.
+Furthermore, we investigate the stability of the GMG compared to the DST.
+Computing the GMG using Algorithm 1 involves solving an optimization problem through a GW solver.
+This experiment aims to demonstrate that our proposed method for solving the GW problem leads to a stable loss in the GMG .
+We assess the stability by measuring the alignment of GMG gradients $\nabla_{\theta}\mathrm{GMG}_{r_{n}}(T_{\theta})$ across 5 randomly sampled batches from each dataset $r_{n}\sim\mathcal{D}$, for a fixed neural map $T_{\theta}$.
+We repeat this procedure for each of the four datasets $\mathcal{D}$ considered and apply the same methodology to the DST.
+Figure 4 presents results for our best-performing configuration, which regularizes the decoder (i.e., $T_{\theta}=d_{\theta}$) and employs the $\mathbf{Cos}$ cost.
+Results for encoder regularization (i.e., $T_{\theta}=e_{\phi}$) are reported in Appendix D.1.
+We observe that the GMG’s gradients exhibit significantly higher alignment compared to those of the DST, demonstrating greater stability of our proposed regularizer.
+This suggests that the GMG ’s formulation, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
 5 Conclusion
-In this work, we introduce an optimal transport (OT) perspective on unsupervised disentangled representation learning to incorporate general latent geometrical constraints.
-We derive the Gromov-Monge gap (GMG), a provably weakly convex OT regularizer that measures the preservation of geometrical properties by a transport map $T$.
-By formulating disentangled representation learning as a transport problem, we integrate the GMG into standard training objectives, allowing for incorporating and studying various geometric constraints on the disentanglement of learned representation spaces.
-Including these geometry preserving regularization offers significant performance benefits across four standard disentanglement benchmarks when applied to existing disentanglement methods.
-Moreover, we show promising results on decoder-free unsupervised disentanglement.
-We demonstrate that optimizing for geometric constraints through the OT lens can provide sufficient training signal and regularization on the model encoder to achieve measurable disentanglement without explicit reconstruction objectives.
-This opens a possible door towards more scalable unsupervised disentanglement and bridges to weakly- & self-supervised encoder-only representation learning efforts.
+In this work, we introduce an OT perspective on unsupervised disentangled representation learning to incorporate general latent geometrical constraints.
+We derive the GMG, a provably weakly convex regularizer that measures whether a map $T$ transports a fixed reference distribution with minimal distortion of some predefined geometric features.
+By formulating disentangled representation learning as a transport problem, we integrate the GMG into standard training objectives, allowing for incorporating and studying various geometric constraints on the learned representation spaces.
+We show significant performance benefits of our approach on four standard disentanglement benchmarks.
+6 Reproduciblity
+In this work, we introduce the GMG, computed as detailed in Algorithm 1.
+To facilitate reproducibility, we provide the implementation code for computing the GMG on source and target batches in Appendix E.
+Comprehensive proofs, including all underlying assumptions, are presented in Appendix B.
+For our experiments on disentanglement benchmarks, we adhere to standard practices, employing streamlined preprocessing across all datasets.
+Detailed descriptions of these procedures are available in Appendix C.
+All experiments described in this paper can be conducted using a single RTX 2080TI GPU, ensuring accessibility and replicability of our results.
 Acknowledgements
@@
 Appendix
-Appendix A Proofs
-A.1 Positivity of the Entropic GMG
-Recall that
-$\begin{split}\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}(T)&\vcentcolon=\tfrac{1}{n}\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathcal{D}_{r_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
-For any coupling $\mathbf{P}\in U_{n}$, since $-\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
-$\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})<\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
-As a result, applying minimization on both sides yields that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({r}_{n},T\sharp{r}_{n})<\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}({r}_{n},T\sharp{r}_{n})$, and therefore:
-$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}(T)>\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}(T)=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0.$
-A.2 Reminders on Monge and Kantorovich OT
+The Appendix is organized as follows:
+  • Section A provides additional background information to supplement the main text.
+  • Section B presents all theoretical proofs, including detailed assumptions.
+  • Section C outlines comprehensive experimental details, ensuring reproducibility.
+  • Section D offers supplementary empirical results that further support our findings.
+  • Section E includes the implementation code for computing the GMG.
+Appendix A Additional Background
+A.1 Reminders on Monge and Kantorovich OT
 In this section, we recall the Monge and Kantorovich formulations of OT, which we will use to prove various results.
@@
 This assumes that we have a meaningful way to compare elements $\mathbf{x},\mathbf{y}$ from the source and target domains.
-The Monge [49] problem (MP) between $p\in\mathcal{P}(\mathcal{X})$ and $p\in\mathcal{P}(\mathcal{Y})$ consists of finding a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $p$, while minimizing the average displacement cost quantified by $c$
+The Monge (1781) problem (MP) between $p\in\mathcal{P}(\mathcal{X})$ and $p\in\mathcal{P}(\mathcal{Y})$ consists of finding a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $p$, while minimizing the average displacement cost quantified by $c$
 $\inf_{T:T\sharp p=p}\int_{\mathcal{X}}c(\mathbf{x},T(\mathbf{x}))\mathop{}\!\mathrm{d}p(\mathbf{x})\,.$ (MP)
@@
 Instead of transport maps, the Kantorovich problem (KP) seeks a couplings $\pi\in\Pi(p,q)$:
-$\mathrm{W}_{c}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\,.$ (KP)
+$\mathrm{W}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\,.$ (KP)
 An optimal coupling $\pi^{\star}$ solution of (KP), always exists.
@@
 Indeed, when (MP) is feasible, the Monge and Kantorovich formulations coincide and $\pi^{\star}=(\mathrm{Id},T^{\star})\sharp p$.
-A.3 Conditionally Positive Kernels
+A.2 Conditionally Positive Kernels
 In this section, we recall the definition of a conditionally positive kernel, which is involved in multiple proofs relying on the linearization of the Gromov-Wasserstein problem as a Kantorovich problem.
 Definition A.1.
-A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is conditionally positive if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t.
+A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is CPD, i.e., conditionally positive, if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t.
 $\mathbf{a}^{\top}\mathbf{1}_{n}=0$, one has
 $\sum_{i,j=1}^{n}\mathbf{a}_{i}\mathbf{a}_{j}\,k(\mathbf{x}_{i},\mathbf{x}_{j})\geq 0$
-Conditionally positive kernels include all positive kernels, such as the inner-product $k(\mathbf{x},\mathbf{y})=\langle\mathbf{x},\mathbf{y}\rangle$, the cosine similarity $k(\mathbf{x},\mathbf{y})=\textrm{cos-sim}(\mathbf{x},\mathbf{y})=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{y}}{\|\mathbf{y}\|_{2}}\rangle$, but also the negative squared Euclidean distance $k(\mathbf{x},\mathbf{y})=-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$.
-Therefore, each of the costs of interest is either a conditionally positive kernel - for the inner product and the cosine distance - or its opposite is - the squared Euclidean distance.
-Appendix B Proofs of § 3.2
+CPD include all positive kernels, such as the inner-product $k(\mathbf{x},\mathbf{x}^{\prime})=\langle\mathbf{x},\mathbf{x}^{\prime}\rangle$, or the cosine similarity $k(\mathbf{x},\mathbf{x}^{\prime})=\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{x}^{\prime}}{\|\mathbf{x}^{\prime}\|_{2}}\rangle$, but also the (scaled) negative squared Euclidean distance $k(\mathbf{x},\mathbf{x}^{\prime})=-\alpha\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$, $\alpha>0$.
+Therefore, each of the costs of interest is either a conditionally positive kernel - for the inner product and the cosine distance - or its opposite is - for the squared Euclidean distance.
+Additionally, CPD kernels also include more exotic cost functions, such as:
+  • The RBF kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/\gamma),\gamma>0$.
+  • The power kernel $k(\mathbf{x},\mathbf{x}^{\prime})=-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{p},0<p<2$.
+  • The thin plate spline kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}\log(\|\mathbf{x}-\mathbf{x}^{\prime})$
+  • The inverse multi-quadratic kernel $k(\mathbf{x},\mathbf{x}^{\prime})=1/\sqrt{\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}+c^{2}},c\in\mathbb{R}$.
+As a result, the family of CPD kernels includes a large variety of cost functions, which can be used to define various GMGs.
+Appendix B Proofs
+B.1 The GMG characterizes Gromov-Monge Optimality
+We show here that if $\mathrm{GMG}_{r}(T)=0$, then $T$ is a Gromov-Monge map between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+As the set of deterministic couplings $\{\pi_{F}\vcentcolon=(\mathrm{I}_{d},F)\sharp r|F:\mathcal{X}\to\mathcal{Y}\,,F\sharp r=T\sharp r\}\subset\Pi(r,T\sharp r\}$, we immediately get that
+$\inf_{F\sharp r=T\sharp r}\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(F(\mathbf{x}),F(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\geq\mathrm{GW}(r,T\sharp r)$ (7)
+On the other hand, if $\mathrm{GMG}_{r}(T)=0$, one has
+$\mathrm{GW}(r,T\sharp r)=\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})$ (8)
+Therefore, combining Eq. (7) and Eq. (8), we get
+$\begin{split}&\inf_{F\sharp r=T\sharp r}\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(F(\mathbf{x}),F(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\\ =&\int_{\mathcal{X}\times\mathcal{X}}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\end{split}$ (9)
+Finally, as $T$ naturally satisfies the marginal constraint, we conclude that $T$ is a Gromov-Monge map between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$.
+B.2 On rescaling the costs matrices in the entropic GW solver
+We remind, from Eq. (EGWP), that
+$\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(\mathbf{C}_{\mathcal{X}_{i,i^{\prime}}}-\mathbf{C}_{\mathcal{Y}_{j,j^{\prime}}})^{2}\,\mathbf{P}_{i,j}\mathbf{P}_{i^{\prime},j^{\prime}}-\varepsilon H(\mathbf{P})\,.$ (10)
+By developping each terms, and using the fact that $\mathbf{P}\in U_{n}$, we get
+$\begin{split}\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})&=\min_{\mathbf{P}\in U_{n}}\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{X}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{Y}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\\ &=\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{X}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\tfrac{1}{n^{2}}\langle\mathbf{C}_{\mathcal{Y}}^{\odot 2}\mathbf{1}_{n},\mathbf{1}_{n}\rangle+\min_{\mathbf{P}\in U_{n}}-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\end{split}$ (11)
+where $\mathbf{C}_{\mathcal{X}}^{\odot 2}=\mathbf{C}_{\mathcal{X}}\odot\mathbf{C}_{\mathcal{X}}$, with $\odot$ the Hadamard (i.e., elementwise) product, and similarly for $\mathbf{C}_{\mathcal{Y}}^{\odot 2}$.
+As we can see that the two terms on the left do not depend on $\mathbf{P}$, they do not impact the minimization, an OT coupling $\mathbf{P}^{\star}$ solving the problem satisfies:
+$\begin{split}\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,.\end{split}$ (12)
+As a result, if we now replace $\mathbf{C}_{\mathcal{X}}$ and $\mathbf{C}_{\mathcal{Y}}$ by $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$, respectively, the new OT coupling $\tilde{\mathbf{P}}^{\star}$ solving the problem satisfies
+$\begin{split}&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\frac{\mathbf{C}_{\mathcal{X}}}{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\mathbf{P}\frac{\mathbf{C}_{\mathcal{Y}}}{\texttt{stat}(\mathbf{C}_{\mathcal{Y}})},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-\frac{2}{{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})}\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})\cdot\varepsilon H(\mathbf{P})\,,\end{split}$ (13)
+where in the last line, we use the fact that $\texttt{stat}(\mathbf{C}_{\mathcal{X}}),\texttt{stat}(\mathbf{C}_{\mathcal{Y}})>0$.
+This yields the desired equivalence on scaling the cost matrix and adapting the entropic regularization strength.
+B.3 Positivity of the Entropic GMG estimator
+Recall that
+$\begin{split}\mathrm{GMG}_{{r}_{n},\varepsilon}(T)&=\mathrm{DST}_{r_{n}}(T)-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathrm{DST}_{r_{n}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
+For any coupling $\mathbf{P}\in U_{n}$, since $-\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
+$\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})<\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
+As a result, applying minimization on both sides yields that $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})<\mathrm{GW}_{0}({r}_{n},T\sharp{r}_{n})=\mathrm{GW}({r}_{n},T\sharp{r}_{n})$, and therefore:
+$\mathrm{GMG}_{r_{n},\varepsilon}(T)>\mathrm{GMG}_{r_{n},0}(T)=\mathrm{GMG}_{r_{n}}(T)\geq 0.$
+B.4 Proofs of Prop. 3.3
 See 3.3
@@
 Then, $\pi^{r}:=(\mathrm{Id},T)\sharp r$ is an optimal Gromov-Wasserstein coupling, solution of Problem (GWP) between $r$ and $T\sharp r$ for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$.
-Therefore, from [65, Theorem.
-3], $\pi^{r}$ is an optimal Kantorvich coupling, solution of Problem (KP) between $r$ and $T\sharp r$ for the linearized cost:
+Therefore, from (Séjourné et al., 2023, Theorem.
+3), $\pi^{r}$ is an optimal Kantorvich coupling, solution of Problem (KP) between $r$ and $T\sharp r$ for the linearized cost:
 $\tilde{c}:(\mathbf{x},\mathbf{y})\in\mathcal{X}\times\mathcal{Y}\mapsto\int_{\mathcal{X}\times\mathcal{Y}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi^{r}(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ (14)
@@
 Afterward, since $\pi^{r}$ has finite mass, by Lebesgue’s dominated convergence Theorem, it follows that $\tilde{c}$ is continuous, and hence uniformly continuous, again since $\mathcal{X}\times\mathcal{Y}$ is compact.
-Afterwards, by virtue of [61, Theorem 1.38], $\mathrm{Spt}\left(\pi^{r}\right)$ is a $\tilde{c}$-cyclically monotone (CM) set (see [61, Definition.
-1.36]).
+Afterwards, by virtue of (Santambrogio, 2015, Theorem 1.38), $\mathrm{Spt}\left(\pi^{r}\right)$ is a $\tilde{c}$-cyclically monotone (CM) set (see (Santambrogio, 2015, Definition.
+1.36)).
 From the definition of cyclical monotonicity, this property translates to subsets.
 Then, by defining $\pi^{s}=(\mathrm{Id},T)\sharp s$, as $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, one has $\mathrm{Spt}(\pi^{s})=\mathrm{Spt}((\mathrm{Id},T)\sharp s)\subset\mathrm{Spt}((\mathrm{Id},T)\sharp r)=\mathrm{Spt}(\pi^{r})$, so $\mathrm{Spt}(\pi^{s})$ is $\tilde{c}$-CM.
-Finally, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $\tilde{c}$ is uniformly continuous, the $\tilde{c}$-cyclical monotonicity of its support implies that the coupling $\pi^{p}$ is a Kantorovich optimal coupling between its marginals for cost $\tilde{c}$, thanks to [61, Theorem 1.49].
-By re-applying [65, Theorem.
-3], we get that $\pi^{s}$ solves the Gromov-Wasserstein problem between its marginals for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$.
-In other words, $\pi^{s}=(\mathrm{Id},T)\sharp s$ is Gromov-Wasserstein optimal coupling between $s$ and $T\sharp s$ so $T$ is a Gromov-Monge map between $s$ and $T\sharp s$ and $\mathcal{GM}_{s}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$.
+Finally, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $\tilde{c}$ is uniformly continuous, the $\tilde{c}$-cyclical monotonicity of its support implies that the coupling $\pi^{p}$ is a Kantorovich optimal coupling between its marginals for cost $\tilde{c}$, thanks to (Santambrogio, 2015, Theorem 1.49).
+By re-applying (Séjourné et al., 2023, Theorem.
+3), we get that $\pi^{s}$ solves the Gromov-Wasserstein problem between its marginals for costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$.
+In other words, $\pi^{s}=(\mathrm{Id},T)\sharp s$ is Gromov-Wasserstein optimal coupling between $s$ and $T\sharp s$ so $T$ is a Gromov-Monge map between $s$ and $T\sharp s$ and $\mathrm{GMG}_{s}(T)=0$.
 ∎
-Appendix C Proofs of § 3.3
-See 3.4
+B.5 Proofs of Thm. 3.5
+See 3.5
+Before proving Thm. 3.5, we first demonstrate some technical results that will be useful later.
+Reformulation of the empirical GMG using permutations.
+We start by showing that $\mathrm{GMG}_{{r}_{n}}(T)$ is always the sub-optimality gap of $T$ in Prob.
+(GMP) between $r_{n}$ and $T\sharp r_{n}$.
+This occurs because Prob.
+(GMP) and Prob.
+(GWP) coincide when applied between empirical measures on the same number of points.
+In other words, we can reformulate Prob.
+(GWP) between $r_{n}$ and $T\sharp r_{n}$ using permutation matrices, instead of (plain) couplings.
+Proposition B.1.
+The empirical GMG reads
+$\begin{split}\mathrm{GMG}_{{r}_{n}}(T)=\mathrm{DST}_{{r}_{n}}(T)-\min_{\sigma\in\mathcal{S}_{n}}\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\left(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{\sigma(i)}),T(\mathbf{x}_{\sigma(j)}))\right)^{2}\end{split}$ (15)
 Proof.
-We start by showing a more general results, stating that when $c_{\mathcal{X}},c_{\mathcal{Y}}$ are conditionally positive kernels (see A.1), the discrete GW couplings between uniform, empirical distributions supported on the same number of points, ae permutation matrices.
-Proposition C.1 (Equivalence between Gromov-Monge and Gromov-Wasserstein problems in the discrete case.).
+We first show a more general results, stating that when $c_{\mathcal{X}},c_{\mathcal{Y}}$ are conditionally positive kernels (see A.1), the discrete GW couplings between uniform, empirical distributions supported on the same number of points, ae permutation matrices.
+Proposition B.2 (Equivalence between Gromov-Monge and Gromov-Wasserstein problems in the discrete case.).
 Let ${p}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and ${q}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{y}_{i}}$ two uniform, empirical measures, supported on the same number of points.
@@
 Then, the GM and GW formulations coincide, in the sense that we can restrict the GW problem to permutations, namely
-$\begin{split}\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({p}_{n},{p}_{n})&=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\sigma\in\mathcal{S}_{n}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{\sigma(i)},\mathbf{y}_{\sigma(j)}))^{2}\end{split}$ (9)
+$\begin{split}\mathrm{GW}({p}_{n},{p}_{n})&=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\\ &=\tfrac{1}{n^{2}}\min_{\sigma\in\mathcal{S}_{n}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{\sigma(i)},\mathbf{y}_{\sigma(j)}))^{2}\end{split}$ (16)
 Proof.
@@
 that always exists by continuity of the GW objective function on the compact $U_{n}$. We show that $\mathbf{P}^{\star}$ can be chosen as a (rescaled) permutation matrix without loss of generality.
-As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from [65, Theorem. 3], $\mathbf{P}^{\star}$ also solves:
+As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem.
+3), $\mathbf{P}^{\star}$ also solves:
 $\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{Q}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})-c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}}))^{2}\mathbf{P}^{\star}_{ij}\mathbf{Q}_{i^{\prime}j^{\prime}}$ (17)
@@
 $\mathbf{P}^{\star}\in\operatorname*{arg\,min}_{\mathbf{Q}\in U_{n}}\langle\tilde{\mathbf{C}},\mathbf{Q}\rangle$ (18)
-Birkhoff’s theorem states that the extremal points of $U_{n}$ are equal to the permutation matrices $P_{n}$.
-Moreover, a seminal theorem of linear programming [6, Theorem 2.7] states that the minimum of a linear objective on a bounded polytope, if finite, is reached at an extremal point of the polyhedron.
+Birkhoff’s theorem states that the extremal points of $U_{n}$ are the permutation matrices $P_{n}$.
+Moreover, a seminal theorem of linear programming (Bertsimas and Tsitsiklis, 1997, Theorem 2.7) states that the minimum of a linear objective on a bounded polytope, if finite, is reached at an extremal point of the polyhedron.
 Therefore, as $\mathbf{P}^{\star}$ solves Eq.
@@
 Therefore, the equivalence between GW and GM follows.
-∎
-To conclude the proof of Prop. 3.4, we simply remark that:
-  • $r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $T\sharp r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$ are uniform, empirical distribution, and supported on the same number of points;
-  • The costs of interests [i] or [ii] are either conditionally positive, or their opposite is, as detailed below  Def (A.1).
-See 3.5
+To conclude the proof of Prop.
+B.1, we simply remark that $r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $T\sharp r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$ are uniform, empirical distribution, and supported on the same number of points.
+Consistency of the empirical GMG.
+We continue by proving a consistency result for the empirical GMG, which we will later use to deduce the asymptotic weak convexity constant from the finite-sample case.
+Proposition B.3.
+For both $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$ and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, one has $\mathrm{GMG}_{{r}_{n}}(T)\rightarrow\mathrm{GMG}_{r}(T)$ almost surely.
 Proof.
-We first note that the empirical estimator of the distortion is consistent, as both costs [i] or [ii] are continuous, and $\mathcal{X}$ is compact.
-We then need to study, in both cases, the convergence of $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r_{n})$ to $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r)$.
+We first note that the empirical estimator of the distortion is consistent, as both costs are continuous, and $\mathcal{X}$ is compact.
+We then need to study, in both cases, the convergence of $\mathrm{GW}(r_{n},T\sharp r_{n})$ to $\mathrm{GW}(r_{n},T\sharp r)$.
 To that end, we first remark that as, almost surely, $r_{n}\to r$ in distribution, one also has that, almost surely, $T\sharp r_{n}\to T\sharp r$ in distribution.
@@
 Afterwards, one can simply adapt the proof of the almost sure weak convergence of empirical measure based on the strong law of large numbers to show that, almost surely, $T\sharp{r}_{n}\rightarrow T\sharp r$ in distribution.
-See for instance [39, Theorem 10.4.1].
-[i] We start by the (scaled) squared euclidean distances.
-Up to replacing $r$ by $\alpha^{2}\sharp r$ and $T$ by $T\circ(\tfrac{1}{\alpha^{2}})$, and similarly for $r_{n}$, we can assume without loss of generality that $\alpha=1$.
-As, almost surely, both $r_{n}\to r$ and $T\sharp{r}_{n}\to T\sharp r$ in distribution, the results follows from [48, Thm 5.1, (e)].
-[ii] We continue with the cosine similarity.
-To that end, we first consider the inner product, i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, and show that if $p_{n}\to p$ and $q_{n}\to q$ in distribution, then $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$.
-As noticed by Rioux et al.
-[58, Lemma 2]–in the first version of the paper– the GW for inner product costs can be reformulated as:
+See for instance (Le Gall, , Theorem 10.4.1).
+We start with the squared Euclidean distance. As, almost surely, both $r_{n}\to r$ and $T\sharp{r}_{n}\to T\sharp r$ in distribution, the results follows from (Mémoli, 2011, Thm 5.1, (e)).
+We continue with the inner product. As noticed by Rioux et al. (2023, Lemma 2)–in the first version of the paper – the GW for inner product costs can be reformulated as:
 $\begin{split}\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)=&\int_{\mathcal{X}\times\mathcal{X}}\langle\mathbf{x},\mathbf{x}^{\prime}\rangle\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})+\int_{\mathcal{Y}\times\mathcal{Y}}\langle\mathbf{y},\mathbf{y}^{\prime}\rangle\mathop{}\!\mathrm{d}q(\mathbf{y})\mathop{}\!\mathrm{d}q(\mathbf{y}^{\prime})\\ +&\min_{\mathbf{M}\in\mathcal{M}}\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+4\|\mathbf{M}\|_{2}^{2}\,,\end{split}$ (19)
@@
 (KP) induced by $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$.
-By virtue of [46, Theorem 2], there exists a constant $C>0$, s.t.
+By virtue of (Manole and Niles-Weed, 2024, Theorem 2), there exists a constant $C>0$, s.t.
 we can uniformly bound
@@
 As a result, we finally get $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$.
-Finally, we remark that for any $p,q$, $\mathrm{GW}^{\textrm{cos-sim}}(p,q)=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)$, where $\mathrm{proj}_{S^{d-1}}(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$.
-Using similar arguments invoked previously, as $p_{n}\to p$ in distribution $p$, $\mathrm{proj}_{S^{d-1}}\sharp p_{n}\to\mathrm{proj}_{S^{d-1}}\sharp p$ in distribution, and similarly $\mathrm{proj}_{S^{d-1}}\sharp q_{n}\to\mathrm{proj}_{S^{d-1}}\sharp q$ in distribution.
-As a result:
-$\begin{split}\mathrm{GW}^{\textrm{cos-sim}}(p_{n},q_{n})&=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p_{n},\mathrm{proj}_{S^{d-1}}\sharp q_{n})\\ &\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)\\ &=\mathrm{GW}^{\textrm{cos-sim}}(p,q)\end{split}$ (15)
-which yields the desired convergence by using $p_{n}=r_{n}$ and $q_{n}=T\sharp r_{n}$.
-Appendix D Proofs of § 3.4
-See 3.7
-We start by recalling the standard definition of weakly convex function on $\mathbb{R}^{d}$, along with technical lemmas that we will in the proof of Thm. (3.7).
-Definition D.1.
+Weak convexity.
+Finally, we demonstrate some useful results on weakly convex functions on $\mathbb{R}^{d}$.
+Definition B.4.
 A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\gamma$-weakly convex if $f+\gamma\|\cdot\|_{2}^{2}$ is convex.
-Lemma D.2.
+From the definition, we see that if $f$ is $\gamma$-weakly convex, than $f$ is also $\gamma^{\prime}$ weakly convex for any $\gamma^{\prime}\geq\gamma$.
+This naturally extends to weakly convex functionals $\mathcal{F}$ on $L_{2}(r)$.
+Lemma B.5.
 Let $\mathbf{A}\in S_{d}(\mathbb{R})$ a symmetric matrix and define the quadratic form $f_{\mathbf{A}}:\mathbf{x}\in\mathbb{R}^{d}\mapsto\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$.
@@
 We use the fact that a twice continuously differentiable function is convex i.f.f.
-its hessian is positive semi-definite [8, §(3.1.4)].
+its hessian is positive semi-definite (Boyd and Vandenberghe, 2004, §(3.1.4)).
 Therefore, $f_{\mathbf{A}}$ is convex i.f.f.
@@
 ∎
-Lemma D.3.
+Lemma B.6.
 Let $(f_{i})_{i\in I}$ a family of $\gamma$-weakly convex functions, with potentially infinite $I$.
@@
 Proof.
-As the $f_{i}$ are $\gamma$-weakly convex, $f_{i}+\tfrac{1}{2}\gamma$ is convex, so $\mathbf{x}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})+\tfrac{1}{2}\gamma\|x\|_{2}^{2}=(\sup_{i\in I}f_{i}(\mathbf{x}))+\tfrac{1}{2}\gamma\|x\|_{2}^{2}$ is convex [8, Eq.
-(3.7)].
+As the $f_{i}$ are $\gamma$-weakly convex, $f_{i}+\tfrac{1}{2}\gamma$ is convex, so $\mathbf{x}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})+\tfrac{1}{2}\gamma\|x\|_{2}^{2}=(\sup_{i\in I}f_{i}(\mathbf{x}))+\tfrac{1}{2}\gamma\|x\|_{2}^{2}$ is convex (Boyd and Vandenberghe, 2004, Eq.
+(3.7)).
 Therefore, the $\gamma$-weak convexity of $f$ follows ∎
+Let’s now proceed to prove the main Thm. (3.5).
 Proof of Thm. (3.5)..
 Finite sample. We first study the weak convexity of $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$, i.e. the Gromov-Monge gap for the inner product. For a map $T\in L_{2}(r)$, it reads
-$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
-As ${r}_{n}$ and $T\sharp{r}_{n}$ are uniform empirical supported on the same number of points, using Prop. C.1, we can reformulate the RHS with permutation matrices, which yields
-$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
-From this expression, $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ can be reformulated as a matrix input function.
+$\begin{split}\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
+As ${r}_{n}$ and $T\sharp{r}_{n}$ are uniform empirical supported on the same number of points, using Prop. B.2, we can reformulate the RHS with permutation matrices, which yields
+$\begin{split}\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
+From this expression, $\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ can be reformulated as a matrix input function.
 Indeed, it only depends on the map $T$ via its values on the support of ${r}_{n}$, namely $\mathbf{x}_{1},...,\mathbf{x}_{n}$.
 Therefore, we write $\mathbf{t}_{i}\vcentcolon=T(\mathbf{x}_{i})$, and define $\mathbf{X},\mathbf{T}\in\mathbb{R}^{n\times d}$ which contain observations $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ respectively, stored as rows.
-Then, studying $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ remains to study
+Then, studying $\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ remains to study
 $f(\mathbf{T})\vcentcolon=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle\mathbf{t}_{i},\mathbf{t}_{j}\rangle|^{2}-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle\mathbf{t}_{j},\mathbf{t}_{j^{\prime}}\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
@@
 To study the convexity of this matrix input function, we vectorize it.
-From [53, Eq.
-(520)], we note that, for any $\mathbf{M}\in\mathbb{R}^{n\times n}$
+From (Petersen and Pedersen, 2008, Eq.
+(520)), we note that, for any $\mathbf{M}\in\mathbb{R}^{n\times n}$
 $\begin{split}\langle\mathbf{M}\mathbf{T},\mathbf{T}\rangle=\mathbf{vec}(\mathbf{T})^{\top}\mathbf{vec}(\mathbf{M}\mathbf{T})=\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{M}\otimes I_{n})\mathbf{vec}(\mathbf{T})\end{split}$
@@
 This remains to study the (semi-) positive definiteness of the matrices $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$.
-As each $\mathbf{A}_{\mathbf{X},\mathbf{P}}\in\mathbb{R}^{n\times n}$ is symmetric and square, $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is also symmetric and from [53, Eq.
-(519)] its eigenvalues are the outer products of the eigenvalues of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$ and $I_{n}$, namely
+As each $\mathbf{A}_{\mathbf{X},\mathbf{P}}\in\mathbb{R}^{n\times n}$ is symmetric and square, $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is also symmetric and from (Petersen and Pedersen, 2008, Eq.
+(519)) its eigenvalues are the outer products of the eigenvalues of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$ and $I_{n}$, namely
 $\begin{split}\mathbf{eig}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})&=\{\lambda_{i}(\mathbf{A}_{\mathbf{X},\mathbf{P}})\cdot\lambda_{j}(I_{n})\}_{1\leq i,j\leq n}\\ &=\{\underbrace{\lambda_{1}(\mathbf{A}_{\mathbf{X},\mathbf{P}}),\dots,\lambda_{1}(\mathbf{A}_{\mathbf{X},\mathbf{P}})}_{\textrm{n times}},\dots,\underbrace{\lambda_{n}(\mathbf{A}_{\mathbf{X},\mathbf{P}}),\dots,\lambda_{n}(\mathbf{A}_{\mathbf{X},\mathbf{P}})}_{\textrm{n times}}\}\end{split}$ (23)
@@
 We then remind that each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}$ is the quadratic form defined by $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$, so by applying Prop.
-D.2, it is $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$-weakly convex, and hence $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex.
+B.5, it is $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$-weakly convex, and hence $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex.
 Therefore, applying Prop.
-(D.3), $r$ is $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex, in $\mathbb{R}^{d}$.
+(B.6), $r$ is $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex, in $\mathbb{R}^{d}$.
 Reminding that $\gamma_{\mathrm{inner}}=\tfrac{1}{n}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$, $r$ is $\tfrac{1}{n}\gamma_{\mathrm{inner}}$ weakly convex.
@@
 By reminding that $\mathbf{T}$ stores the $T(\mathbf{x}_{i})$ as rows, $\tfrac{1}{n}\|\mathbf{T}\|_{2}^{2}=\|T\|_{L_{2}(r_{n})}$.
-Consequently, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is $\gamma_{\mathrm{inner}}$ in $L_{2}(r_{n})$.
-We then study the convexity of $\mathcal{GM}_{r_{n}}^{2}$. We follow exactly the same approach. One has:
-$\begin{split}\mathcal{GM}^{2}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
-Similarly, studying the convexity of $\mathcal{GM}^{2}_{{r}_{n}}(T)$ remains to study the convexity of the matrix input function:
+Consequently, $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is $\gamma_{\mathrm{inner}}$ in $L_{2}(r_{n})$.
+We then study the convexity of $\mathrm{GMG}_{r_{n}}^{2}$. We follow exactly the same approach. One has:
+$\begin{split}\mathrm{GMG}^{2}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
+Similarly, studying the convexity of $\mathrm{GMG}^{2}_{{r}_{n}}(T)$ remains to study the convexity of the matrix input function:
 $\begin{split}g(\mathbf{T})&\vcentcolon=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|\mathbf{t}_{i}-\mathbf{t}_{j}\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|\mathbf{t}_{i}-\mathbf{t}_{j}\|_{2}^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
@@
 $\begin{split}\lambda_{\min}(\mathbf{B}_{\mathbf{X},\mathbf{P}})&=\lambda_{\min}\left(\tfrac{1}{2n}(\tfrac{1}{2}(\mathbf{P}^{\top}+\mathbf{P})-I_{n})\otimes\mathbf{D}_{\mathbf{X}}\right)\\ &\geq\lambda_{\min}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)+\lambda_{\min}\left(-\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\\ &\geq\lambda_{\min}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)-\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\end{split}$ (26)
-For both terms, we apply again [53, Eq. (519)]. For the LHS, one has:
+For both terms, we apply again (Petersen and Pedersen, 2008, Eq. (519)). For the LHS, one has:
 $\mathbf{eig}\left(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P})\otimes\mathbf{D}_{\mathbf{X}}\right)=\{\lambda_{i}(\tfrac{1}{4n}(\mathbf{P}^{\top}+\mathbf{P}))\lambda_{j}(\mathbf{D}_{\mathbf{X}})\}_{1\leq i,j\leq n}$ (27)
@@
 As a result, since convexity is preserved under pointwise convergence and by virtue of Prop.
-(3.5), we study the (almost sure) convergence of $\gamma_{\textrm{inner},n}$ and $\gamma_{2,n}$.
+(B.3), we study the (almost sure) convergence of $\gamma_{\textrm{inner},n}$ and $\gamma_{2,n}$.
 We start by $\gamma_{\textrm{inner},n}$.
@@
 We first remark that $\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}\leq\sup_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
-As a result, by defining $\tilde{\gamma}_{2,n}=\gamma_{\textrm{inner},n}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$, $\mathcal{GM}^{2}_{{r}_{n}}$ is also $\tilde{\gamma}_{2,n}$-weakly convex.
+As a result, by defining $\tilde{\gamma}_{2,n}=\gamma_{\textrm{inner},n}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$, $\mathrm{GMG}^{2}_{{r}_{n}}$ is also $\tilde{\gamma}_{2,n}$-weakly convex.
 Moreover, $\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ does not depends on $n$, $\tilde{\gamma}_{2,n}\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ almost surely, which also provides the desired asymptotic result.
-Appendix E Additional Empirical Results
-Table 4: Disentanglement of regularizing the Encoder and the Encoder and Decoder as measured by DCI-D on two different datasets.
-We highlight best, second best, and third best results for each method and dataset.
-DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
-Shapes3D [33]
-Base $67.7$$\pm 7.8$ $75.6$$\pm 8.7$ $88.1$$\pm 7.4$ $89.5$$\pm 7.9$
-+ Enc-(DST) $69.2$$\pm 9.1$ $77.2$$\pm 7.5$ $87.7$$\pm 7.7$ $90.5$$\pm 5.9$
-+ Enc-(GMG) $70.9$$\pm 9.5$ 79.6 ±$\pm 6.6$ 92.5 ±$\pm 5.9$ 93.5 ±$\pm 6.9$
-+ Dec-(DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ 91.9 ±$\pm 9.4$
-+ Dec-(GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
-+ Enc-Dec-(GMG) 72.8 ±$\pm 7.7$ $79.3$$\pm 13.9$ 93.3 ±$\pm 5.0$ $91.8$$\pm 7.3$
-DSprites [27]
-Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
-+ Enc-(DST) 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
-+ Enc-(GMG) $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
-+ Dec-(DST) $28.6$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
-+ Dec-(GMG) 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
-+ Enc-Dec-(GMG) 33.1 ±$\pm 14.9$ 40.2 ±$\pm 7.0$ $28.7$$\pm 14.6$ $46.0$$\pm 11.3$
-Appendix F Experimental Details
-All our experiments build on python 3 and the jax-framework [2], alongside ott-jax for optimal transport utilities.
-To effectively conduct comprehensive and representative research on disentangled representation learning, we convert the public PyTorch framework proposed in  [60] to an equivalent jax variant.
+Appendix C Experimental Details
+All our experiments build on python and the jax-framework (Babuschkin et al., 2020), alongside ott-jax for optimal transport utilities.
+C.1 Details on Disentanglement Benchmark
+To effectively conduct comprehensive and representative research on disentangled representation learning, we convert the public PyTorch framework proposed in  Roth et al.
+(2023) to an equivalent jax variant.
 We verify our implementation through replications of baseline and HFS results in Roth et al.
-[60], mainting relative performance orderings and close absolute disentanglement scores (as measured using DCI-D, whose implementation directly follows from [43] and leverages gradient boosted tree implementations from scikit-learn).
+(2023), mainting relative performance orderings and close absolute disentanglement scores (as measured using DCI-D, whose implementation directly follows from Locatello et al.
+(2019a) and leverages gradient boosted tree implementations from scikit-learn).
 For exact and fair comparison, we utilize standard hyperparamater choices from Roth et al.
-[60] (which leverages hyerparameters directly from [43], [44] and https://github.com/google-research/disentanglement_lib).
-Consequently, the base VAE architecture utilized across all experiment is the same as the one utilized in [60] and [44]: With image input sizes of $64\times 64\times N_{c}$ (with $N_{c}$ the number of input image channels, usually 3).
+(2023) (which leverages hyerparameters directly from Locatello et al.
+(2019a), Locatello et al.
+(2020) and https://github.com/google-research/disentanglement_lib).
+Consequently, the base VAE architecture utilized across all experiment is the same as the one utilized in Roth et al.
+(2023) and Locatello et al.
+(2020): With image input sizes of $64\times 64\times N_{c}$ (with $N_{c}$ the number of input image channels, usually 3).
 The latent dimensionality, if not otherwise specified, is set to 10.
@@
   • Decoder: MLP(256), [upconv($64,4\times 4$, stride 2) + ReLU] $\times$ 2, [upconv($32,4\times 4$, stride 2) + ReLU], [upconv($n_{c},4\times 4$, stride 2) + ReLU]
-Similar, we retain all training hyperparameters from [60] and [44]: Using an Adam optimizer ([34], $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) and a learning rate of $10^{-4}$.
-Similarly, we utilize a batch-size of $64$, for which we also ablate all baseline methods.
+Similar, we retain all training hyperparameters from (Roth et al., 2023) and (Locatello et al., 2020): Using an Adam optimizer ((Kingma and Ba, 2014), $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) and a learning rate of $10^{-4}$.
+Following Locatello et al.
+(2020), Roth et al.
+(2023) we utilize a batch-size of $64$, for which we also ablate all baseline methods.
 The total number of training steps is set to $300000$.
-As commonly done for this setting [43, 44, 60], we also perform a small grid search over all the hyperparameters. We report the full details in Tab. 5.
-For $\lambda_{enc}$ and $\lambda_{dec}$, we set $\lambda_{enc}=0$ and $\lambda_{dec}=1$ for the Decoder setting and $\lambda_{enc}=1$ and $\lambda_{dec}=0$ for the Encoder setting.
-For the joint one, we set $\lambda_{enc}=0.5$ and $\lambda_{dec}=0.5$.
-All runs run on an RTX 2080TI GPU.
 Table 3: Hyperparameter grid searches for different baseline and proposed methods.
@@
 + GMG $\lambda$ [0.1, 1, 5, 10, 20]
++ Jac $\lambda$ [0.1, 1, 5, 10, 20]
+As commonly done for this setting (Locatello et al., 2019a; 2020, Roth et al., 2023), we also perform a small grid search over all the hyperparameters. We report the full details in Tab. 3.
+For $\lambda_{e}$ and $\lambda_{d}$, we set $\lambda_{e}=0$ for the Decoder setting and $\lambda_{d}=0$ for the Encoder setting while altering the weighting for the other $\lambda$.
+All experiments run on a single RTX 2080TI GPU.
+C.2 Stability Analysis
+[…]
+Appendix D Additional Empirical Results
+In this section, we report additional empirical results revolving around the regularization of encoder.
+First in  D.1, we conduct the stability analysis when regularizing the encoder.
+Then, we report further results of encoder and decoder regularization on DSprites.
+Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.3.
+D.1 Encoder Stability Analysis
+[…]
+Figure 4: Gradient stability analysis on the DST and GMG as Encoder regularizations.
+The cosine similarity is computed between all pairs of gradients $\nabla_{\theta}$ obtained through 5 randomly sampled batches and a fixed network $T_{\theta}$ for each dataset.
+D.2 Encoder Analysis on DSprites
+Table 4: Disentanglement of regularizing the Encoder and the Encoder and Decoder as measured by DCI-D on DSprites. We highlight best, second best, and third best results for each method and dataset.
+DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
+DSprites Higgins et al. (2017)
+Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
++ Enc-DST 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
++ Enc-GMG $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
++ Dec-DST $28.6$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
++ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
+D.3 Towards Decoder-free Disentanglement
+Table 5: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D (Burgess and Kim, 2018).
+[…]
+Recently, works such as Burns et al.
+(2021), von Kügelgen et al.
+(2021), Eastwood et al.
+(2023), Matthes et al.
+(2023), Aitchison and Ganev (2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023).
+However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
+As the goal of geometry preservation connects the data manifold and the latent domain through a minimal distortion objective and is applicable to both the encoder and decoder of a VAE (§3, Table 4), we posit that its application may provide sufficient training signal to learn meaningful representations and encourage disentanglement, eliminating the need for a reconstruction loss and decoder.
+Table 5 shows preliminary results on unsupervised decoder-free disentangled representation learning on the Shapes3D benchmark, where the decoder and associated reconstruction objective have been removed.
+Standard approaches such as $\beta$-VAE or $\beta$-TCVAE collapse and do not achieve measurable disentanglement (DCI-D of $0.0$).
+However, the inclusion of either DST or GMG significantly raises achievable disentanglement and, combined with the $\beta$-TCVAE matching objective, can achieve DCI-D scores of up to $53.5$ without needing any decoder or reconstruction loss.
+While these are preliminary insights, we believe they offer promise for more scalable approaches to unsupervised disentangled representation learning and potential bridges to popular and scalable self-supervised representation learning approaches.
+Note, that here the distortion loss significantly outperforms the GMG.
+This is expected due to the nature of the GMG, as the distortion loss offers a more restrictive and, thus, stronger signal for learning representations, which is necessary in the absence of a reconstruction objective.
+This highlights that while in most scenarios (§ 2.1, Figure 2), the GMG is preferable over the distortion loss, there also exist settings where a more restrictive optimization signal is desirable.
+Appendix E Python Code for the Computation of the Gromov-Monge Gap
+{python}
+import jax import jax.numpy as jnp
+from ott.geometry import costs, geometry from ott.solvers.quadratic import gromov_wasserstein from ott.problems.quadratic import quadratic_problem
+def gromov_monge_gap_from_samples( source: jax.Array, target: jax.Array, cost_fn: costs.CostFn = costs.Cosine(), epsilon: float = 0.1, stat_fn: Callable, # usually computes the mean of the cost matrix **kwargs, ) -> float: """Gromov Monge gap regularizer on samples."""
+# define source and target geometries cost_matrix_x = cost_fn.all_pairs(x=source, y=source) scale_cost_x = stat_fn(scale_cost_x) cost_matrix_x = cost_matrix_x / jax.lax.stop_gradient(scale_cost_x) geom_xx = geometry.Geometry(cost_matrix=cost_matrix_x)
+cost_matrix_y = cost_fn.all_pairs(x=target, y=target) scale_cost_y = stat_fn(cost_matrix_y) cost_matrix_y = cost_matrix_y / jax.lax.stop_gradient(scale_cost_y) geom_yy = geometry.Geometry(cost_matrix=cost_matrix_y)
+# define and solve entropic GW problem prob = quadratic_problem.QuadraticProblem(geom_xx, geom_yy)
+solver = gromov_wasserstein.GromovWasserstein( epsilon=epsilon, **kwargs ) out = solver(prob)
+# compute the distortion induced by the map distortion_cost = jnp.nanmean( (geom_xx.cost_matrix - geom_yy.cost_matrix)**2 )
+# compute optimal (entropic) gromov-monge displacement reg_gw_cost = out.reg_gw_cost ent_reg_gw_cost = reg_gw_cost - 2 * epsilon * jnp.log(len(source))
+# compute gromov-monge gap loss = distortion_cost - ent_reg_gw_cost
+return loss * jax.lax.stop_gradient(scale_cost_x * scale_cost_y)

### v3

@@
 We demonstrate the effectiveness of our approach for disentanglement across four standard benchmarks, outperforming other methods leveraging geometric considerations.
+Code is available at: https://github.com/ExplainableML/GMG
 1 Introduction
@@
 In this work, we demonstrate how to effectively combine geometric desiderata with prior matching within the VAE framework, using optimal transport (OT) theory (Santambrogio, 2015, Peyré and Cuturi, 2019).
-By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2020, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports.
+By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2023, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports.
 Specifically, we consider cost functions $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ on $\mathcal{X}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$ on $\mathcal{Y}$ that encode geometric features such as scaled distances or angles.
@@
   • (iv)
-    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods—from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023)—notably outperforming other methods that leverage geometric constraints.
-2 Background and Related Works
+    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods, from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023).
+2 Background: On Disentanglement, Quadratic-OT and Distortion
 2.1 Disentangled Representation Learning
@@
 Disentanglement through a Geometric Lens.
-Recent research suggests that disentanglement can arise by imposing geometric constraints on representation spaces (Horan et al., 2021).
-Most works (Karras et al., 2020, Huh et al., 2023) promote latent representations to preserve scaled distances between data points.
-Closest to our work, Lee et al.
-(2022) use a Jacobian-based regularization within a VAE, while Nakagawa et al.
-(2023) employ a distortion-based regularization, directly comparing pairwise squared Euclidean distances, within a Wasserstein AE (WAE, Tolstikhin et al.
-(2019)).
+Recent studies have revealed a fundamental connection between geometric structure preservation and disentanglement in learned representations (Gropp et al., 2020, Chen et al., 2020a, Lee et al., 2022, Nakagawa et al., 2023, Huh et al., 2023).
+This connection was theoretically established by Horan et al.
+(2021) proving that unsupervised disentanglement is always feasible when the generative factors are sufficiently non-Gaussian and maintain local isometry to the data.
+Our work builds directly on this insight by developing a learning framework that promotes representations that are as close as possible to being isometric to the data.
+To quantify geometric preservation between spaces of different dimensions, we leverage quadratic OT theory, which originated in Koopmans and Beckmann (1957) and was formalized by Mémoli (2011) as a framework for measuring isometric correspondence between metric spaces.
+We detail these tools in Section 2.2 before showing how they can be used to learn representations in Section 2.3.
+Additionally, in a concurrent work, Sotiropoulou and Alvarez-Melis (2024) also introduced the Gromov-Monge gap.
+However, they do not apply it for disentangled representational learning.
 2.2 Quadratic Optimal Transport
+OT (Peyré and Cuturi, 2019) theory studies efficient ways to map a probability distribution onto another.
+Linear OT formulations, such as the Monge (1781) problem, require domains $\mathcal{X},\mathcal{Y}$ that can be directly compared through a cost function $c(x,y)$ defined between their elements.
+When these distributions lie on incomparable domains, one must instead rely on quadratic formulations of OT (Q-OT), which instead compare geometric structure through intra-domain costs, also known as the Gromov-Monge (GM) and GW problems.
+In the context of representation learning, representation and data spaces are incomparable by design, which necessitates the use of Q-OT in this work.
 Gromov-{Monge, Wasserstein} Formulations.
-OT (Peyré and Cuturi, 2019) involves transferring a probability distribution to another with inductive biases.
-When these distributions lie on incomparable domains, one relies on the Gromov-Monge (GM) and GW problems, also known as quadratic formulations of OT.
-Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$.
-For technical reasons, we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are CPD kernels (see Def.
+Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, equipped with intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$.
+We assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are CPD kernels (Def.
 (A.1)).
-For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
+For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$, two distributions supported on each domain, the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
 $\inf_{T:T\sharp p=q}\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))|^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})\,.$ (GMP)
@@
 Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e.
-distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2020)
+distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2023)
 $\mathrm{GW}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
@@
 As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$.
-In addition to yielding greater convexity (Rioux et al., 2023) and better statistical (Zhang et al., 2023) and properties, entropic regularization improves computational performance.
-We can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (Peyré et al., 2016).
+Entropic regularization improves computational performance, as we can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (see Appendix D.3 for full details).
 This solver has $\mathcal{O}(n^{2})$ memory complexity.
@@
 2).
-Neural GW.
-While for classical OT, numerous neural methods have been proposed (Makkuva et al., 2020, Korotin et al., 2022, Tong et al., 2023, Eyring et al., 2024), the GW setting has received less attention.
-Carrasco et al.
-(2024) propose an approach for solving GM with inner product costs, that involves a min-max-min optimization procedure.
-Klein et al.
-(2024) propose a framework to computeentropic GW couplings.
-Yet, these formulations are end-to-end, and not based on a regularizer that can be flexibly added to a loss function.
-Therefore, they cannot be combined with other models, as we do in this work with the $\beta$-VAE.
-In a concurrent work, Sotiropoulou and Alvarez-Melis (2024) introduced the Gromov-Monge gap.
-However, (i) they do not apply it for disentangled representational learning, and (ii) they do not investigate its properties, as we do in this work.
-3 Disentanglement with the Gromov-Monge gap
-This section outlines our novel OT-based approach to disentanglement through geometric considerations within the VAE framework.
-First, we explore how an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains can be encouraged to preserve predefined geometric features.
+2.3 Distortion in Representation Learning
+Given an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$, we consider how it can be learned to preserve predefined geometric features.
 In a VAE, $T$ may represent either the encoder $e_{\phi}$, which generates latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from these codes.
-In the case of the encoder, the source domain $\mathcal{X}$ corresponds to the data, and the target domain $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped.
+In the case of the encoder, $\mathcal{X}$ corresponds to the data, and $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped.
 Assuming that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, i.e., $e_{\phi}\circ d_{\theta}=\mathrm{I}d$, the preservation of geometric features by either $e_{\phi}$ or $d_{\theta}$ becomes equivalent.
 Therefore, in the following sections, we refer to $T$ as either the encoder or the decoder without loss of generality.
-Outline.
-Building on this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1.
-We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion.
-§3.3 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework.
-This transitions into §3.4 studying (weak) convexity properties of the GMG, as an operator.
-Finally, in §3.5, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
-3.1 From the distortion…
-We encode the geometric features of interest using two cost functions, one for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$.
-We aim for $T$ to preserve the geometric feature, which means that $T$ preserves these costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$.
-In practice, two types of cost function from this family are particularly relevant:
+We encode geometric features using a cost function for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$.
+Ideally, $T$ should preserve geometry, which means that $T$ preserves costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$.
+In practice, two types of costs are often used:
   • [i]
@@
 When $\alpha=1$, we recover the standard definition of an isometry.
-  • [ii]
-[…]
+  • [$i$]
+    Cosine-Similarity: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})\vcentcolon=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{x}^{\prime}}{\|\mathbf{x}^{\prime}\|_{2}}\rangle$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})=\textrm{cos-sim}(\mathbf{y},\mathbf{y}^{\prime})$ similarly.
+On has $\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})=\cos(\theta_{\mathbf{x},\mathbf{x}^{\prime}})$ where $\theta_{\mathbf{x},\mathbf{x}^{\prime}}$ is the angle between $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
+A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ then preserves the angles between the points, i.e.
+it is a conformal map.
+Note that if $T$ is (scaled) isometry (see above), it is a conformal map.
 We refer to these costs via $\mathbf{L2^{2}}$ for [i] with $\alpha=1$, $\mathbf{ScL2^{2}}$ for [i] with $\alpha\neq 1$ and $\mathbf{Cos}$ for [ii].
@@
 $\mathrm{DST}_{r}(T)$ quantifies how much $T$ distorts geometric features encoded by $c_{\mathcal{X}},c_{\mathcal{Y}}$ on the support of $r$, that is, when $\mathrm{DST}_{r}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$.
+Distortion as a Loss for Representation Learning.
 Nakagawa et al.
 (2023) suggest promoting geometry preservation by regularizing the encoder $d_{\theta}$ using the DST, with $\mathbf{ScL2^{2}}$ as costs, and the latent representation as reference distribution, namely $r=e_{\phi}\sharp p_{\textrm{data}}$.
-While they use it within a WAE, translated in our VAE setting, this corresponds to minimizing
+While they use it within a WAE, we translate their objective to VAE setting adopted later in the paper.
+This results in:
 $\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda\,\mathrm{DST}_{r}(d_{\theta})}_{\textrm{(iii) geom.
-preservation}},$ (2)
-where $\lambda>0$.
+preservation}},\quad\lambda>0$ (2)
 Given the choice of the costs, $d_{\theta}$ is distortion-free (i.e., $\mathrm{DST}_{r}(d_{\theta})=0$), if it is a scaled isometry.
+Challenges Arising from a Mixed Loss.
 Since a scaled isometry that maps the prior onto the data distribution may not exist, there is an inherent trade-off between minimizing terms (ii) and (iii), which are responsible for achieving practical disentanglement.
-As these terms cannot be simultaneously minimized to 00, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations.
+As these terms cannot be simultaneously minimized to 0, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations.
 This naturally raises the question of how to avoid this over-penalization.
@@
 In the next section, we introduce a regularizer to fit Gromov-Monge maps, which we will use as a replacement to the DST.
-3.2 …to The Gromov-Monge Gap
+3 Disentanglement with the Gromov-Monge gap
+Building on these geometric preservation principles, we introduce in §3.1 the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion while fitting the marginal constraints.
+§3.2 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework.
+This transitions into §3.3 studying (weak) convexity properties of the GMG, as an operator.
+Finally, in §3.4, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
+3.1 An Efficient Gap Formulation for Distortion
 Figure 1: Sketch of $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})$ for two different maps $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$.
@@
 Recently, Uscidda and Cuturi (2023) introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost.
-Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
+Building on this concept, we replace ”displacement” with ”distortion” to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
 Definition 3.1 (Gromov-Monge gap).
@@
 We consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use plug-in estimators, i.e.
-$\begin{split}\mathrm{DST}_{{r}_{n}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (3)
+[…]
 and $\mathrm{GMG}_{{r}_{n}}(T)=\mathrm{DST}_{{r}_{n}}(T)-\mathrm{GW}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$.
-Algorithm 1 GMG$(\mathbf{x}_{1},\dots\mathbf{x}_{n},T,\varepsilon)$.
 […]
-2:$\mathbf{t}_{1},\dots\mathbf{t}_{n}\leftarrow T(\mathbf{x}_{1}),\dots,T(\mathbf{x}_{n})$.
-3:$\mathbf{C}_{\mathcal{X}}\leftarrow[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{1\leq i,i^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{X}})$
-4:$\mathbf{C}_{\mathcal{Y}}\leftarrow[c_{\mathcal{Y}}(\mathbf{t}_{j},\mathbf{t}_{j^{\prime}})]_{1\leq j,j^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{Y}}$)
-5:$\textsc{DST}\leftarrow\texttt{mean}((\mathbf{C}_{\mathcal{X}}-\mathbf{C}_{\mathcal{Y}})^{2})$triangleright $n^{2}$
-6:$\varepsilon\leftarrow\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$triangleright usually $\mathcal{O}(n^{2})$
-7:$\textsc{GW}\leftarrow\texttt{GW}(\mathbf{C}_{\mathcal{X}},\mathbf{C}_{\mathcal{Y}},\varepsilon)$triangleright $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$
+1:Require: $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\sim r$; map $T$; entropic solver GW, entropic regularization scale $\varepsilon_{0}$ (default =0.1=0.1), statistic operator on cost matrix stat (default $\texttt{mean}$).
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 8:return DST - GW
@@
 We use the implementation provided by ott-jax (Cuturi et al., 2022).
-In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$,$\mathbf{C}_{\mathcal{Y}}$.
+In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$$\mathbf{C}_{\mathcal{Y}}$.
 We define a scale $\varepsilon_{0}$ and set $\varepsilon=\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$.
@@
 B.2.
-It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013), then extended to the entropic GW solver (Cuturi et al., 2022).
-Additionally, as suggested by Scetbon et al.
-(2022), to promote the solver’s stability, we always initialize it with $\mathbf{P}_{\textrm{init}}=\tfrac{1}{n^{2}}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$.
+It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013).
 Computational Complexity.
@@
 1.
+We stress that, in all cases and for any cost function, the time complexity depends linearly on the dimensions of the source and target spaces, $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$, making the GMG scalable to high-dimensional distributions.
 3.3 (Weak) Convexity of the Gromov-Monge gap
-[…]
+As laid out, the GMG can be used as a regularization loss to push any model $T$ to move distributions with minimal distortion.
+A natural question arises: is this regularizer convex?
+In the following, we study the convexity of $T\mapsto\mathrm{GMG}_{r}(T)$ and its finite-sample counterpart $T\mapsto\mathrm{GMG}_{{r}_{n}}(T)$.
+We focus on the costs $\mathbf{L2}$ and $\mathbf{Cos}$.
+For simplicity, we replace $\mathbf{Cos}$ with $\langle\cdot,\cdot\rangle$, as these costs are equivalent, up to normalization of $r$ and $T$.
+We respectively denote by $\mathrm{GMG}_{r}^{2}$ and $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$ the GMG for these costs.
+We start by introducing a weaker notion of convexity, previously defined on $\mathbb{R}^{d}$ (Davis et al., 2018), which we extend here to $L_{2}(r)\vcentcolon=\{T\,|\,\|T\|_{L_{2}(r)}^{2}\vcentcolon=\int_{\mathcal{X}}\|T(\mathbf{x})\|_{2}^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})<+\infty\}$.
 Definition 3.3 (Weak convexity.).
@@
 We explore this strategy in experiments § 4.2.
-[…]
+Figure 2: Learning of geometry-preserving maps with the DST and the GMG.
+Provided a mixture of Gaussian source distribution $p$ and a uniform or circular target distribution $q$, we minimize a fitting loss together with a geometry-preserving regularization.
+Each line correspond to a different data setup and cost.
+As costs on the top line, we use $\mathbf{L2^{2}}=\|\cdot-\cdot\|_{2}^{2}$, to preserve distances between points while on the bottom line, we use $\mathbf{Cos}=\textrm{cos-sim}(\cdot,\cdot)$, to preserve angles.
 4 Experiments
@@
 In Fig.
-2, we start by illustrating the difference between the DST and GMG on synthetic data.
-We train $T_{\theta}$ to map from a $3$D source $p$ to a $2$D target distribution $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019) and $\mathrm{R}$ a geometry-preserving regularizer.
+2, we start by illustrating the difference between the DST and GMG.
+We train $T_{\theta}$ to map from a 33D source $p$ to a 22D target $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019).
 We compare three settings, [i] no regularization $\mathrm{R}=0$, [ii] $\mathrm{R}=\mathrm{DST}_{p}$ as regularizer , and [iii] $\mathrm{R}=\mathrm{GMG}_{p}$.
@@
 Experimental Setup.
-Having demonstrated how the DST enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning.
+Having demonstrated how the GMG enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning.
 Our primary goal is to investigate whether the GMG results in enhanced disentanglement compared to the DST by efficiently combining (i) prior matching with (ii) geometric constraints on the representation space.
-Moreover, we aim to determine which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose in the geometric regularization and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$.
-The experimental setup is described below.
+Moreover, we aim to determine which $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$.
   • Baselines.
@@
 We report mean and standard deviation over 5 seeds.
-  • Datasets. We benchmark over four standard datasets: Shapes3D (Kim and Mnih, 2018), DSprites Higgins et al. (2017), SmallNORB LeCun et al. (2004), and Cars3D Reed et al. (2015).
-[…]
+  • Datasets. We benchmark over four $64\times 64$ image datasets: Shapes3D (Kim and Mnih, 2018), DSprites (Higgins et al., 2017), SmallNORB (LeCun et al., 2004), and Cars3D (Reed et al., 2015).
+  • Hyperparameters.
+To ensure a fair experimental comparison, we follow recent works (Locatello et al., 2019a; 2020, Roth et al., 2023) by using the same architecture and hyperparameters.
+We perform a similar small grid search for the weighting terms: $\beta$ for the KL loss and $\gamma$ for HFS.
+Additionally, we include the weighting $\lambda$ for the geometry-preserving regularizer in the grid search.
+Note that we search over the same loss weightings $\lambda$ for DST, GMG, and Jac.
+For all experiments with the GMG, we compute it with Alg.
+1, and systematically use $\varepsilon_{0}=0.1$ and $\texttt{stat}=\texttt{mean}$.
+As a result, $\varepsilon_{0}$ is not included in the grid search.
+We conduct an ablation study on $\varepsilon_{0}$ in App.
+D.3.
+We use a batch size of $n=64$.
+At this scale, the computational cost of compute the GMG loss for a batch is negligible, about 3 milliseconds.
+See App.
+C for full details on hyperparameters.
 Table 1: Effect of different regularization on disentanglement (DCI-D on Shapes3D).
@@
 + Dec-DST $67.4$ $\pm 7.1$ $77.9$ $\pm 4.5$ $93.2$ $\pm 9.7$ $94.5$ $\pm 6.9$
-+ Dec-GMG 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
-[…]
++ Dec-GMG 70.0 $\pm 5.9$ 81.0 $\pm 3.2$ 93.3 $\pm 8.6$ 96.1 $\pm 3.8$
+$\mathbf{Cos}:c_{\mathcal{X}}=c_{\mathcal{Y}}=\textrm{cos-sim}(\cdot,\cdot)$
 + Enc-DST $69.2$ $\pm 9.1$ $77.2$ $\pm 7.5$ $87.7$ $\pm 7.7$ $90.5$ $\pm 5.9$
 + Enc-GMG $70.9$ $\pm 9.5$ $79.6$ $\pm 6.6$ $92.5$ $\pm 5.9$ $93.5$ $\pm 6.9$
-+ Dec-DST 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
-+ Dec-GMG 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
++ Dec-DST 76.8 $\pm 4.1$ 81.3 $\pm 4.7$ $87.5$ $\pm 3.3$ $91.9$ $\pm 9.4$
++ Dec-GMG 82.1 $\pm 4.5$ 83.7 $\pm 8.8$ 95.7 $\pm 5.8$ 96.9 $\pm 4.9$
 Which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ should we choose?
@@
 With $\mathbf{Cos}$ costs $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
-DSprites Higgins et al. (2017)
-Base $26.2$$\pm 18.5$ ${32.3}$$\pm 19.3$ $33.6$$\pm 17.9$ $48.7$$\pm 10.2$
+DSprites (Higgins et al., 2017)
+Base $26.2$ $\pm 18.5$ 32.3{32.3} $\pm 19.3$ $33.6$ $\pm 17.9$ $48.7$ $\pm 10.2$
 + Dec-DST $\underline{28.6}$ $\pm 19.3$ $32.4$ $\pm 8.5$ $39.3$ $\pm 18.1$ $49.0$ $\pm 11.2$
-+ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
-SmallNORB LeCun et al. (2004)
++ Dec-GMG 39.5 $\pm 15.2$ 42.2 $\pm 3.6$ 46.7 $\pm 2.0$ 50.1 $\pm 8.5$
+SmallNORB (LeCun et al., 2004)
 Base $26.8$ $\pm 0.2$ $29.8$ $\pm 0.4$ $26.8$ $\pm 0.2$ $29.8$ $\pm 0.4$
-+ Dec-DST $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$ $28.2$$\pm 0.3$ 29.9 ±$\pm 0.4$
-+ Dec-GMG 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$ 28.3 ±$\pm 0.6$ 29.9 ±$\pm 0.5$
-Cars3D Reed et al. (2015)
++ Dec-DST $28.2$ $\pm 0.3$ 29.9 $\pm 0.4$ $28.2$ $\pm 0.3$ 29.9 $\pm 0.4$
++ Dec-GMG 28.3 $\pm 0.6$ 29.9 $\pm 0.5$ 28.3 $\pm 0.6$ 29.9 $\pm 0.5$
+Cars3D (Reed et al., 2015)
 Base $29.6$ $\pm 5.7$ $32.3$ $\pm 4.6$ $29.6$ $\pm 5.7$ $32.3$ $\pm 4.6$
 + Dec-DST $26.8$ $\pm 3.6$ $33.7$ $\pm 4.2$ $26.8$ $\pm 3.6$ $33.7$ $\pm 4.2$
-+ Dec-GMG 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$ 30.1 ±$\pm 5.6$ 36.4 ±$\pm 5.7$
++ Dec-GMG 30.1 $\pm 5.6$ 36.4 $\pm 5.7$ 30.1 $\pm 5.6$ 36.4 $\pm 5.7$
 GMG Consistently Enhances Disentanglement.
@@
 We report full results in Table 2.
-Again we observe that the GMG outperforms or performs equally well to its distortion equivalent, confirming the benefits of accounting for the optimal possible mapping in the regularization.
-Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding an HFS regularization and obtained the best results without it.
-We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets versus not using any isometric regularization.
+Again we observe that the GMG outperforms or performs equally well to its distortion equivalent.
+Note that for SmallNORB and Cars3D, we found no benefits with respect to DCI-D in adding HFS and obtained the best results without it.
+We emphasize that using the GMG with $\mathbf{Cos}$ regularization significantly improves results for all datasets.
 This establishes the GMG as an effective tool for enhanced disentanglement.
@@
 We repeat this procedure for each of the four datasets $\mathcal{D}$ considered and apply the same methodology to the DST.
-Figure 4 presents results for our best-performing configuration, which regularizes the decoder (i.e., $T_{\theta}=d_{\theta}$) and employs the $\mathbf{Cos}$ cost.
-Results for encoder regularization (i.e., $T_{\theta}=e_{\phi}$) are reported in Appendix D.1.
+Figure 4 presents results.
 We observe that the GMG’s gradients exhibit significantly higher alignment compared to those of the DST, demonstrating greater stability of our proposed regularizer.
-This suggests that the GMG ’s formulation, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
+This suggests that the GMG, which accounts for the the minimal distortion, effectively mitigates the inherent variability inherent of the distortion, leading to more stable gradient computations.
 5 Conclusion
@@
 Fabian J.
-Theis consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity.
+Theis consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity
 Appendix
@@
 As a result, if we now replace $\mathbf{C}_{\mathcal{X}}$ and $\mathbf{C}_{\mathcal{Y}}$ by $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$, respectively, the new OT coupling $\tilde{\mathbf{P}}^{\star}$ solving the problem satisfies
-$\begin{split}&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\frac{\mathbf{C}_{\mathcal{X}}}{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\mathbf{P}\frac{\mathbf{C}_{\mathcal{Y}}}{\texttt{stat}(\mathbf{C}_{\mathcal{Y}})},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-\frac{2}{{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})}\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})\cdot\varepsilon H(\mathbf{P})\,,\end{split}$ (13)
+[…]
 where in the last line, we use the fact that $\texttt{stat}(\mathbf{C}_{\mathcal{X}}),\texttt{stat}(\mathbf{C}_{\mathcal{Y}})>0$.
@@
 Recall that
-$\begin{split}\mathrm{GMG}_{{r}_{n},\varepsilon}(T)&=\mathrm{DST}_{r_{n}}(T)-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathrm{DST}_{r_{n}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
-For any coupling $\mathbf{P}\in U_{n}$, since $-\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
+[…]
+For any coupling $\mathbf{P}\in U_{n}$, since $\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
 $\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})<\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}$
@@
 $\mathrm{GMG}_{r_{n},\varepsilon}(T)>\mathrm{GMG}_{r_{n},0}(T)=\mathrm{GMG}_{r_{n}}(T)\geq 0.$
-B.4 Proofs of Prop. 3.3
-See 3.3
+B.4 Proofs of Prop. 3.2
+See 3.2
 Proof.
@@
 ∎
-B.5 Proofs of Thm. 3.5
-See 3.5
-Before proving Thm. 3.5, we first demonstrate some technical results that will be useful later.
+B.5 Proofs of Thm. 3.4
+See 3.4
+Before proving Thm. 3.4, we first demonstrate some technical results that will be useful later.
 Reformulation of the empirical GMG using permutations.
@@
 We denote by $P_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n},\exists\sigma\in\mathcal{S}_{n},\mathbf{P}_{ij}:=\delta_{j,\sigma(i)}\}$ the set set of permutation matrices.
-Assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels (see A.1).
+Assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are conditionally positive kernels (see A.1).
 Then, the GM and GW formulations coincide, in the sense that we can restrict the GW problem to permutations, namely
@@
 that always exists by continuity of the GW objective function on the compact $U_{n}$. We show that $\mathbf{P}^{\star}$ can be chosen as a (rescaled) permutation matrix without loss of generality.
-As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem.
+As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem.
 3), $\mathbf{P}^{\star}$ also solves:
@@
 We continue with the inner product. As noticed by Rioux et al. (2023, Lemma 2)–in the first version of the paper – the GW for inner product costs can be reformulated as:
-$\begin{split}\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)=&\int_{\mathcal{X}\times\mathcal{X}}\langle\mathbf{x},\mathbf{x}^{\prime}\rangle\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})+\int_{\mathcal{Y}\times\mathcal{Y}}\langle\mathbf{y},\mathbf{y}^{\prime}\rangle\mathop{}\!\mathrm{d}q(\mathbf{y})\mathop{}\!\mathrm{d}q(\mathbf{y}^{\prime})\\ +&\min_{\mathbf{M}\in\mathcal{M}}\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+4\|\mathbf{M}\|_{2}^{2}\,,\end{split}$ (19)
+[…]
 where we define $\mathcal{M}=[-M/2,M/2]^{d_{\mathcal{X}}\times d_{\mathcal{Y}}}$ with $M=\sqrt{\int_{\mathcal{X}}\|\mathbf{x}\|_{2}^{2}\mathop{}\!\mathrm{d}p(\mathbf{x})\int_{\mathcal{Y}}\|\mathbf{y}\|_{2}^{2}\mathop{}\!\mathrm{d}q(\mathbf{y})}$.
@@
 If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e.
-00-weakly convex.
-Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $-\lambda_{\min}(\mathbf{A})$-weakly convex.
+0-weakly convex.
+Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $\lambda_{\min}(\mathbf{A})-$weakly convex.
 ∎
@@
 Similarly, we have:
-$-\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
+$\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
 from which we deduce that:
@@
 C.2 Stability Analysis
-[…]
+For the gradient stability analysis experiment, we repeat the following experiment for each of the four image datasets $\mathcal{D}$ that we consider.
+We first considered a fixed neural map $T_{\theta}$, which we choose to be a randomly initialized neural network, consisting of encoder and decoder, before any training.
+For $k=1,\dots,5$, we sample a batch $\mathbf{x}_{1}^{k},\dots,\mathbf{x}_{n}^{k}\sim\mathcal{D}$ and let $r_{n}^{k}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}^{k}}$.
+We report the pariwise cosine similarity between the gradients of the DST and the GMG.
+Formally, we compute $\textrm{cos-sim}(\nabla_{\theta}\mathrm{DST}_{r_{n}^{k}}(T_{\theta}),\nabla_{\theta}\mathrm{DST}_{r_{n}^{l}}(T_{\theta}))$, and $\textrm{cos-sim}(\nabla_{\theta}\mathrm{GMG}_{r_{n}^{k}}(T_{\theta}),\nabla_{\theta}\mathrm{GMG}_{r_{n}^{l}}(T_{\theta}))$, for $k,l=1,\dots,5$.
 Appendix D Additional Empirical Results
@@
 Then, we report further results of encoder and decoder regularization on DSprites.
-Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.3.
+Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.6.
 D.1 Encoder Stability Analysis
-[…]
+We repeat the following experiment for each of the four image datasets $\mathcal{D}$ that we consider.
+We first considered a fixed neural map $T_{\theta}$.
+For $i=1,\dots,5$, we sample a batch $\mathbf{x}_{1}^{i},\dots,\mathbf{x}_{n}^{i}\sim\mathcal{D}$ and let $r_{n}^{i}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{n}^{i}}$.
+We report the pariwise cosine similarity between the gradients of the DST.
+Namely, for $i,j=1,\dots,5$, we compute $\textrm{cos-sim}(\nabla_{\theta}\mathrm{DST}_{r_{n}^{i}}(T_{\theta}),\nabla_{\theta}\mathrm{DST}_{r_{n}^{j}}(T_{\theta}))$, and the GMG, $\textrm{cos-sim}(\nabla_{\theta}\mathrm{GMG}_{r_{n}^{i}}(T_{\theta}),\nabla_{\theta}\mathrm{GMG}_{r_{n}^{j}}(T_{\theta}))$.
 Figure 4: Gradient stability analysis on the DST and GMG as Encoder regularizations.
@@
 DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
-DSprites Higgins et al. (2017)
-Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
-+ Enc-DST 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
-+ Enc-GMG $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
+DSprites (Higgins et al., 2017)
+Base $27.6$ $\pm 13.4$ $36.0$ $\pm 5.3$ 38.7 $\pm 15.7$ $48.1$ $\pm 10.8$
++ Enc-DST 32.8 $\pm 15.0$ $36.5$ $\pm 5.9$ $33.9$ $\pm 15.9$ $48.9$ $\pm 11.1$
++ Enc-GMG $27.5$ $\pm 14.3$ 37.4 $\pm 5.8$ $31.0$ $\pm 14.3$ $45.9$ $\pm 10.9$
 + Dec-DST $28.6$ $\pm 19.3$ $32.4$ $\pm 8.5$ $39.3$ $\pm 18.1$ $49.0$ $\pm 11.2$
-+ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
-D.3 Towards Decoder-free Disentanglement
++ Dec-GMG 39.5 $\pm 15.2$ 42.2 $\pm 3.6$ 46.7 $\pm 2.0$ 50.1 $\pm 8.5$
+D.3 Ablation on the Entropic Regularization Strength
+Effect of $\varepsilon$ on the entropic GW solver.
+In this section, we remind some insights about the entropic GW solver (Peyré et al., 2016) introduced in § 2.2 and used to compute the GMG in Alg.
+1.
+We provide its algorithmic details in Alg.
+2.
+This solver naturally provides a trade-off between (i) the approximation of the true GW distance $\mathrm{GW}(p,q)$ and (i) the optimization speed (i.e., convergence rate).
+This trade-off is controlled by the entropic regularization strength $\varepsilon$.
+  • (i)
+    Optimization.
+The solver employs a mirror descent scheme that iteratively linearizes the entropic GW problem (EGWP) and applies the Sinkhorn algorithm.
+Each mirror descent step corresponds to a projection with respect to the KL divergence, which can be efficiently performed using the Sinkhorn algorithm.
+The solver is initialized with $\mathbf{P}_{0}=\tfrac{1}{n^{2}}\mathbf{1}\mathbf{1}^{\top}$ and iterates as $\mathbf{P}_{t+1}\leftarrow\textsc{Sinkhorn}\left(-\mathbf{C}_{\mathcal{X}}\mathbf{P}_{t}\mathbf{C}_{\mathcal{Y}},\varepsilon\right)$.
+Since a larger $\varepsilon$ in Sinkhorn leads to faster convergence (Cuturi, 2013, Altschuler et al., 2018), this subsequently accelerates each mirror descent step.
+In other words, it reduces the number of iterations within each Sinkhorn call, i.e., the number of inner iterations of the solver.
+Furthermore, Rioux et al.
+(2023) recently demonstrated that increasing $\varepsilon$ enhances the convexity of the entropic GW problem, thereby improving the convergence rate of the mirror descent scheme.
+This provides theoretical justification for using a larger $\varepsilon$ to reduce the number of mirror descent steps.
+In other words, it reduces the number of calls to Sinkhorn, i.e., the number of outer iterations of the solver.
+    We have empirically validated this behavior on DSprites with BetaVAE + GMG.
+We plot the mean amount of inner and outer iterations for the first 3 epochs over 5 seeds for six different values of $\varepsilon_{0}$, which we provide in Figure 5.
+We can observe the expected scaling of increased inner and outer iterations with decreased $\varepsilon_{0}$.
+  • (ii)
+    Approximation. Zhang et al. (2023) show that when using inner products and squared Euclidean distances as costs, and for $\varepsilon\in(0,1]$, the approximation error scales as:
+    $|\mathrm{GW}_{\varepsilon}(p,q)-\mathrm{GW}(p,q)|\lesssim\varepsilon\log(1/\varepsilon),$
+    where the constants in $\lesssim$ depend on $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$, that is, the dimensions of the support of $p$ and $q$, as well as their fourth-order moments.
+Given this trade-off, the practical goal is to select an $\varepsilon$ value that is sufficiently large to ensure fast convergence while avoiding any degradation in performance.
+Across all our experiments, we found that $\varepsilon=0.1$ struck the right balance.
+We validate this observation in the next paragraph.
+Algorithm 2 Entropic Gromov-Wasserstein solver (Peyré et al., 2016), (Scetbon et al., 2022, Alg. 2).
+1:Require: samples $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\sim p$; $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\sim q$; cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$; entropic regularization scale $\varepsilon_{0}$ (default =0.1=0.1), statistic operator on cost matrix stat (default $\texttt{mean}$).
+[…]
+[…]
+[…]
+[…]
+6:while converged do
+7:  $\mathbf{C}_{t+1}\leftarrow-\mathbf{C}_{\mathcal{X}}\mathbf{P}_{t}\mathbf{C}_{\mathcal{Y}}$ $\triangleright$ $n^{3}$ or $n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}})$
+8:  $\mathbf{P}_{t+1}\leftarrow\textsc{Sinkhorn}(\mathbf{C}_{t+1},\varepsilon)$ $\triangleright$ $\mathcal{O}(n^{2})$
+9:end while
+10:Compute $\mathrm{GW}_{\varepsilon}(p_{n},q_{n})$ from $\mathbf{P}_{t}$ using Eq. (EGWP) $\triangleright$ $n^{3}$ or $n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}})$
+11:return $\mathrm{GW}_{\varepsilon}(p_{n},q_{n})$
+Effect of $\varepsilon$ in disentanglement.
+We investigate the effect of the entropic regularization strength $\varepsilon$ used to compute GMG on the disentanglement performances.
+The results are presented in Fig.
+5 and show that performance is robust to the choice of entropic regularization scale $\varepsilon_{0}$.
+We observed this both with respect to a setting, where we see major improvements from the GMG (DSprites with BetaVAE) as well as one where we only observe minor improvements (DSprites with BetaTCVAE).
+This validates our choice to use a single reasonable value for all our experiments, namely $\varepsilon_{0}=0.1$.
+This robustness with respect to the entropic strength was also observed in a recent work (Piran et al., 2024) proposing a similar gap regularization, based on the entropic multi-marginal OT problem.
+See (Piran et al., 2024, Fig.
+3) for experiments highlighting this robustness.
+Our intuition is as follows: when using Sinkhorn (in the case of M3G) or GW/quadratic OT (in our work) to compute a training loss, $\varepsilon$ acts as a sharpness parameter, emphasizing certain pairs of points more strongly.
+While the loss value changes with $\varepsilon$, the optimization of network variables on top of this loss appears to be largely unaffected by the sharpness introduced by $\varepsilon$.
+However, when the output of Sinkhorn or GW is used directly for predictions or learning flows, for example, in Monge maps (Pooladian and Niles-Weed, 2021, Kassraie et al., 2024) or Gromov-Monge maps (Klein et al., 2024), the entropic regularization strength $\varepsilon$ has a much stronger influence and requires careful tuning.
+Figure 5: Analysis of the effect of the entropic regularization scale $\varepsilon_{0}$ on the disentanglement performances when learning with the GMG.
+We use the SmallNORB dataset and consider the $\beta$-VAE (in blue) and $\beta$-TCVAE (in orange) settings.
+The GMG is applied to the decoder $d_{\theta}$ with cost functions $c_{\mathcal{X}}=c_{\mathcal{Y}}=\mathrm{cos}\text{-}\mathrm{sim}(\cdot)$, corresponding to the setting that provides better disentanglement.
+We compute the GMG using Algorithm 1 with stat=mean.
+We investigate the effect of the entropic regularization scale $\varepsilon_{0}$ by testing five other values of $\varepsilon_{0}$ besides the one used in all other experiments in the paper (namely, $\varepsilon_{0}=0.1$).
+The values tested are $\varepsilon_{0}\in\{0.05,0.2,0.3,0.5,1\}$.
+We also include the baseline result without using the GMG (dashed line) as a comparison.
+Additionally, we provide a runtime analysis with respect to both the inner and outer iteration of the GW-solver.
+D.4 Scaling to Higher Image Resolutions
+As detailed in § 3.2, the computation of the GMG scales linearly with the data dimension, enabling our method to handle high-dimensional settings effectively.
+To demonstrate this scalability, we benchmark our approach on the Shapes3D dataset upscaled to $128\times 128$ resolution.
+To accommodate the increased image resolution, we extend the Decoder by adding one additional layer.
+Using our best-performing configuration from the $64\times 64$ experiments—Cos costs and the GMG applied to the decoder $d_{\theta}$—we conduct a focused hyperparameter search as described in Table 5.
+The results, summarized in Table 6, compare the four base models Beta(TC)VAE (+ HFS), both with and without the GMG.
+Our findings indicate that the proposed setup scales to higher resolutions while preserving the performance improvements achieved by the GMG.
+Finally, we visually validate these results by plotting the latent traversals for the best-performing configuration, as shown in Figure 6.
+Table 5: Hyperparameter grid search for Shapes3D $128\times 128$.
+Method Parameter Values
+$\beta$-(TC)VAE $\beta$ [10, 16]
++ GMG $\lambda$ [0.1, 1, 10]
+$\beta$-(TC)VAE $\beta$ [2, 4]
++ GMG $\lambda$ [0.1, 1, 10]
++ HFS $\gamma$ [1, 10]
+Table 6: Impact of the GMG, applied with $\mathbf{Cos}$ as the cost function on the decoder $d_{\theta}$, on disentanglement performance using upscaled $128\times 128$ Shapes3D images.
+Performance is evaluated using DCI-D, with the best result highlighted for each method.
+With $\mathbf{Cos}$ costs $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
+Shapes3D ($128\times 128$) (Burgess and Kim, 2018)
+Base $54.3$ $\pm 19.3$ 74.6{74.6} $\pm 17.3$ $87.2$ $\pm 2.7$ $88.6$ $\pm 11.4$
++ Dec-GMG 63.9 $\pm 8.2$ 82.0 $\pm 12.8$ 90.4 $\pm 3.7$ 92.2 $\pm 8.2$
+D.5 Latent Traversal Visualization
+Figure 6: Latent traversal visualization for Shapes3D $128\times 128$ with our best performing setup, BetaTCVAE + HFS + GMG.
+We select the best performing result out of 5 seeds achieving a DCI-D of $99.4$.
+We plot four different latent dimensions while traversing them from 1.0-1.0 to $1.0$.
+As visualized the model has clearly learned to separate wall hue, object hue, scale, and floor hue into different latent dimensions.
+Figure 7: Latent traversal visualization for Shapes3D $64\times 64$ with our best performing setup, BetaTCVAE + HFS + GMG.
+We select the best performing result out of 5 seeds achieving a DCI-D of $100.0$.
+We plot four different latent dimensions while traversing them from 1.0-1.0 to $1.0$.
+As visualized the model has clearly learned to separate wall hue, object hue, scale, and floor hue into different latent dimensions.
+D.6 Towards Decoder-free Disentanglement
 Table 7: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D (Burgess and Kim, 2018).
-[…]
-Recently, works such as Burns et al.
-(2021), von Kügelgen et al.
-(2021), Eastwood et al.
-(2023), Matthes et al.
-(2023), Aitchison and Ganev (2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023).
+Decoder-free $\beta$-VAE $\beta$-TCVAE Base $0.0$ $\pm 0.0$ $0.0$ $\pm 0.0$ $\mathbf{L2^{2}}:\|\cdot-\cdot\|_{2}^{2}$ + DST $38.2$ $\pm 0.8$ $42.7$ $\pm 1.6$ + GMG $13.9$ $\pm 0.4$ $20.5$ $\pm 0.5$ $\mathbf{ScL2^{2}}:\alpha\|\cdot-\cdot\|_{2}^{2}$, $\alpha>0$ learnable + DST 45.6 $\pm 1.2$ 53.5 $\pm 1.0$ + GMG $15.2$ $\pm 0.3$ $25.2$ $\pm 0.6$ $\mathbf{Cos}:\textrm{cos-sim}(\cdot,\cdot)$ + DST $37.0$ $\pm 0.4$ $46.1$ $\pm 1.5$ + GMG $37.0$ $\pm 0.9$ $38.8$ $\pm 1.1$
+Recently, works such as (Burns et al., 2021, von Kügelgen et al., 2021, Eastwood et al., 2023, Matthes et al., 2023, Aitchison and Ganev, 2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023).
 However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
@@
 from ott.geometry import costs, geometry from ott.solvers.quadratic import gromov_wasserstein from ott.problems.quadratic import quadratic_problem
-def gromov_monge_gap_from_samples( source: jax.Array, target: jax.Array, cost_fn: costs.CostFn = costs.Cosine(), epsilon: float = 0.1, stat_fn: Callable, # usually computes the mean of the cost matrix **kwargs, ) -> float: """Gromov Monge gap regularizer on samples."""
+def gromov_monge_gap_from_samples( source: jax.Array, target: jax.Array, cost_fn: costs.CostFn = costs.Cosine(), epsilon: float = 0.1, stat_fn: Callable, # usually computes the mean of the cost matrix **kwargs, ) -¿ float: ”””Gromov Monge gap regularizer on samples.”””
 # define source and target geometries cost_matrix_x = cost_fn.all_pairs(x=source, y=source) scale_cost_x = stat_fn(scale_cost_x) cost_matrix_x = cost_matrix_x / jax.lax.stop_gradient(scale_cost_x) geom_xx = geometry.Geometry(cost_matrix=cost_matrix_x)